### v1

# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

Abstract
Fine-tuning large language models (LLMs) with low-rank adapters (LoRAs) has become common practice, often yielding numerous copies of the same LLM differing only in their LoRA updates. This paradigm presents challenges for systems that serve real-time responses to queries that each involve a different LoRA. Prior works optimize the design of such systems but still require continuous loading and offloading of LoRAs, as it is infeasible to store thousands of LoRAs in GPU memory. To mitigate this issue, we investigate the efficacy of compression when serving LoRA adapters. We consider compressing adapters individually via SVD and propose a method for joint compression of LoRAs into a shared basis paired with LoRA-specific scaling matrices. Our experiments with up to 500 LoRAs demonstrate that compressed LoRAs preserve performance while offering major throughput gains in realistic serving scenarios with over a thousand LoRAs, maintaining 75% of the throughput of serving a single LoRA.
1MIT CSAIL, 2MIT-IBM Watson AI Lab
1 Introduction
The myriad uses for foundation models (FMs) has led to a proliferation of specialized models, each fine-tuned to perform downstream tasks. The growing number of parameters of these models, however, incurs significant costs both for the fine-tuning and for serving these models.
To avoid fine-tuning giant foundation models, parameter-efficient fine-tuning (PEFT) algorithms update a smaller set of values that define an edit to the original model. An especially successful PEFT method is low-rank adaptation (LoRA) [8], which achieves parameter-efficient fine-tuning by learning low-rank additive changes to neural network matrices. Because of the low-rank parameterization, these matrices (called adapter weights) contain orders-of-magnitude fewer parameters than the base model. LoRA can achieve performance on par with full fine-tuning [8]. In addition, by merging the adapter weights into the base model, the fine-tuned model does not suffer from increased latency.
LoRA’s popularity has triggered a growing need to serve large collections of LoRA adapters at the scale where serving multiple copies of the base model becomes infeasible [25]. In the extreme, one may wish to serve hundreds to thousands of LoRAs, where each incoming request requires a different LoRA adapter, e.g., for a personalized chat assistant. To this end, S-LoRA [25] separates the base model computation from individual LoRA computations and optimizes the system’s inner workings via custom CUDA kernels and memory management to increase the throughput when serving multiple LoRAs. Despite the optimized system design, serving LoRAs still has a fundamental limitation: when the number of adapters is large, they need to be constantly loaded and offloaded from GPU memory to accommodate incoming requests, degrading throughput.
In this work, we consider the problem of compressing a collection of LoRAs. We have two key objectives: (1) preserving the performance of the original LoRAs and (2) improving the throughput of serving many LoRAs. We formulate LoRA compression as a reconstruction problem, where the goal is to approximate the original adapters via collections of matrices of a smaller total size. We investigate two approaches: compressing each LoRA individually by lowering its rank via SVD and compressing LoRAs jointly by finding a shared basis and LoRA-specific scaling matrices. The former is inspired by [11], who demonstrated that lowering LoRA ranks is beneficial for multi-task learning and model merging. For the latter, we propose a joint compression algorithm inspired by models for joint diagonalization.
Figure 1: Throughput gains when serving 1000s of compressed LoRAs with vLLM.
Our individual and joint compression methods present interesting trade-offs. For example, joint compression can achieve extreme parameter count reduction when representing any specific LoRA by using a shared basis that can be pre-loaded onto the GPU. The shared basis, however, will typically increase the rank of LoRAs. On the other hand, individual compression can reduce the rank, but is limited in terms of its compression efficiency. We investigate how these trade-offs affect throughput when serving 1000s of LoRAs using vLLM [12], a state-of-the-art LLM serving system. In Figure 1 we illustrate the benefits of compression and the effect of these trade-offs. When serving over 1000 LoRAs, compression increases throughput over 2 times and maintains 75% of the throughput of serving the base LLM (or a single LoRA merged into the LLM).
We summarize our main contributions below:
  • We formulate the problem of compressing a collection of LoRAs and propose two solutions: individual and joint compression.
  • We establish theoretical guarantees for the reconstruction error central to our compression formulation and verify the relation between reconstruction loss and performance empirically.
  • We train a collection of 500 high-quality LoRAs for Mistral-7B-Instruct-v0.2 [10] on 500 natural instruction tasks [27] and demonstrate that our compression techniques can preserve the performance of the original LoRAs. We will release the 500 LoRAs to facilitate future work on LoRA compression as well as the code to compress new models.
  • We incorporate LoRA compression into a state-of-the-art LLM serving system and demonstrate that it is possible to serve over 1000 LoRAs across thousands of asynchronous requests with throughput comparable to serving a single LoRA.
  • We analyze the trade-offs between individual and joint compression, presenting opportunities for further improvement of LLM serving systems.
2 Related Work
Parameter-efficient fine-tuning (PEFT) has become a prevalent method for updating foundation models thanks to the need for efficiency in training and communication [14]. Many PEFT methods have been proposed, e.g. [7, 16], but the most used in practice is LoRA [8], partially due to the ease of switching between LoRAs in inference time.
Several works propose improvements to LoRA [17, 26], sometimes with algebraic methods like SVD [19, 34, 11] or by leveraging its statistical properties [35, 33]. Relatively few methods, however, accelerate inference times. S-LoRA [25] provides an efficient means of switching between LoRAs. F-LoRA [28] adapts training to reduce batch multiplications and thus to accelerate inference. Our method achieves a similar outcome (see Appendix E) without changing the LoRA formulation or requiring that LoRAs be trained in a dedicated way; from this perspective, an advantage of our work is that future generic improvements to LoRA will also benefit (e.g., [19]).
There are many efforts to compress models [1, 5, 24, 13]—including some that consider LoRAs specifically—to accelerate inference. Predominantly, pruning and sparsification methods delete some of the weights [31], and quantization methods reduce the precision of the weights [4]. Some works compress weights to reduce model size but typically require decompression and hence do not save GPU memory [6]. Similarly to our work, while most methods increase speed at the cost of performance, a few note increased performance and generalization after compression [31, 21, 6, 24].
Our work also relates to model merging [3, 30, 18] and mixture of experts methods [20]. Such methods reuse models trained by others [2, 22], serving them together as one compressed model. Despite this similarity, these methods create a single general model that acts on any input, while our model allows for more performant per-task solutions.
3 Rank-Based LoRA Compression
LoRA updates are parameterized by pairs of matrices $A,B$, whose product $BA$ updates the fixed weight matrices $W_{0}\in\mathbb{R}^{d_{B}\times d_{A}}$ of a neural network foundation model. Given an input $x$ to a layer, the output of the LoRA-updated model at this layer is $(W_{0}+BA)x$.
In formulating our compression algorithms, we consider a collection of given LoRA adapters $\{(A_{i},B_{i})\}_{i=1}^{n}$ that we would like to serve. We let $r_{i}$ refer to the rank of the LoRA adapter-pair $(A_{i},B_{i})$, i.e., $B_{i}\in\mathbb{R}^{d_{B}\times r_{i}}$, $A_{i}\in\mathbb{R}^{r_{i}\times d_{A}}$. After laying out the basic considerations in designing a compression scheme in §3.1, we consider a simple baseline parameter reduction approach §3.2 as well as a more aggressive joint diagonalization method in §3.3.
3.1 Desiderata
While our compression technique has access only to a collection of $\{(A_{i},B_{i})\}_{i=1}^{n}$ pairs, in our experiments we will assess the efficacy of compression by comparing how the compressed matrices perform relative to the uncompressed LoRAs on typical data. For this reason, although in this section we optimize a Frobenius norm reconstruction error relative to the product $B_{i}A_{i}$, in reality this is a proxy for the nonlinear and complex way that compression errors in the adapters impact transformer performance. Our experimental evaluation will thus focus on the performance of the compressed LoRAs against the uncompressed versions on real data in §6.
Our compression methods significantly reduce the overall number of parameters. Reducing parameters through compression theoretically accelerates storage and serving processes for a collection of LoRAs. This reduction, however, alters the computational dynamics during inference, so parameter reduction alone does not immediately imply faster throughput. In light of the complexities of GPU optimization, we experimentally assess the throughput under realistic conditions in §6.3.
3.2 Independent Compression via SVD
The simplest approach is to replace each rank-$r_{i}$ LoRA adapter $B_{i}A_{i}$ with a reduced rank-$r$ approximation, where typically $r<\frac{1}{n}\sum_{i=1}^{n}r_{i}$:
$\operatorname{SVD}_{r}(B_{i}A_{i})=U_{i}\Sigma_{i}V_{i}^{\top},\quad\forall i=1,\dots,n~{}~{}~{}~{}~{}~{}(\mathsf{r-SVD})$ (1)
Thanks to the Eckart-Young Theorem, this strategy finds the rank-$r$ approximation that best fits $B_{i}A_{i}$ in terms of Frobenius norm. As $\Sigma_{i}V_{i}^{\top}$ can be saved as a single matrix, this approach reduces the number of parameters used from $\sum_{i}(d_{A}+d_{B})r_{i}$ to $rn(d_{A}+d_{B})$. While this approximation is effective, it is limited in its compression abilities as $n$ increases, since it does not share information across LoRAs. This shortcoming motivates our next approach.
3.3 Joint Diagonalization
Next, we suggest a Joint Diagonalization (JD) method, which optimizes a shared basis onto which we can project the set of $n$ LoRAs. This will allow structure to be shared, implicitly clustering the collection of LoRAs.
In this model, each LoRA-product $B_{i}A_{i}$ is factorized into the form $U\Sigma_{i}V$, where $U$ and $V$ are shared across all LoRAs and $\Sigma_{i}$ is specific to each LoRA. In this formulation, every $\Sigma_{i}$ shares the same rank $r$. This allows $U$ and $V$ to be pre-loaded onto the GPU, with $\Sigma_{i}$ loaded when necessary for each batch. The matrices $\Sigma_{i}$ can be either diagonal or small square matrices, accelerating the forward pass compared to conventional multi-LoRA serving configurations.
Objective function. Motivated by the relationship of singular value decomposition to minimizing the Frobenius norm of the reconstruction error, we also propose to minimize the Frobenius norm of the adapter matrix approximation error. Specifically, we use the following objective function:
$\min_{\{\Sigma_{i}\}_{i=1}^{n},U,V}\sum_{i=1}^{n}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}.$ (2)
Using the Frobenius norm has the added benefit of making the objective convex in each argument separately, suggesting the possibility of efficient optimization. This objective function is underdetermined, however, so we consider two constrained regimes below.
Full $\Sigma_{i}$ approximation. The first method we call $\mathsf{JD-Full}$. Without loss of generality, $U$ and $V$ can be constrained to be orthogonal, so long as $\Sigma_{i}$ remains an unconstrained full matrix. $\mathsf{JD-Full}$ adopts this restriction to make the optimization better posed, but note it does not restrict the expressiveness of the objective (2). This setting yields the following optimization problem:
$\operatorname{JD-Full}_{r}(\{B_{i}A_{i}\}_{i=1}^{n})=\underset{U^{\top}U=V^{\top}V=I_{r}}{\underset{\{\Sigma_{i}\}_{i=1}^{n}}{\operatorname{argmin}}}\sum_{i=1}^{n}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}~{}~{}~{}~{}(\mathsf{JD-Full})$ (3)
An efficient alternating algorithm to solve this objective function can be found in Appendix B.
Diagonal $\Sigma_{i}$ approximation. As an alternative, we can leave $U$, $V$ unconstrained (other than to have $r$ columns) and instead constrain the matrices $\Sigma_{i}$ to be diagonal (but not necessarily positive). This formulation yields the following optimization problem:
$\operatorname{JD-Diag}_{r}(\{B_{i}A_{i}\}_{i=1}^{n})={\underset{\{\Sigma_{i}\}_{i=1}^{n},U,V}{\operatorname{argmin}}}\sum_{i=1}^{n}\|B_{i}A_{i}-U\mathrm{diag}(\Sigma_{i})V^{\top}\|_{\mathrm{Fro}}^{2}~{}~{}~{}~{}(\mathsf{JD-Diag})$ (4)
An efficient alternating least squares algorithm to optimize this objective can be found in Appendix B. This diagonal version has some per-LoRA parameter savings when compared to $\mathsf{JD-Full}$, since the diagonal $\Sigma_{i}$ only needs $r$ parameters instead of $r^{2}$.
4 Theoretical Analysis
While SVD-based approaches are relatively well-understood, in this section, we seek to better understand the role of the joint diagonalization method presented in §3.3. We will focus on the full-$\Sigma_{i}$ case with orthogonal $U$, $V$ matrices. Note that, for the same $r$, the $r\!\!-\!\!\mathsf{JD\!-\!Diag}$ has at least as large reconstruction error as $r\!\!-\!\!\mathsf{JD\!-\!Full}$ since it imposes an additional constraint on the $\Sigma_{i}$.
Firstly, note that perfect reconstruction can be achieved if and only if $r$ is large enough:
Proposition 1.
Suppose that for all $i$, $\mathrm{rank}(B_{i}A_{i})=r_{i}$, and let
$\tilde{r}=\max\left\{\mathrm{rank}([A_{1},\dots,A_{n}]),\mathrm{rank}([B_{1}^{\top}\dots,B_{n}^{\top}])\right\}.$
Note $\max_{i}r_{i}\leq\tilde{r}\leq\sum_{i=1}^{n}r_{i}$. Then $\mathsf{JD-Full}$ (3) with $r=\tilde{r}$ achieves lossless compression (perfect reconstruction), and using $r<\tilde{r}$ will give nonzero reconstruction error.
Proof.
There exist $U,V$ such that all the $B_{i}$, $A_{i}$ are in the spans of $U,V$ resp. if and only if $r\geq\tilde{r}$. ∎
Due to training noise, $\tilde{r}$ will equal $\sum_{i=1}^{n}r_{i}$ almost always. This implies that in most realistic settings, the joint diagonalization approach is a lossy reconstruction.
This reconstruction loss can be significant, as the following theorem shows (proved in Appendix C):
Theorem 1.
Consider $n$ LoRAs ($\{A_{i},B_{i}\}_{i=1}^{n}$) with $r,n\leq d^{2}$, and form the matrix
$L=\left[\begin{array}[]{ccc}\mathrm{vec}(B_{1}A_{1})&\cdots&\mathrm{vec}(B_{n}A_{n})\end{array}\right].$
Let $\sigma_{j}$ be the singular values of $L$, sorted from largest to smallest, and let $\bar{\sigma}_{j}$ be the singular values of $\sum_{i=1}^{n}B_{i}A_{i}$. Then, using $\mathsf{JD-Full}$ (3),
$\sum_{j=1}^{r}\bar{\sigma}^{2}_{j}\leq\sum_{i=1}^{n}\|\Sigma_{i}\|^{2}_{\mathrm{Fro}}=\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}\|^{2}_{\mathrm{Fro}}\leq\sum_{j=1}^{\min(r^{2},n)}\sigma_{j}^{2},$
implying the sum of squared Frobenius norms of the reconstructed LoRAs satisfies
$\frac{\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}\|^{2}_{\mathrm{Fro}}}{\sum_{i=1}^{n}\|B_{i}A_{i}\|_{\mathrm{Fro}}^{2}}\!\leq\!\frac{\sum_{j=1}^{\min(r^{2},n)}\!\!\sigma_{j}^{2}}{\sum_{j=1}^{n}\sigma_{j}^{2}}\leq 1,\>\>{and}\>\>\>\frac{\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}\!\!-\!B_{i}A_{i}\|^{2}_{\mathrm{Fro}}}{\sum_{i=1}^{n}\|B_{i}A_{i}\|_{\mathrm{Fro}}^{2}}\geq 1-\frac{\sum_{j=1}^{\min(r^{2},n)}\!\!\sigma_{j}^{2}}{\sum_{j=1}^{n}\sigma_{j}^{2}}.$
In other words, if the singular values of $L$ are not concentrated in the top $r^{2}$ entries, significant reconstruction error is unavoidable.
Remark 1 (Lower bound and merging).
The lower bound $\sum_{j=1}^{r}\bar{\sigma}^{2}_{j}$ could be achieved by setting all the $\Sigma_{i}$ equal, i.e., using a fully merged model instead of only merging the subspaces $U$, $V$ and allowing $\Sigma_{i}$ to vary with $i$.
Remark 2 (Upper bound and clustering).
The upper bound is smallest when the LoRAs are relatively clustered, i.e., when groups of vectors $\mathrm{vec}(B_{i}A_{i})$ are similar. This situation raises the magnitude of the largest singular values of $L$, raising the upper bound in the proposition. As the LoRAs are $d\times d$ matrices that can be thought of as points in $d^{2}$ dimensional space, for typical values of $d$ well into the hundreds, it is likely that unrelated LoRAs will be unclustered, i.e., they will have relatively low inner products with each other.
For the case of orthogonal LoRAs, the singular values of $L$ are the norms of the LoRAs, and we immediately have the following corollary: [^1] A result for isotropic Gaussian LoRAs could be obtained via the quantiles of the Marchenko-Pastur Law.
Corollary 1.
Suppose (e.g., due to normalization) that the inputs to the joint diagonalization algorithm all have unit Frobenius norm, i.e., $\|B_{i}A_{i}\|_{\mathrm{Fro}}=1$. Moreover, assume that the LoRAs are all orthogonal in the sense $\mathrm{tr}((B_{i}A_{i})(B_{j}A_{j})^{\top})=0$ for $i\neq j$. Then, using the $\mathsf{JD-Full}$ method (3), we have $1\leq\sum_{i=1}^{n}\|\Sigma_{i}\|^{2}_{\mathrm{Fro}}\leq\min(r^{2},n),$ implying that the sum of squared Frobenius norms of the reconstructed LoRAs satisfies
$1-\frac{1}{n}\geq\frac{\sum_{i=1}^{n}\|U\Sigma_{i}V^{\top}-B_{i}A_{i}\|^{2}_{\mathrm{Fro}}}{\sum_{i=1}^{n}\|B_{i}A_{i}\|_{\mathrm{Fro}}^{2}}\leq 1-\min\left(\frac{r^{2}}{n},1\right).$
This implies that for the common setting where $r^{2}\ll n$, the reconstructed LoRAs will be significantly smaller than the original LoRAs and necessarily have significant reconstruction error.
The results in this section illustrate the tradeoffs of using joint diagonalization. If the LoRAs are similar or well-clustered, reconstruction error will be low. On the other hand, if the LoRAs are random and orthogonal, reconstruction error will be high.
Note that since the loss space of transformers is highly complex, increasing reconstruction error for the weights does not necessarily imply degrading ability to approximate the desired LLM outputs. Interestingly, in an experiment shown in Figure 4 below, we indeed see that while large reconstruction error rapidly decreases performance, moderate (but still relatively large, at around 0.4) reconstruction error does not damage performance and may even slightly outperform the zero reconstruction error setting. This observation motivates our focus on minimizing weight reconstruction error, while also pointing towards the possibility that our approach is capable of achieving something deeper than simply low-loss compression. Specifically, the tendency of the Frobenius-norm based joint diagonalization is to find subspaces that are shared with many LoRAs when $r$ is large, and merge subspaces together towards the mean when $r$ is small. When $r$ is severely below the natural rank of the problem, this tendency towards averaging all or some of the LoRA signals directly connects to the concept of merging LoRAs, whose empirical success in works such as [23, 9] could explain the success of our procedure despite the nonlinearity of the underlying transformers.
Experiments in Appendix F.10 explore this idea further, comparing reconstruction of real-world LoRAs to reconstruction of randomly sampled LoRAs. The reconstruction error is generally large, but significantly lower than the reconstruction error for random noise, indicating that there is a major shared component between the LoRAs that is being successfully retained.
5 Training LoRAs & Evaluating Task Performance
5.1 Training
We trained LoRA adapters on 500 natural instruction tasks [27] using Mistral-7B-Instruct-v0.2 [10] as the base model. All LoRA adapters were configured with a rank of 16, i.e., $\forall i,r_{i}=16$.
We selected 10 diverse tasks (Table 2 in Appendix D) manually for consistent evaluation across experiments and randomly sampled an additional 490 tasks, resulting in a total of 500 tasks. These tasks were exclusively in English (both input and output), ensuring higher quality and thorough review [27]. Each task dataset was divided into training, validation, and test sets.
Hyperparameters, such as early stopping, were tuned using the validation sets. Evaluation on the test sets demonstrated that LoRA consistently outperformed the base model in terms of both Rouge scores and loss metrics, as shown in Table 1. Details are provided in Appendix D.
Table 1: Comparison of metrics before and after LoRA training across 500 tasks.
Metric Base Model LoRA
Loss $4.99\pm 3.11$ $0.43\pm 0.57$
Exact Match $2.28\pm 7.89$ $66.66\pm 34.34$
Rouge1 $20.38\pm 18.90$ $76.74\pm 24.89$
RougeL $19.66\pm 18.16$ $76.22\pm 25.27$
5.2 Evaluation
We evaluated multiple metrics for the natural instruction tasks, including cross-entropy loss, Rouge-1, Rouge-L [15], exact match, and agreement between uncompressed and compressed LoRA. Here, agreement measures the exact match in task-generations between the uncompressed LoRA model and the compressed LoRA model, rather than comparing to ground truth data. While detailed results and discussions for all metrics are provided in Appendix F, our primary focus in the main text is on Rouge-L. We find that all metrics correlate, but Rouge-L correlates most strongly with downstream utility. This finding aligns with prior work [27], which demonstrates that Rouge-L correlates well with classification accuracy.
While cross-entropy is used for optimization during training, identical generation outputs across models can yield different cross-entropy losses. Exact match is too rigid and does not account for the variability in task responses. Similarly, agreement does not capture the inexactness associated with most of our tasks, nor does it account for the performance gains or losses of the compressed LoRAs. Arguably, practitioners are primarily concerned with task performance in the settings for which the LoRA was designed, rather than exact generational agreement between models.
Joint diagonalization optimizes reconstruction error measured by the Frobenius norm, and our theoretical analysis in §4 bounds this reconstruction error. Empirically, reconstruction error and downstream Rouge-L performance correlate.
Instead of listing the absolute performance of different methods, we compute the performance difference between the base model and the LoRA model for each task. We present the ratio
$\text{Performance relative to LoRA}:=\frac{\text{method-performance}}{\text{LoRA-performance}}$
for the specific method in question, highlighting relative improvement with respect to the uncompressed LoRA and the base model.
6 Experiments
6.1 Task Performance
For each method, we vary the number of $n$ LoRAs that are compressed and the compression rank $r$. We run each experiment three times with different random seeds and report the mean and standard deviation. See Table 3 for results where we evaluate on the same ten manually selected tasks (Table 2) across settings. Every compressed collection of LoRAs contains these 10 tasks (i.e., in-distribution tasks), and each collection contains the smaller collections as subsets.
We normalize each LoRA adapter to have a Frobenius norm of one prior to running joint diagonalization. This normalization enhances performance and reduces the variance in reconstruction error. We restore the original norms of the LoRA adapters before reconstruction and testing.
Figure 2(a) plots the mean performance improvement relative to uncompressed LoRAs ($y$-axis) against the GPU workload parameter saved ratio, i.e., the reduction in per-LoRA parameters ($x$-axis). Similarly, Figure 2(b) relates mean performance improvement to the total parameter saved ratio. Both the independent compression via SVD and joint diagonalization methods significantly compress the LoRAs while preserving—and occasionally enhancing—performance. Notably, the JD methods are unique in their ability to approach the compression efficacy of a single LoRA, although this aggressive reduction in size may decrease performance in larger LoRA collections. As shown in Figure 4, a JD variant achieves the highest throughput. These results provide practitioners with a spectrum of methods and configurations, enabling them to optimize the balance between throughput, compression, and performance according to their specific requirements.
For efficiency, we limited the JD methods to ten iterations instead of pursuing full convergence. While the alternating algorithm quickly reaches an approximation of the minimizer, squeezing out the last few digits of precision takes many more iterations with limited to no performance gain. Appendix F.12 also evaluates an alternative eigenvalue iteration algorithm that more rapidly converges once $U,V$ are close to a minimizer of the optimization problem, with minimal performance differences.
(a) Performance with GPU parameter saved ratio
(b) Performance with total parameter saved rate
Figure 2: We qualitatively illustrate the performance of JD-LoRA. The GPU workload parameter saved ratio indicates the number of parameters saved compared to LoRA during inference when switching between different LoRAs. For JD-Diagonal methods, it is computed as $r_{pergpu}:=1-\frac{r_{JD}}{r_{LoRA}*d*2}$. The total parameter saved ratio depicts the number of parameters saved for a system with a large number $K$ of LoRAs for different tasks. It is computed as: $r_{total}:=1-\frac{2*r_{JD}*d+K*r_{JD}}{K*r_{LoRA}*d*2}$, where $r_{JD}$ and $r_{LoRA}$ are the rank for JD-LoRA and LoRA respectively, $d$ is the weight dimension, and $K$ is the number of LoRAs served.
6.2 Performance and Reconstruction Error
Figure 4 relates reconstruction error and performance. The $y$-axis measures mean performance improvement of Rouge-L relative to uncompressed LoRA, and the $x$-axis quantifies the mean reconstruction error between the compressed reconstruction of the product $BA$ and the original uncompressed product $BA$. Although the relationship between performance and reconstruction error is nonlinear, it demonstrates a generally decreasing, somewhat exponential trend. Notably, the minimal reconstruction error does not correlate with optimal performance, indicating that a degree of lossy reconstruction may be advantageous for enhancing generalization.
Figure 3: Reconstruction error vs. performance.
Figure 4: Throughput when serving varying number of LoRAs with vLLM.
6.3 Throughput of Serving Compressed LoRAs
Figure 4 studies the efficacy of compression in a real-world serving scenario. We consider a varying number of rank-16 LoRAs, using a dataset of Shakespeare sonnets as inputs [^2] https://www.kaggle.com/datasets/shivamshinde123/william-shakespeares-sonnet/data arriving asynchronously. We measure throughput, i.e., the number of requests served per second when generating ten tokens per request.
In these experiments, the base model is Mistral 7B Instruct as in the previous experiments; we simulate random LoRAs and assign inputs to LoRAs at random. Experiments were conducted on A100 80GB GPU capped at 30% memory consumption. This is done to reflect cost concerns in practical situations where a service provider might want to serve many LoRAs from cheaper hardware with lower memory than higher-end GPUs. This setting also takes into account the scenario where the LLM is large compared to the size of GPU and yet a provider may want to serve many LoRAs efficiently using the same device.
Although LoRAs are meant as parameter-efficient adapters, when serving over 200 LoRAs, the throughput starts to degrade rapidly, illustrating the challenge of serving many LoRAs. When serving over 500 LoRAs, our compression methods provide over 2×2$\times$2 × improvement in throughput, and at over 1000 LoRAs, JD-Full-64 serves at 75% of the throughput of the base LLM (equivalent to serving a single LoRA after merging it with the LLM).
Next, we analyze trade-offs between the compression methods. Each compression method except JD-Full-64 and JD-Diag-128 starts to degrade faster at a certain point as the number of LoRAs increases; this point is the threshold at which compressed LoRAs no longer fit into the GPU memory and the throughput degrades due to scheduling. The JD LoRAs are the most efficient in terms of memory consumption, and we can simultaneously load all 1024 LoRAs. Due to their increased rank (e.g., each LoRA is of rank 128 for JD-Diag-128), however, the throughput starts lower, but degrades at a lower rate. We conclude that for the current vLLM design, when serving a moderate number of LoRAs, SVD compression provides favorable throughput as long as it can fit all LoRAs into memory. When serving a large number of LoRAs that cannot fit into memory even with SVD compression, JD is the better option, as it can fit all LoRAs onto the GPU at the cost of a slight throughput reduction due to its larger rank.
Our analysis above suggests that scheduling LoRAs is an important bottleneck. Compressed LoRAs, especially with JD, significantly reduce the amount of parameters that need to be loaded onto a GPU to process a request requiring an offloaded LoRA. Designing a custom scheduler that can leverage this property is a promising avenue to further improve the throughput when serving compressed LoRAs.
We also note that vLLM extensively uses custom CUDA kernels. To accommodate our compression techniques, we minimally adjusted the vLLM code to generate additional kernels needed by the compressed LoRAs. There likely is room for improvement to optimize the newly added kernels.
Fast LoRA [28] is a recent work that aims to alleviate the batched matrix multiplication (BMM) bottleneck when serving many LoRAs. They propose an adapter parameterization that replaces addition with elementwise multiplication, avoiding BMM and improving LoRA throughput at lower ranks. Our JD LoRA formulation also circumvents or heavily reduces the impact of BMM (see Appendix E), and both individual and joint compression methods can be applied to Fast LoRAs.
7 Discussion
This study introduces approaches to LoRA compression, addressing significant challenges facing foundation models and large language models. Our contributions include theoretical formulations, empirical validation, and practical implementations that enhance understanding and application of LLMs in scalable environments.
The implications of our findings are manifold. Our theoretical guarantees for reconstruction error not only increase confidence in the use of compressed models but also lay a groundwork for future explorations in this area. Demonstrating that our compression techniques can preserve up to 100% of the original LoRAs’ performance highlights the effectiveness of our methods. Furthermore, integrating LoRA compression into state-of-the-art LLM serving systems demonstrates the potential for resource optimization, with throughput for thousands of LoRAs nearing that of a single LoRA.
A primary limitation involves integration of new LoRAs. Our findings indicate that while compressing a new LoRA within the original space improves upon the baseline, it can degrades performance compared to uncompressed LoRAs (see Appendix F.11). Thus, the optimal strategy varies with the frequency of new additions. Infrequent additions justify recompressing all LoRAs given the minimal impact relative to inference costs. Conversely, in scenarios where a proportion of new LoRAs are often introduced but seldom served, one should compress only frequently-served LoRAs and otherwise use the original LoRAs.
The promising results of our study suggest several future research directions. Extending our compression techniques to a wider array of models and tasks could further substantiate and expand the generalizability of our findings. This includes out-of-distribution evaluations, which can highlight generalization improvements associated with compression. Moreover, more sophisticated optimization algorithms could improve the balance between compression and performance.
In conclusion, our research significantly advances the deployment of LLMs by providing robust, scalable, and efficient compression solutions. The ability of compressed LoRAs to maintain high performance while facilitating substantial resource savings opens new avenues for the broader application and adoption of LLMs across various industries. We encourage the community to build upon our findings and the shared LoRAs to further explore and enhance the utility of these technologies.
8 Acknowledgements
The MIT Geometric Data Processing Group acknowledges the generous support of Army Research Office grants W911NF2010168 and W911NF2110293, from the CSAIL Systems that Learn program, from the MIT–IBM Watson AI Laboratory, from the Toyota–CSAIL Joint Research Center, and from an Amazon Research Award.
Appendix A Limitations
Compression techniques, while reducing time or memory constraints, invariably lose information, which can increase with the number of models compressed. Consequently, while our approach seeks to minimize this loss, it is axiomatic that compression will diminish information to some degree.
Our methodology compresses a collection of known models. Introducing new models may necessitate either recompression, projecting into a suboptimal basis, or opting to use some models in an uncompressed state.
Empirically, our methods achieve reasonable balances between the number of models compressed, the amount of information retained, and throughput. Nonetheless, these benefits can be accompanied by increased latency as more models are compressed.
Appendix B Joint Diagonalization Algorithms
B.1 Alternating Methods
Our goal is to derive algorithms that optimize (2). Common to both methods, we expand the objective functional:
$\sum_{i}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}$ $=\sum_{i}\mathrm{tr}((B_{i}A_{i}-U\Sigma_{i}V^{\top})(B_{i}A_{i}-U\Sigma_{i}V^{\top})^{\top})\textrm{ by definition}$
$=\sum_{i}\left[\mathrm{tr}(B_{i}A_{i}A_{i}^{\top}B_{i}^{\top})-2\mathrm{tr}(B_{i}A_{i}V\Sigma_{i}^{\top}U^{\top})+\mathrm{tr}(U\Sigma_{i}V^{\top}V\Sigma_{i}^{\top}U^{\top})\right]$
[…]
Using this expansion, we now consider the two settings discussed in §3.3.
Case 1: Non-diagonal $\Sigma_{i}$, orthogonal $U,V$. Setting the derivative of (5) with respect to $\Sigma_{i}$ to zero, we find
$\Sigma_{i}=\Sigma^{\ast}_{i}(U,V)=U^{\top}B_{i}A_{i}V.$ (6)
We simplify our objective function after plugging in this expression:
[…]
$=-\sum_{i}\mathrm{tr}(B_{i}A_{i}VV^{\top}A_{i}^{\top}B_{i}^{\top}UU^{\top}).$
Substituting (6), we find
$U_{opt},V_{opt}=\arg\max_{\begin{subarray}{c}U^{\top}U=I\\ VV^{\top}=I\end{subarray}}\ \sum_{i=1}^{n}\|U^{\top}B_{i}A_{i}V\|_{\mathrm{Fro}}^{2}=\arg\max_{\begin{subarray}{c}U^{\top}U=I\\ VV^{\top}=I\end{subarray}}\ \sum_{i=1}^{n}\|\Sigma^{\ast}_{i}(U,V)\|_{\mathrm{Fro}}^{2}.$ (7)
Note that
$\sum_{i=1}^{n}\|U^{\top}B_{i}A_{i}V\|_{\mathrm{Fro}}^{2}$ $=\mathrm{tr}\left(\left(\sum_{i=1}^{n}B_{i}A_{i}VV^{\top}A_{i}^{\top}B_{i}^{\top}\right)UU^{\top}\right)$
$=\mathrm{tr}\left(\left(\sum_{i=1}^{n}B_{i}^{\top}A_{i}^{\top}UU^{\top}A_{i}B_{i}\right)VV^{\top}\right),$
by the identity $\|A\|_{\mathrm{Fro}}^{2}=\mathrm{tr}(A^{\top}A)$. Hence, we optimize (7) by alternating between $U$ and $V$:
  • $U$ iteration: Define $M\coloneqq\sum_{i}B_{i}A_{i}VV^{\top}A_{i}^{\top}B_{i}^{\top}$. Parenthesizing this expression properly requires only $O((m+n)r)$ storage/computation time. With this definition, we maximize $\mathrm{tr}(MUU^{\top})$ over $U$ satisfying $U^{\top}U=I$. Since $M$ is positive semidefinite, the optimum is to take $U$ to be the $r$ eigenvectors of $M$ with largest eigenvalue, equivalent to an SVD problem.
  • $V$ iteration: Define $N\coloneqq\sum_{i}A_{i}^{\top}B_{i}^{\top}UU^{\top}B_{i}A_{i}$. Similarly to the previous step, we take $V$ to contain the $r$ eigenvectors of $N$ with largest eigenvalue, again solvable using an SVD.
This method decreases the objective in each step.
Case 2: Diagonal $\Sigma_{i}$. If constrain $\Sigma_{i}$ to be diagonal, we interpret our objective function (2) as a “triple least squares” problem. We compute gradients:
$\nabla_{U}\sum_{i}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}$ $=2\sum_{i}(U\Sigma_{i}V^{\top}-B_{i}A_{i})V\Sigma_{i}^{\top}$
$\nabla_{V}\sum_{i}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}$ $=2\sum_{i}(V\Sigma_{i}^{\top}U^{\top}-A_{i}^{\top}B_{i}^{\top})U\Sigma_{i}$
$\nabla_{\Sigma_{i}}\sum_{i}\|B_{i}A_{i}-U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}$ $=2U^{\top}(U\Sigma_{i}V^{\top}-B_{i}A_{i})V$
These expressions suggest efficient $r\times r$ linear systems to solve for $U,V$:
$U$ $=\left(\sum_{i}B_{i}A_{i}V\Sigma_{i}^{\top}\right)\left(\sum_{i}\Sigma_{i}V^{\top}V\Sigma_{i}^{\top}\right)^{-1}$
$V$ $=\left(\sum_{i}A_{i}^{\top}B_{i}^{\top}U\Sigma_{i}\right)\left(\sum_{i}\Sigma_{i}^{\top}U^{\top}U\Sigma_{i}\right)^{-1}.$
For $\Sigma_{i}$, we extract the diagonal from our gradient above:
$\mathrm{diag}(U^{\top}U\Sigma_{i}V^{\top}V)_{j}$ $=(U^{\top}U\Sigma_{i}V^{\top}V)_{jj}$
$=\sum_{m}(U^{\top}U)_{jm}\Sigma_{imm}(V^{\top}V)_{mj}$
$=(U^{\top}U\circ V^{\top}V)\mathrm{diag}(\Sigma_{i})$
$\mathrm{diag}(U^{\top}B_{i}A_{i}V)_{j}$ $=\sum_{m}(U^{\top}B_{i})_{jm}(A_{i}V)_{mj}$
$=\sum_{m}(U^{\top}B_{i})_{jm}(V^{\top}A_{i}^{\top})_{jm}$
$=(U^{\top}B_{i}\circ V^{\top}A_{i}^{\top})\mathbbm{1}$
$\implies\mathrm{diag}(\Sigma_{i})$ $=(U^{\top}U\circ V^{\top}V)^{-1}(U^{\top}B_{i}\circ V^{\top}A_{i}^{\top})\mathbbm{1}$
Here $\circ$ denotes the Hadamard product.
Combining these expressions, we use a simple coordinate descent algorithm cycling between the following three steps:
  1. 1.
    Solve for $U$
  2. 2.
    Solve for $V$
  3. 3.
    Solve for the $\Sigma_{i}$’s
  4. 4.
    Optionally, normalize so $\sum_{i}\|\Sigma_{i}\|_{\mathrm{Fro}}^{2}=1$
B.2 Additional Eigenvalue Iteration Algorithm
For the first case in §B.1, we introduce an alternative algorithm that eschews the use of SVD. This alternative is optimized for GPU execution, enabling tractable runs to convergence.
To derive this algorithm, we employ Lagrange multipliers to formulate the derived objective from (7):
$U_{opt},V_{opt}=\arg\max_{\begin{subarray}{c}U^{\top}U=I\\ VV^{\top}=I\end{subarray}}\ \sum_{i=1}^{n}\|U^{\top}B_{i}A_{i}V\|_{\mathrm{Fro}}^{2},$ (8)
yielding the expression
$\Lambda=-\frac{1}{2}\|U^{\top}B_{i}A_{i}V\|_{\mathrm{Fro}}^{2}-\frac{1}{2}\text{tr}(X^{\top}(I-U^{\top}U))-\frac{1}{2}\text{tr}(Y^{\top}(I-V^{\top}V)).$ (9)
Taking the derivatives gives
$\nabla_{U}\Lambda=$ $-\sum_{i}B_{i}(A_{i}V)(V^{\top}A_{i}^{\top})(B_{i}^{\top}U)+UX$ (10)
$\nabla_{V}\Lambda=$ $-\sum_{i}A_{i}^{\top}(B_{i}^{\top}U)(U^{\top}B_{i})(A_{i}V)+VY$ (11)
Setting these derivatives to zero shows
$\sum_{i}B_{i}(A_{i}V)(V^{\top}A_{i}^{\top})(B_{i}^{\top}U)$ $=UX$ (12)
$\sum_{i}A_{i}^{\top}(B_{i}^{\top}U)(U^{\top}B_{i})(A_{i}V)$ $=VY.$ (13)
Here, one can show that the Lagrange multiplier matrices $X$ and $Y$ are diagonal and nonnegative, since the problem reduces to an eigenvalue problem when either $U$ or $V$ is fixed; this is essentially the argument behind the alternating algorithm in Appendix B. Hence, taking inspiration from classical eigenvalue iteration, we use the following updates to improve our estimates of $U$ and $V$:
$U_{0}^{(k+1)}$ $\leftarrow\sum_{i}B_{i}(A_{i}V^{(k)})((V^{(k)})^{\top}A_{i}^{\top})(B_{i}^{\top}U^{(k)})$ (14)
$V_{0}^{(k+1)}$ $\leftarrow\sum_{i}A_{i}^{\top}(B_{i}^{\top}U^{(k)})((U^{(k)})^{\top}B_{i})(A_{i}V^{(k)})$ (15)
$U^{(k+1)}$ $\leftarrow\texttt{orthogonalize}(U_{0}^{(k+1)})$ (16)
$V^{(k+1)}$ $\leftarrow\texttt{orthogonalize}(V_{0}^{(k+1)})$ (17)
Here, the function orthogonalize orthogonalizes the columns of a matrix, e.g. by using the $Q$ part of the reduced-size $QR$ factorization. Although we lack a formal convergence proof, in practice we find that this method reliably reaches a local optimum of our problem.
By executing matrix operations in the specified sequence, these computations can be rapidly performed on GPUs. Note the expressions above are parenthesized to avoid constructing a large matrix product as an intermediate computation.
Appendix C Proof of Theorem 1
Proof.
For the lower bound, note that by Jensen’s inequality,
$\sum_{i=1}^{n}\|U^{\top}B_{i}A_{i}V\|^{2}_{\mathrm{Fro}}\geq\left\|U^{\top}\sum_{i=1}^{n}B_{i}A_{i}V\right\|_{\mathrm{Fro}}^{2},$
for any $U,V$. Hence,
$\sup_{U,V\in\mathrm{St}(k,d)}\sum_{i=1}^{n}\|U^{\top}B_{i}A_{i}V\|^{2}_{\mathrm{Fro}}\geq\sup_{U,V\in\mathrm{St}(k,d)}\left\|U^{\top}\sum_{i=1}^{n}B_{i}A_{i}V\right\|_{\mathrm{Fro}}^{2}.$ (18)
By the definition of singular value decomposition, the right hand side of (18) is maximized with $U$, $V$ being the top $r$ singular vectors of $\sum_{i=1}^{n}B_{i}A_{i}$, yielding $\left\|U^{\top}\sum_{i=1}^{n}B_{i}A_{i}V\right\|_{\mathrm{Fro}}^{2}=\sum_{i=1}^{r}\bar{\sigma}^{2}_{i}$. Recalling that $\Sigma_{i}=U^{\top}B_{i}A_{i}V$ yields the lower bound.
For the upper bound, recall that $\Sigma_{i}=U^{\top}B_{i}A_{i}V$. Rearranging,
$\mathrm{vec}(\Sigma_{i})=(V^{\top}\otimes U^{\top})\mathrm{vec}(B_{i}A_{i}).$
Define
$\bar{\Sigma}\coloneqq[\mathrm{vec}(\Sigma_{1}),\dots,\mathrm{vec}(\Sigma_{n})].$
By our previous simplification,
$\bar{\Sigma}=(V^{\top}\otimes U^{\top})L.$
Now
$\sum_{i=1}^{n}\|\Sigma_{i}\|^{2}_{\mathrm{Fro}}=\|\bar{\Sigma}\|_{\mathrm{Fro}}^{2}=\mathrm{tr}\left(((V\otimes U)(V\otimes U)^{\top})(LL^{\top})\right)$
Since $U,V$ are orthogonal and size $d\times r$, the top $r^{2}$ eigenvalues of the symmetric matrix $(V\otimes U)(V\otimes U)^{\top}$ will be equal to 1, and the rest will equal 0. The eigenvalues of the symmetric matrix $LL^{\top}$ will be equal to the squared singular values of $L$. We can then apply the Von Neumann trace inequality to obtain the upper bound.
The last statement follows from the Pythagorean theorem and the fact that the $\Sigma_{i}$ is a projection of $B_{i}A_{i}$ to the $U,V$ subspace. ∎
Note that we have only used the fact that the matrix $(V\otimes U)$ has singular values equal to 1; we have not used the fact that it has Kronecker product structure. On the other hand, each vector $\mathrm{vec}(B_{i}A_{i})$ is a sum of $r_{i}$ Kronecker products and cannot be expressed as a Kronecker product. As a result, while the upper bound in the Von Neumann trace inequality is achieved if the eigenvectors of the two matrices align, the Kronecker product structure is a severe constraint and the upper bound we have provided is generous.
Appendix D Training LoRAs
We trained LoRA adapters on 500 natural instruction tasks [27] using Mistral-7B-Instruct-v0.2 [10] as the base model. All LoRA adapters were configured with a rank of 16, i.e., $\forall i,r_{i}=16$. We selected 10 diverse tasks manually for consistent evaluation across experiments and randomly sampled an additional 490 tasks, resulting in a total of 500 tasks. These tasks were exclusively in English (both input and output), ensuring higher quality and thorough review [27]. Each task dataset was divided into training, validation, and test sets (80-10-10). Hyperparameters, such as early stopping, were tuned using the validation sets; that is, we train for five epochs and take the best-performing epoch-checkpoint per validation loss. Evaluation on the test sets demonstrated that LoRA consistently outperformed the base model in terms of both Rouge scores and loss metrics (see Table 1).
Table 2: Main Evaluation Tasks
Task Number Name Type Domain task280 stereoset_classification_stereotype_type classification stereoset task190 snli_classification snli image captions task391 causal_relationship commonsense cause and effect task290 tellmewhy_question_answerability answerability story task1391 winogrande_easy_answer_generation commonsense social and physical task1342 amazon_us_reviews_title title generation amazon reviews task442 com_qa_paraphrase_question_generation question generation wikipedia task620 ohsumed_medical_subject_headings_answer_generation keyword tagging scientific task1598 nyc_long_text_generation data to text restaurants task039 qasc_find_overlapping_words overlap extraction natural science
We use Huggingface [29] in our implementation. For the base model, we use quantization with configuration:
    BitsAndBytesConfig(
        load_in_4bit=True,
        bnb_4bit_use_double_quant=True,
        bnb_4bit_quant_type="nf4",
        bnb_4bit_compute_dtype=torch.bfloat16,
and LoRA configuration:
    LoraConfig(
        r=16,
        lora_alpha=32,
        target_modules=["q_proj", "k_proj", "v_proj"],
        lora_dropout=0.05,
        bias="none",
        task_type="CAUSAL_LM",
        init_lora_weights=init_lora_weights,
Appendix E Avoiding Batched Matrix Multiplication (BMM)
In the envisioned deployment scenario, a service provider hosts a large collection of LoRAs. Upon receiving a request, each user specifies both the input data and the desired LoRA identifier. The provider then processes the base model augmented with the specified LoRA for each user’s data. As a provider is batching a collection of requests for GPU parallelization, they can expect to frequently have more than one unique LoRA identifier per batch.
Traditionally, a specific LoRA is integrated into the base model by transforming $W_{0}\rightarrow W_{0}+B_{i}A_{i}$. Serving multiple LoRAs conventionally would necessitate maintaining and executing a separate copy of the base model for each LoRA, bringing substantial computational overhead. Alternatively, the computation for $W_{0}x$ and $B_{i}A_{i}x$ can be performed independently and subsequently merged. This strategy necessitates only a single instance of $W_{0}x$ computation and storage of LoRA-specific parameters rather than the entire base model.
Consider the batch processing of $\mathbf{B}\mathbf{A}\mathbf{x}$, where boldface indicates that $B_{i},A_{i}$ are stacked into tensors of dimensions $(b\times m\times r)$ and $(b\times r\times n)$ respectively, with batched data $\mathbf{x}$ shaped $(b\times l\times n)$:
$\mathbf{A}\mathbf{x}\leftrightarrow(b\times r\times n)\times(b\times l\times n)\rightarrow(b\times l\times r)\text{\ \ bmm}$
$\mathbf{B(Ax)}\leftrightarrow(b\times m\times r)\times(b\times l\times r)\rightarrow(b\times l\times m)\text{\ \ bmm}.$
Here, “bmm” denotes batched matrix multiplication, a known bottleneck in both throughput and latency. Consider the corresponding operations for our joint compression scheme, $U\mathbf{\Sigma}V^{\top}x$:
$V^{\top}\mathbf{x}\leftrightarrow(\tilde{r}\times n)\times(b\times l\times n)\rightarrow(b\times l\times\tilde{r})\text{\ \ broadcasted}$
$\mathbf{\Sigma}(V^{\top}\mathbf{x})\leftrightarrow(b\times\tilde{r})\times(b\times l\times\tilde{r})\rightarrow(b\times l\times\tilde{r})\text{\ \ broadcasted}$
$U(\mathbf{\Sigma}V^{\top}\mathbf{x})\leftrightarrow(m\times\tilde{r})\times(b\times l\times\tilde{r})\rightarrow(b\times l\times m)\text{\ \ broadcasted}$
In our optimized setup, batched matrix multiplications can be completely circumvented if the $\Sigma_{i}$ matrices are diagonal. If not, given that $\tilde{r}\ll m,n$, any required batched matrix multiplication remains computationally inexpensive.
Appendix F Additional Results
This section elaborates on the results that underpin the figures presented in the main text and showcases a consistent correlation across various evaluation metrics. Additionally, we assess the significance of achieving convergence and the performance of compression on new unseen LoRA models.
F.1 Relative Rouge-L Performance and Compression Rate
Table 3 presents comprehensive results from the experiments underlying Figure 2(b) for each evaluation task. Additionally, we incorporate results using the Ties-merging benchmark [32], which consolidates all LoRA-adapters into a single adapter of identical configuration and parameter count; this integration significantly compromises performance.
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 0.26 $\pm$ 0.00 0.02 $\pm$ 0.00 0.19 $\pm$ 0.00 0.42 $\pm$ 0.00 0.11 $\pm$ 0.00 0.47 $\pm$ 0.00 0.11 $\pm$ 0.00 0.23 $\pm$ 0.00 0.19 $\pm$ 0.00 0.77 $\pm$ 0.00 0.28 $\pm$ 0.21 1.00/1.00 lora 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00/0.00 TIES 10 0.81 $\pm$ 0.00 0.57 $\pm$ 0.02 0.45 $\pm$ 0.04 0.10 $\pm$ 0.01 0.83 $\pm$ 0.01 0.47 $\pm$ 0.00 0.69 $\pm$ 0.01 0.57 $\pm$ 0.00 0.82 $\pm$ 0.01 0.85 $\pm$ 0.00 0.62 $\pm$ 0.23 1.00 / 1.00 50 0.59 $\pm$ 0.00 0.41 $\pm$ 0.00 0.18 $\pm$ 0.05 0.03 $\pm$ 0.01 0.91 $\pm$ 0.01 0.31 $\pm$ 0.00 0.65 $\pm$ 0.00 0.62 $\pm$ 0.00 0.32 $\pm$ 0.04 0.84 $\pm$ 0.00 0.48 $\pm$ 0.28 1.00 / 1.00 100 0.55 $\pm$ 0.00 0.40 $\pm$ 0.00 0.20 $\pm$ 0.05 0.01 $\pm$ 0.02 0.88 $\pm$ 0.00 0.33 $\pm$ 0.00 0.64 $\pm$ 0.00 0.57 $\pm$ 0.02 0.01 $\pm$ 0.00 0.82 $\pm$ 0.00 0.44 $\pm$ 0.30 1.00 / 1.00 500 0.37 $\pm$ 0.00 0.26 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.83 $\pm$ 0.00 0.29 $\pm$ 0.00 0.57 $\pm$ 0.00 0.37 $\pm$ 0.00 0.01 $\pm$ 0.00 0.43 $\pm$ 0.00 0.31 $\pm$ 0.26 1.00 / 1.00 SVD SVD 2 0.98 $\pm$ 0.03 1.07 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01 1.00 $\pm$ 0.01 1.00 $\pm$ 0.10 1.00 $\pm$ 0.01 1.00 $\pm$ 0.01 1.00 $\pm$ 0.04 0.88 / 0.88 SVD 4 0.99 $\pm$ 0.04 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.02 0.99 $\pm$ 0.08 0.99 $\pm$ 0.01 1.00 $\pm$ 0.01 1.00 $\pm$ 0.03 0.75 / 0.75 SVD 8 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 1.01 $\pm$ 0.01 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.50 / 0.50 SVD 16 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 / 0.00 10 diagonal (D) 16 D 1.02 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.99 $\pm$ 0.00 0.96 $\pm$ 0.00 1.02 $\pm$ 0.02 1.13 $\pm$ 0.03 0.99 $\pm$ 0.02 0.98 $\pm$ 0.01 1.01 $\pm$ 0.05 1.00 / 0.90 32 D 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.01 $\pm$ 0.01 0.99 $\pm$ 0.00 0.97 $\pm$ 0.01 1.05 $\pm$ 0.03 1.00 $\pm$ 0.01 1.00 $\pm$ 0.01 1.00 $\pm$ 0.03 1.00 / 0.80 64 D 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.99 $\pm$ 0.01 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 / 0.60 128 D 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 / 0.20 256 D 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 / -0.60 10 full (F) 16 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.00 1.01 $\pm$ 0.02 1.07 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.03 1.00/0.90 32 F 1.02 $\pm$ 0.01 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.96 $\pm$ 0.01 1.00 $\pm$ 0.02 1.00 $\pm$ 0.01 1.01 $\pm$ 0.00 1.00 $\pm$ 0.02 0.99/0.79 64 F 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.98 $\pm$ 0.01 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.97/0.57 128 F 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.88/0.07 256 F 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.50/-1.10 50 diagonal (D) 16 D 0.98 $\pm$ 0.04 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.92 $\pm$ 0.06 0.84 $\pm$ 0.07 0.92 $\pm$ 0.02 0.68 $\pm$ 0.05 0.87 $\pm$ 0.10 0.88 $\pm$ 0.07 0.83 $\pm$ 0.02 0.89 $\pm$ 0.10 1.00 / 0.98 32 D 1.00 $\pm$ 0.02 1.02 $\pm$ 0.02 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.84 $\pm$ 0.02 1.00 $\pm$ 0.13 0.98 $\pm$ 0.01 0.88 $\pm$ 0.01 0.96 $\pm$ 0.07 1.00 / 0.96 64 D 1.02 $\pm$ 0.00 1.05 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.97 $\pm$ 0.00 0.99 $\pm$ 0.01 1.09 $\pm$ 0.03 1.01 $\pm$ 0.01 0.90 $\pm$ 0.01 1.00 $\pm$ 0.05 1.00 / 0.92 128 D 1.01 $\pm$ 0.01 1.08 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.00 0.98 $\pm$ 0.01 1.11 $\pm$ 0.03 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.01 $\pm$ 0.04 1.00 / 0.84 256 D 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 0.97 $\pm$ 0.03 1.01 $\pm$ 0.03 1.00 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.02 1.00 / 0.68 50 full (F) 16 F 0.99 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.94 $\pm$ 0.01 0.64 $\pm$ 0.10 1.01 $\pm$ 0.15 0.97 $\pm$ 0.02 0.87 $\pm$ 0.00 0.93 $\pm$ 0.12 1.00/0.98 32 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01 0.96 $\pm$ 0.00 0.95 $\pm$ 0.01 1.09 $\pm$ 0.02 1.01 $\pm$ 0.02 0.89 $\pm$ 0.01 0.99 $\pm$ 0.05 0.99/0.95 64 F 1.02 $\pm$ 0.01 1.06 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.01 1.03 $\pm$ 0.01 1.11 $\pm$ 0.00 1.00 $\pm$ 0.01 0.98 $\pm$ 0.02 1.02 $\pm$ 0.04 0.97/0.89 128 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.98 $\pm$ 0.00 0.98 $\pm$ 0.01 1.03 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.03 0.88/0.72 256 F 1.00 $\pm$ 0.00 1.02 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.50/0.18 100 diagonal (D) 16 D 0.80 $\pm$ 0.07 0.89 $\pm$ 0.06 0.93 $\pm$ 0.03 0.96 $\pm$ 0.01 0.50 $\pm$ 0.09 0.78 $\pm$ 0.01 0.28 $\pm$ 0.07 0.52 $\pm$ 0.10 0.78 $\pm$ 0.03 0.81 $\pm$ 0.02 0.72 $\pm$ 0.22 1.00 / 0.99 32 D 0.95 $\pm$ 0.06 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.91 $\pm$ 0.06 0.80 $\pm$ 0.14 0.89 $\pm$ 0.06 0.60 $\pm$ 0.10 0.77 $\pm$ 0.26 0.91 $\pm$ 0.02 0.83 $\pm$ 0.02 0.86 $\pm$ 0.14 1.00 / 0.98 64 D 1.01 $\pm$ 0.03 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.02 0.96 $\pm$ 0.01 0.94 $\pm$ 0.01 0.88 $\pm$ 0.05 1.11 $\pm$ 0.08 0.96 $\pm$ 0.02 0.87 $\pm$ 0.03 0.97 $\pm$ 0.07 1.00 / 0.96 128 D 1.01 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.97 $\pm$ 0.00 1.00 $\pm$ 0.03 1.11 $\pm$ 0.02 0.99 $\pm$ 0.01 0.89 $\pm$ 0.02 1.00 $\pm$ 0.05 1.00 / 0.92 256 D 1.00 $\pm$ 0.00 1.06 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00 1.00 $\pm$ 0.01 1.11 $\pm$ 0.03 1.00 $\pm$ 0.01 0.98 $\pm$ 0.01 1.01 $\pm$ 0.04 1.00 / 0.84 100 full (F) 16 F 0.95 $\pm$ 0.01 0.97 $\pm$ 0.03 0.97 $\pm$ 0.03 0.97 $\pm$ 0.03 0.93 $\pm$ 0.01 0.92 $\pm$ 0.01 0.64 $\pm$ 0.03 0.89 $\pm$ 0.16 0.87 $\pm$ 0.02 0.83 $\pm$ 0.01 0.89 $\pm$ 0.11 1.00/0.99 32 F 1.00 $\pm$ 0.02 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.97 $\pm$ 0.01 0.95 $\pm$ 0.00 0.86 $\pm$ 0.03 1.12 $\pm$ 0.03 0.96 $\pm$ 0.01 0.87 $\pm$ 0.00 0.97 $\pm$ 0.07 0.99/0.97 64 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.99 $\pm$ 0.01 1.09 $\pm$ 0.01 0.99 $\pm$ 0.02 0.89 $\pm$ 0.00 0.99 $\pm$ 0.05 0.97/0.93 128 F 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 1.03 $\pm$ 0.01 1.10 $\pm$ 0.01 1.01 $\pm$ 0.00 0.99 $\pm$ 0.01 1.02 $\pm$ 0.04 0.88/0.80 256 F 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00 1.00 $\pm$ 0.03 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.50/0.34 500 diagonal (D) 16 D 0.57 $\pm$ 0.07 0.55 $\pm$ 0.03 0.83 $\pm$ 0.04 0.78 $\pm$ 0.16 0.85 $\pm$ 0.04 0.68 $\pm$ 0.07 0.24 $\pm$ 0.01 0.43 $\pm$ 0.01 0.76 $\pm$ 0.06 0.79 $\pm$ 0.01 0.65 $\pm$ 0.20 1.00 / 1.00 32 D 0.61 $\pm$ 0.12 0.55 $\pm$ 0.08 0.83 $\pm$ 0.02 0.84 $\pm$ 0.12 0.91 $\pm$ 0.02 0.71 $\pm$ 0.05 0.29 $\pm$ 0.05 0.47 $\pm$ 0.08 0.79 $\pm$ 0.04 0.79 $\pm$ 0.01 0.68 $\pm$ 0.20 1.00 / 1.00 64 D 0.73 $\pm$ 0.02 0.63 $\pm$ 0.11 0.89 $\pm$ 0.04 0.97 $\pm$ 0.00 0.94 $\pm$ 0.00 0.83 $\pm$ 0.05 0.45 $\pm$ 0.09 0.50 $\pm$ 0.07 0.82 $\pm$ 0.02 0.80 $\pm$ 0.02 0.76 $\pm$ 0.18 1.00 / 0.99 128 D 0.84 $\pm$ 0.00 0.92 $\pm$ 0.02 0.97 $\pm$ 0.03 0.98 $\pm$ 0.01 0.94 $\pm$ 0.00 0.88 $\pm$ 0.02 0.60 $\pm$ 0.15 0.53 $\pm$ 0.01 0.85 $\pm$ 0.05 0.80 $\pm$ 0.02 0.83 $\pm$ 0.15 1.00 / 0.98 256 D 0.99 $\pm$ 0.03 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.96 $\pm$ 0.00 0.92 $\pm$ 0.03 0.66 $\pm$ 0.06 0.84 $\pm$ 0.14 0.92 $\pm$ 0.02 0.84 $\pm$ 0.01 0.91 $\pm$ 0.11 1.00 / 0.97 500 full (F) 16 F 0.57 $\pm$ 0.01 0.43 $\pm$ 0.07 0.78 $\pm$ 0.01 0.97 $\pm$ 0.00 0.96 $\pm$ 0.00 0.83 $\pm$ 0.01 0.64 $\pm$ 0.00 0.53 $\pm$ 0.03 0.83 $\pm$ 0.01 0.83 $\pm$ 0.00 0.75 $\pm$ 0.17 1.00/1.00 32 F 0.79 $\pm$ 0.05 0.54 $\pm$ 0.04 0.93 $\pm$ 0.02 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.90 $\pm$ 0.01 0.69 $\pm$ 0.01 0.50 $\pm$ 0.00 0.86 $\pm$ 0.02 0.83 $\pm$ 0.01 0.81 $\pm$ 0.16 0.99/0.99 64 F 1.02 $\pm$ 0.00 0.96 $\pm$ 0.01 0.94 $\pm$ 0.01 1.00 $\pm$ 0.01 0.96 $\pm$ 0.00 0.97 $\pm$ 0.01 0.73 $\pm$ 0.01 0.54 $\pm$ 0.01 0.91 $\pm$ 0.01 0.86 $\pm$ 0.00 0.89 $\pm$ 0.14 0.97/0.96 128 F 1.03 $\pm$ 0.01 0.97 $\pm$ 0.02 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.87 $\pm$ 0.01 1.07 $\pm$ 0.02 0.98 $\pm$ 0.00 0.87 $\pm$ 0.00 0.97 $\pm$ 0.06 0.88/0.86 256 F 1.03 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.97 $\pm$ 0.01 0.99 $\pm$ 0.02 1.03 $\pm$ 0.01 1.00 $\pm$ 0.01 0.87 $\pm$ 0.00 0.99 $\pm$ 0.05 0.50/0.47
Table 3: Relative In-Distribution ROUGE-L scores for various tasks and methods
F.2 Absolute Rouge-L Performance and Compression Rate
Table 4 provides the full results behind Table 3, but with Rouge-L scores instead of relative performance compared to LoRA.
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 24.44 $\pm$ 0.00 1.60 $\pm$ 0.00 19.13 $\pm$ 0.00 39.22 $\pm$ 0.00 10.27 $\pm$ 0.00 35.46 $\pm$ 0.00 7.85 $\pm$ 0.00 6.22 $\pm$ 0.00 17.82 $\pm$ 0.00 38.87 $\pm$ 0.00 20.24 $\pm$ 13.27 1.00/1.00 lora 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 74.88 $\pm$ 0.00 74.40 $\pm$ 0.00 26.68 $\pm$ 0.00 95.00 $\pm$ 0.00 50.32 $\pm$ 0.00 78.87 $\pm$ 22.56 0.00/0.00 TIES 10 76.50 $\pm$ 0.00 49.00 $\pm$ 1.73 44.33 $\pm$ 4.04 9.80 $\pm$ 0.58 78.56 $\pm$ 0.96 35.24 $\pm$ 0.00 51.37 $\pm$ 0.67 15.26 $\pm$ 0.12 77.67 $\pm$ 1.15 42.72 $\pm$ 0.01 48.05 $\pm$ 23.61 1.00 / 1.00 50 55.80 $\pm$ 0.00 35.00 $\pm$ 0.00 18.00 $\pm$ 5.20 2.42 $\pm$ 0.50 85.78 $\pm$ 0.96 23.03 $\pm$ 0.00 48.03 $\pm$ 0.00 16.50 $\pm$ 0.00 30.00 $\pm$ 3.46 42.47 $\pm$ 0.02 35.70 $\pm$ 23.01 1.00 / 1.00 100 52.43 $\pm$ 0.00 34.00 $\pm$ 0.00 19.67 $\pm$ 4.62 1.09 $\pm$ 1.66 83.33 $\pm$ 0.00 24.89 $\pm$ 0.00 47.52 $\pm$ 0.00 15.18 $\pm$ 0.42 1.00 $\pm$ 0.00 41.19 $\pm$ 0.03 32.03 $\pm$ 24.50 1.00 / 1.00 500 35.18 $\pm$ 0.00 22.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 78.00 $\pm$ 0.00 21.46 $\pm$ 0.00 42.22 $\pm$ 0.04 9.93 $\pm$ 0.13 1.00 $\pm$ 0.00 21.50 $\pm$ 0.03 23.27 $\pm$ 23.64 1.00 / 1.00 SVD SVD 2 93.15 $\pm$ 2.77 92.24 $\pm$ 1.85 99.09 $\pm$ 0.18 93.44 $\pm$ 0.14 93.89 $\pm$ 0.35 73.74 $\pm$ 0.51 74.55 $\pm$ 0.98 26.80 $\pm$ 2.79 95.06 $\pm$ 1.35 50.21 $\pm$ 0.44 79.11 $\pm$ 22.72 0.88 / 0.88 SVD 4 94.01 $\pm$ 3.60 89.21 $\pm$ 0.71 99.05 $\pm$ 0.09 93.65 $\pm$ 0.03 94.66 $\pm$ 0.63 74.89 $\pm$ 0.33 73.61 $\pm$ 1.15 26.34 $\pm$ 2.13 93.98 $\pm$ 0.77 50.47 $\pm$ 0.54 78.90 $\pm$ 22.68 0.75 / 0.75 SVD 8 95.00 $\pm$ 0.00 87.40 $\pm$ 0.59 99.05 $\pm$ 0.09 93.65 $\pm$ 0.03 94.36 $\pm$ 0.38 74.58 $\pm$ 0.12 75.07 $\pm$ 0.00 26.71 $\pm$ 0.27 95.51 $\pm$ 1.09 50.89 $\pm$ 0.07 81.01 $\pm$ 21.74 0.50 / 0.50 SVD 16 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 74.90 $\pm$ 0.03 74.23 $\pm$ 0.18 26.68 $\pm$ 0.00 95.00 $\pm$ 0.00 50.30 $\pm$ 0.02 78.36 $\pm$ 22.97 0.00 / 0.00 10 diagonal (D) 16 D 96.67 $\pm$ 0.58 87.00 $\pm$ 1.00 99.00 $\pm$ 0.00 94.00 $\pm$ 0.67 93.11 $\pm$ 0.38 72.08 $\pm$ 0.06 76.26 $\pm$ 1.19 30.11 $\pm$ 0.79 94.00 $\pm$ 1.73 49.30 $\pm$ 0.46 79.15 $\pm$ 22.18 1.00 / 0.90 32 D 95.67 $\pm$ 0.58 90.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.00 $\pm$ 0.33 94.89 $\pm$ 0.51 73.86 $\pm$ 0.31 71.92 $\pm$ 0.84 27.89 $\pm$ 0.70 94.67 $\pm$ 0.58 50.36 $\pm$ 0.26 79.13 $\pm$ 22.75 1.00 / 0.80 64 D 95.00 $\pm$ 0.00 88.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.78 $\pm$ 0.38 74.61 $\pm$ 0.13 74.97 $\pm$ 0.58 26.35 $\pm$ 0.25 96.00 $\pm$ 0.00 50.99 $\pm$ 0.06 79.37 $\pm$ 22.94 1.00 / 0.60 128 D 95.00 $\pm$ 0.00 86.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 74.92 $\pm$ 0.13 74.96 $\pm$ 0.51 26.45 $\pm$ 0.23 95.00 $\pm$ 0.00 50.21 $\pm$ 0.12 79.02 $\pm$ 22.84 1.00 / 0.20 256 D 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 74.88 $\pm$ 0.00 74.40 $\pm$ 0.00 26.68 $\pm$ 0.00 95.00 $\pm$ 0.00 50.27 $\pm$ 0.02 78.92 $\pm$ 22.77 1.00 / -0.60 10 full (F) 16 F 97.00 $\pm$ 0.00 91.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.56 $\pm$ 0.69 73.60 $\pm$ 0.36 74.94 $\pm$ 1.25 28.66 $\pm$ 0.03 96.00 $\pm$ 1.00 50.15 $\pm$ 0.20 79.75 $\pm$ 22.72 1.00/0.90 32 F 96.67 $\pm$ 0.58 89.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.22 $\pm$ 0.19 94.44 $\pm$ 0.19 74.11 $\pm$ 0.19 71.74 $\pm$ 0.59 26.74 $\pm$ 0.50 94.67 $\pm$ 0.58 50.63 $\pm$ 0.24 79.06 $\pm$ 23.01 0.99/0.79 64 F 95.00 $\pm$ 0.00 88.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.56 $\pm$ 0.38 74.56 $\pm$ 0.13 75.47 $\pm$ 0.58 26.26 $\pm$ 0.34 96.00 $\pm$ 0.00 50.89 $\pm$ 0.17 79.41 $\pm$ 22.97 0.97/0.57 128 F 95.00 $\pm$ 0.00 86.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 75.04 $\pm$ 0.03 74.40 $\pm$ 0.00 26.53 $\pm$ 0.13 95.00 $\pm$ 0.00 50.36 $\pm$ 0.03 79.00 $\pm$ 22.81 0.88/0.07 256 F 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 74.90 $\pm$ 0.03 74.29 $\pm$ 0.19 26.68 $\pm$ 0.00 95.00 $\pm$ 0.00 50.30 $\pm$ 0.03 78.92 $\pm$ 22.77 0.50/-1.10 50 diagonal (D) 16 D 92.76 $\pm$ 3.53 84.67 $\pm$ 1.15 99.00 $\pm$ 0.00 86.17 $\pm$ 5.81 79.68 $\pm$ 6.21 69.07 $\pm$ 1.54 50.65 $\pm$ 3.97 23.27 $\pm$ 2.60 83.90 $\pm$ 6.43 41.86 $\pm$ 0.96 71.10 $\pm$ 23.99 1.00 / 0.98 32 D 95.33 $\pm$ 2.08 87.33 $\pm$ 2.08 99.00 $\pm$ 0.00 92.60 $\pm$ 0.29 90.32 $\pm$ 1.04 71.16 $\pm$ 1.47 62.51 $\pm$ 1.64 26.60 $\pm$ 3.54 93.33 $\pm$ 1.15 44.35 $\pm$ 0.41 76.25 $\pm$ 23.81 1.00 / 0.96 64 D 97.00 $\pm$ 0.00 90.33 $\pm$ 1.53 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 93.00 $\pm$ 0.58 72.37 $\pm$ 0.35 73.39 $\pm$ 0.93 29.06 $\pm$ 0.80 95.67 $\pm$ 0.58 45.43 $\pm$ 0.34 78.90 $\pm$ 23.29 1.00 / 0.92 128 D 96.33 $\pm$ 0.58 92.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.00 $\pm$ 0.58 73.32 $\pm$ 0.24 73.03 $\pm$ 1.09 29.51 $\pm$ 0.93 95.00 $\pm$ 0.00 50.16 $\pm$ 0.74 79.56 $\pm$ 22.51 1.00 / 0.84 256 D 95.67 $\pm$ 0.58 88.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 94.67 $\pm$ 0.67 74.82 $\pm$ 0.24 72.36 $\pm$ 2.07 26.90 $\pm$ 0.75 95.33 $\pm$ 0.58 50.73 $\pm$ 0.46 79.14 $\pm$ 22.90 1.00 / 0.68 50 full (F) 16 F 94.06 $\pm$ 3.54 85.67 $\pm$ 1.15 98.67 $\pm$ 0.58 90.35 $\pm$ 1.37 89.90 $\pm$ 1.91 70.32 $\pm$ 0.66 47.62 $\pm$ 7.28 26.88 $\pm$ 3.96 92.33 $\pm$ 1.53 43.68 $\pm$ 0.24 73.95 $\pm$ 24.73 1.00/0.98 32 F 97.00 $\pm$ 0.00 85.67 $\pm$ 1.53 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 92.22 $\pm$ 0.69 71.88 $\pm$ 0.30 71.01 $\pm$ 1.02 29.07 $\pm$ 0.65 95.67 $\pm$ 1.53 44.97 $\pm$ 0.41 78.02 $\pm$ 23.18 0.99/0.95 64 F 96.67 $\pm$ 0.58 91.00 $\pm$ 2.00 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.22 $\pm$ 0.51 73.16 $\pm$ 0.41 76.28 $\pm$ 0.51 29.67 $\pm$ 0.12 95.33 $\pm$ 0.58 49.31 $\pm$ 1.00 79.72 $\pm$ 22.50 0.97/0.89 128 F 97.00 $\pm$ 0.00 91.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.33 $\pm$ 0.00 94.11 $\pm$ 0.51 73.51 $\pm$ 0.23 73.17 $\pm$ 0.58 27.53 $\pm$ 1.12 95.00 $\pm$ 1.00 50.56 $\pm$ 0.06 79.42 $\pm$ 22.93 0.88/0.72 256 F 95.00 $\pm$ 0.00 88.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.44 $\pm$ 0.19 74.25 $\pm$ 0.21 74.97 $\pm$ 0.58 26.79 $\pm$ 0.09 96.00 $\pm$ 0.00 50.86 $\pm$ 0.19 79.30 $\pm$ 22.82 0.50/0.18 100 diagonal (D) 16 D 76.43 $\pm$ 7.07 76.67 $\pm$ 4.93 91.61 $\pm$ 2.75 89.99 $\pm$ 1.07 47.55 $\pm$ 8.56 58.08 $\pm$ 0.72 20.77 $\pm$ 5.50 13.90 $\pm$ 2.79 73.93 $\pm$ 3.13 40.74 $\pm$ 0.85 58.97 $\pm$ 26.83 1.00 / 0.99 32 D 90.10 $\pm$ 5.85 84.00 $\pm$ 1.00 99.00 $\pm$ 0.00 85.52 $\pm$ 5.34 75.69 $\pm$ 12.75 66.62 $\pm$ 4.18 44.66 $\pm$ 7.26 20.49 $\pm$ 7.07 86.67 $\pm$ 1.86 42.01 $\pm$ 0.94 69.48 $\pm$ 25.14 1.00 / 0.98 64 D 95.56 $\pm$ 2.49 86.67 $\pm$ 0.58 99.00 $\pm$ 0.00 92.24 $\pm$ 1.68 90.89 $\pm$ 1.17 70.35 $\pm$ 0.45 65.62 $\pm$ 4.03 29.58 $\pm$ 2.02 91.67 $\pm$ 2.31 43.64 $\pm$ 1.36 76.52 $\pm$ 23.02 1.00 / 0.96 128 D 96.00 $\pm$ 0.00 87.33 $\pm$ 1.15 99.00 $\pm$ 0.00 93.89 $\pm$ 0.19 93.00 $\pm$ 0.58 72.70 $\pm$ 0.30 74.34 $\pm$ 2.07 29.66 $\pm$ 0.54 93.67 $\pm$ 0.58 44.82 $\pm$ 0.89 78.44 $\pm$ 22.87 1.00 / 0.92 256 D 95.00 $\pm$ 0.00 91.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.11 $\pm$ 0.19 73.05 $\pm$ 0.20 74.52 $\pm$ 0.95 29.67 $\pm$ 0.67 95.33 $\pm$ 0.58 49.42 $\pm$ 0.65 79.37 $\pm$ 22.38 1.00 / 0.84 100 full (F) 16 F 90.70 $\pm$ 1.07 83.00 $\pm$ 2.65 96.00 $\pm$ 3.00 91.22 $\pm$ 2.94 87.94 $\pm$ 0.54 68.72 $\pm$ 1.05 47.57 $\pm$ 2.54 23.75 $\pm$ 4.33 82.33 $\pm$ 2.08 41.51 $\pm$ 0.67 71.27 $\pm$ 24.23 1.00/0.99 32 F 95.33 $\pm$ 1.53 85.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.50 $\pm$ 0.22 91.44 $\pm$ 0.84 70.94 $\pm$ 0.02 63.64 $\pm$ 1.98 29.82 $\pm$ 0.81 91.67 $\pm$ 0.58 43.94 $\pm$ 0.18 76.43 $\pm$ 23.01 0.99/0.97 64 F 97.00 $\pm$ 0.00 85.67 $\pm$ 1.53 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 92.56 $\pm$ 0.19 72.11 $\pm$ 0.08 73.29 $\pm$ 0.64 29.15 $\pm$ 0.24 94.33 $\pm$ 1.53 44.97 $\pm$ 0.05 78.18 $\pm$ 23.03 0.97/0.93 128 F 96.33 $\pm$ 0.58 90.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.00 $\pm$ 0.00 93.89 $\pm$ 0.19 73.11 $\pm$ 0.36 76.50 $\pm$ 1.01 29.45 $\pm$ 0.35 96.00 $\pm$ 0.00 49.81 $\pm$ 0.34 79.74 $\pm$ 22.47 0.88/0.80 256 F 96.33 $\pm$ 0.58 88.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.89 $\pm$ 0.19 74.40 $\pm$ 0.16 72.90 $\pm$ 0.12 26.77 $\pm$ 0.68 96.00 $\pm$ 0.00 50.83 $\pm$ 0.09 79.35 $\pm$ 23.04 0.50/0.34 500 diagonal (D) 16 D 54.44 $\pm$ 6.87 47.00 $\pm$ 2.83 82.21 $\pm$ 3.59 73.38 $\pm$ 14.97 80.08 $\pm$ 3.71 51.02 $\pm$ 5.31 17.49 $\pm$ 1.10 11.58 $\pm$ 0.21 72.67 $\pm$ 6.03 39.65 $\pm$ 0.28 53.16 $\pm$ 24.97 1.00 / 1.00 32 D 58.08 $\pm$ 11.52 47.00 $\pm$ 7.07 82.06 $\pm$ 1.69 78.62 $\pm$ 11.23 85.57 $\pm$ 1.48 52.98 $\pm$ 3.81 21.73 $\pm$ 3.95 12.53 $\pm$ 2.26 75.33 $\pm$ 4.04 39.78 $\pm$ 0.42 55.66 $\pm$ 25.48 1.00 / 1.00 64 D 69.21 $\pm$ 2.03 54.50 $\pm$ 9.19 88.33 $\pm$ 4.04 91.11 $\pm$ 0.38 88.78 $\pm$ 0.38 62.36 $\pm$ 3.52 33.36 $\pm$ 6.69 13.34 $\pm$ 1.86 77.67 $\pm$ 2.31 40.42 $\pm$ 0.98 62.16 $\pm$ 26.05 1.00 / 0.99 128 D 79.77 $\pm$ 0.37 79.50 $\pm$ 2.12 95.89 $\pm$ 2.83 91.89 $\pm$ 1.39 88.67 $\pm$ 0.00 65.92 $\pm$ 1.79 44.98 $\pm$ 10.98 14.14 $\pm$ 0.19 81.00 $\pm$ 5.00 40.34 $\pm$ 0.80 67.82 $\pm$ 26.35 1.00 / 0.98 256 D 93.83 $\pm$ 2.52 85.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 90.56 $\pm$ 0.38 68.95 $\pm$ 1.92 49.39 $\pm$ 4.36 22.33 $\pm$ 3.78 87.33 $\pm$ 2.31 42.15 $\pm$ 0.73 72.83 $\pm$ 25.93 1.00 / 0.97 500 full (F) 16 F 54.30 $\pm$ 1.13 37.00 $\pm$ 5.66 77.67 $\pm$ 0.58 91.00 $\pm$ 0.00 90.56 $\pm$ 0.19 62.47 $\pm$ 0.79 47.56 $\pm$ 0.29 14.18 $\pm$ 0.67 79.00 $\pm$ 1.00 41.58 $\pm$ 0.23 60.31 $\pm$ 24.42 1.00/1.00 32 F 75.10 $\pm$ 4.92 46.50 $\pm$ 3.54 91.67 $\pm$ 1.53 91.56 $\pm$ 0.19 91.56 $\pm$ 0.38 67.37 $\pm$ 0.83 51.17 $\pm$ 0.81 13.44 $\pm$ 0.02 81.67 $\pm$ 1.53 41.92 $\pm$ 0.42 65.84 $\pm$ 25.64 0.99/0.99 64 F 96.94 $\pm$ 0.42 82.50 $\pm$ 0.71 93.33 $\pm$ 0.58 93.89 $\pm$ 0.69 90.67 $\pm$ 0.00 72.30 $\pm$ 0.71 54.63 $\pm$ 0.79 14.49 $\pm$ 0.27 86.33 $\pm$ 0.58 43.16 $\pm$ 0.08 72.49 $\pm$ 26.64 0.97/0.96 128 F 97.67 $\pm$ 0.58 83.50 $\pm$ 2.12 98.00 $\pm$ 0.00 93.56 $\pm$ 0.19 92.00 $\pm$ 0.00 71.92 $\pm$ 0.19 65.02 $\pm$ 0.81 28.49 $\pm$ 0.55 93.00 $\pm$ 0.00 43.85 $\pm$ 0.12 76.47 $\pm$ 23.77 0.88/0.86 256 F 98.00 $\pm$ 0.00 88.50 $\pm$ 0.71 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 93.00 $\pm$ 0.88 72.45 $\pm$ 0.38 73.77 $\pm$ 1.21 27.59 $\pm$ 0.39 95.33 $\pm$ 0.58 43.81 $\pm$ 0.17 78.18 $\pm$ 24.16 0.50/0.47
Table 4: Absolute In-Distribution ROUGE-L scores for various tasks and methods
F.3 Relative Rouge-1 Performance and Compression Rate
Table 5 provides full results for relative performance of Rouge-1, which shows the same trends as the results for relative performance of Rouge-L (Table 3).
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 0.26 $\pm$ 0.00 0.02 $\pm$ 0.00 0.19 $\pm$ 0.00 0.42 $\pm$ 0.00 0.11 $\pm$ 0.00 0.51 $\pm$ 0.00 0.11 $\pm$ 0.00 0.26 $\pm$ 0.00 0.19 $\pm$ 0.00 0.80 $\pm$ 0.00 0.29 $\pm$ 0.22 1.00/1.00 lora 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00/0.00 TIES 10 0.81 $\pm$ 0.00 0.57 $\pm$ 0.02 0.45 $\pm$ 0.04 0.10 $\pm$ 0.01 0.83 $\pm$ 0.01 0.52 $\pm$ 0.00 0.71 $\pm$ 0.01 0.58 $\pm$ 0.00 0.82 $\pm$ 0.01 0.80 $\pm$ 0.00 0.62 $\pm$ 0.22 1.00 / 1.00 50 0.59 $\pm$ 0.00 0.41 $\pm$ 0.00 0.18 $\pm$ 0.05 0.03 $\pm$ 0.01 0.91 $\pm$ 0.01 0.34 $\pm$ 0.00 0.67 $\pm$ 0.00 0.62 $\pm$ 0.00 0.32 $\pm$ 0.04 0.78 $\pm$ 0.00 0.48 $\pm$ 0.27 1.00 / 1.00 100 0.55 $\pm$ 0.00 0.40 $\pm$ 0.00 0.20 $\pm$ 0.05 0.01 $\pm$ 0.02 0.88 $\pm$ 0.00 0.36 $\pm$ 0.00 0.65 $\pm$ 0.00 0.57 $\pm$ 0.02 0.01 $\pm$ 0.00 0.78 $\pm$ 0.00 0.44 $\pm$ 0.29 1.00 / 1.00 500 0.37 $\pm$ 0.00 0.26 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.83 $\pm$ 0.00 0.31 $\pm$ 0.00 0.58 $\pm$ 0.00 0.37 $\pm$ 0.00 0.01 $\pm$ 0.00 0.41 $\pm$ 0.00 0.32 $\pm$ 0.26 1.00 / 1.00 SVD SVD 2 0.98 $\pm$ 0.03 1.07 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.10 1.00 $\pm$ 0.01 0.99 $\pm$ 0.01 1.00 $\pm$ 0.04 0.88 / 0.88 SVD 4 0.99 $\pm$ 0.04 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.99 $\pm$ 0.08 0.99 $\pm$ 0.01 1.01 $\pm$ 0.00 1.00 $\pm$ 0.03 0.75 / 0.75 SVD 8 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 1.01 $\pm$ 0.01 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.50 / 0.50 SVD 16 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 / 0.00 10 diagonal (D) 16 D 1.02 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.99 $\pm$ 0.00 0.97 $\pm$ 0.00 1.03 $\pm$ 0.02 1.12 $\pm$ 0.03 0.99 $\pm$ 0.02 0.99 $\pm$ 0.00 1.01 $\pm$ 0.04 1.00 / 0.90 32 D 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.01 $\pm$ 0.01 0.99 $\pm$ 0.00 0.97 $\pm$ 0.01 1.04 $\pm$ 0.03 1.00 $\pm$ 0.01 1.01 $\pm$ 0.01 1.01 $\pm$ 0.02 1.00 / 0.80 64 D 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.99 $\pm$ 0.01 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 / 0.60 128 D 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 / 0.20 256 D 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 / -0.60 10 full (F) 16 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.99 $\pm$ 0.00 1.01 $\pm$ 0.02 1.07 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.02 $\pm$ 0.03 1.00/0.90 32 F 1.02 $\pm$ 0.01 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.96 $\pm$ 0.01 1.00 $\pm$ 0.02 1.00 $\pm$ 0.01 1.01 $\pm$ 0.00 1.00 $\pm$ 0.02 0.99/0.79 64 F 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.98 $\pm$ 0.01 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.97/0.57 128 F 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.88/0.07 256 F 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.50/-1.10 50 diagonal (D) 16 D 0.98 $\pm$ 0.04 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.92 $\pm$ 0.06 0.85 $\pm$ 0.06 0.94 $\pm$ 0.02 0.69 $\pm$ 0.05 0.88 $\pm$ 0.10 0.88 $\pm$ 0.07 0.86 $\pm$ 0.01 0.90 $\pm$ 0.10 1.00 / 0.98 32 D 1.00 $\pm$ 0.02 1.02 $\pm$ 0.02 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.85 $\pm$ 0.02 1.00 $\pm$ 0.12 0.98 $\pm$ 0.01 0.90 $\pm$ 0.00 0.97 $\pm$ 0.06 1.00 / 0.96 64 D 1.02 $\pm$ 0.00 1.05 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.97 $\pm$ 0.01 0.99 $\pm$ 0.01 1.09 $\pm$ 0.03 1.01 $\pm$ 0.01 0.94 $\pm$ 0.00 1.01 $\pm$ 0.04 1.00 / 0.92 128 D 1.01 $\pm$ 0.01 1.08 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.00 0.98 $\pm$ 0.02 1.10 $\pm$ 0.03 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.02 $\pm$ 0.04 1.00 / 0.84 256 D 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 0.97 $\pm$ 0.03 1.00 $\pm$ 0.03 1.00 $\pm$ 0.01 1.01 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 / 0.68 50 full (F) 16 F 0.99 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01 0.65 $\pm$ 0.09 1.01 $\pm$ 0.15 0.97 $\pm$ 0.02 0.88 $\pm$ 0.01 0.94 $\pm$ 0.11 1.00/0.98 32 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01 0.97 $\pm$ 0.00 0.96 $\pm$ 0.01 1.09 $\pm$ 0.03 1.01 $\pm$ 0.02 0.93 $\pm$ 0.00 0.99 $\pm$ 0.04 0.99/0.95 64 F 1.02 $\pm$ 0.01 1.06 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.00 1.03 $\pm$ 0.01 1.11 $\pm$ 0.00 1.00 $\pm$ 0.01 0.99 $\pm$ 0.01 1.02 $\pm$ 0.04 0.97/0.89 128 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.98 $\pm$ 0.00 0.98 $\pm$ 0.01 1.03 $\pm$ 0.04 1.00 $\pm$ 0.01 1.01 $\pm$ 0.00 1.01 $\pm$ 0.02 0.88/0.72 256 F 1.00 $\pm$ 0.00 1.02 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.50/0.18 100 diagonal (D) 16 D 0.80 $\pm$ 0.07 0.89 $\pm$ 0.06 0.93 $\pm$ 0.03 0.96 $\pm$ 0.01 0.51 $\pm$ 0.09 0.81 $\pm$ 0.02 0.30 $\pm$ 0.07 0.54 $\pm$ 0.11 0.78 $\pm$ 0.03 0.83 $\pm$ 0.02 0.73 $\pm$ 0.21 1.00 / 0.99 32 D 0.95 $\pm$ 0.06 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.91 $\pm$ 0.06 0.80 $\pm$ 0.13 0.91 $\pm$ 0.05 0.62 $\pm$ 0.10 0.78 $\pm$ 0.25 0.91 $\pm$ 0.02 0.85 $\pm$ 0.01 0.87 $\pm$ 0.14 1.00 / 0.98 64 D 1.01 $\pm$ 0.03 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.90 $\pm$ 0.05 1.11 $\pm$ 0.07 0.96 $\pm$ 0.02 0.88 $\pm$ 0.02 0.98 $\pm$ 0.07 1.00 / 0.96 128 D 1.01 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.00 1.00 $\pm$ 0.03 1.11 $\pm$ 0.02 0.99 $\pm$ 0.01 0.92 $\pm$ 0.00 1.00 $\pm$ 0.05 1.00 / 0.92 256 D 1.00 $\pm$ 0.00 1.06 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00 1.00 $\pm$ 0.01 1.11 $\pm$ 0.03 1.00 $\pm$ 0.01 0.99 $\pm$ 0.02 1.01 $\pm$ 0.04 1.00 / 0.84 100 full (F) 16 F 0.95 $\pm$ 0.01 0.97 $\pm$ 0.03 0.97 $\pm$ 0.03 0.97 $\pm$ 0.03 0.93 $\pm$ 0.01 0.93 $\pm$ 0.01 0.66 $\pm$ 0.03 0.90 $\pm$ 0.16 0.87 $\pm$ 0.02 0.85 $\pm$ 0.01 0.90 $\pm$ 0.10 1.00/0.99 32 F 1.00 $\pm$ 0.02 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.97 $\pm$ 0.01 0.96 $\pm$ 0.00 0.87 $\pm$ 0.03 1.12 $\pm$ 0.03 0.96 $\pm$ 0.01 0.89 $\pm$ 0.00 0.98 $\pm$ 0.07 0.99/0.97 64 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.99 $\pm$ 0.01 1.10 $\pm$ 0.01 0.99 $\pm$ 0.02 0.93 $\pm$ 0.01 1.00 $\pm$ 0.04 0.97/0.93 128 F 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 1.03 $\pm$ 0.01 1.10 $\pm$ 0.01 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 1.02 $\pm$ 0.03 0.88/0.80 256 F 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 1.00 $\pm$ 0.03 1.01 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.01 0.50/0.34 500 diagonal (D) 16 D 0.57 $\pm$ 0.07 0.55 $\pm$ 0.03 0.83 $\pm$ 0.04 0.78 $\pm$ 0.16 0.85 $\pm$ 0.04 0.73 $\pm$ 0.07 0.24 $\pm$ 0.02 0.45 $\pm$ 0.01 0.76 $\pm$ 0.06 0.81 $\pm$ 0.00 0.66 $\pm$ 0.20 1.00 / 1.00 32 D 0.61 $\pm$ 0.12 0.55 $\pm$ 0.08 0.83 $\pm$ 0.02 0.84 $\pm$ 0.12 0.91 $\pm$ 0.02 0.75 $\pm$ 0.05 0.30 $\pm$ 0.05 0.49 $\pm$ 0.07 0.79 $\pm$ 0.04 0.82 $\pm$ 0.01 0.69 $\pm$ 0.20 1.00 / 1.00 64 D 0.73 $\pm$ 0.02 0.63 $\pm$ 0.11 0.89 $\pm$ 0.04 0.97 $\pm$ 0.00 0.94 $\pm$ 0.00 0.86 $\pm$ 0.03 0.46 $\pm$ 0.09 0.51 $\pm$ 0.07 0.82 $\pm$ 0.02 0.83 $\pm$ 0.01 0.77 $\pm$ 0.18 1.00 / 0.99 128 D 0.84 $\pm$ 0.00 0.92 $\pm$ 0.02 0.97 $\pm$ 0.03 0.98 $\pm$ 0.01 0.94 $\pm$ 0.00 0.90 $\pm$ 0.02 0.62 $\pm$ 0.14 0.54 $\pm$ 0.01 0.85 $\pm$ 0.05 0.83 $\pm$ 0.01 0.84 $\pm$ 0.15 1.00 / 0.98 256 D 0.99 $\pm$ 0.03 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.96 $\pm$ 0.00 0.93 $\pm$ 0.02 0.68 $\pm$ 0.05 0.85 $\pm$ 0.14 0.92 $\pm$ 0.02 0.85 $\pm$ 0.00 0.92 $\pm$ 0.11 1.00 / 0.97 500 full (F) 16 F 0.57 $\pm$ 0.01 0.43 $\pm$ 0.07 0.78 $\pm$ 0.01 0.97 $\pm$ 0.00 0.96 $\pm$ 0.00 0.86 $\pm$ 0.01 0.65 $\pm$ 0.00 0.55 $\pm$ 0.02 0.83 $\pm$ 0.01 0.84 $\pm$ 0.00 0.76 $\pm$ 0.17 1.00/1.00 32 F 0.79 $\pm$ 0.05 0.54 $\pm$ 0.04 0.93 $\pm$ 0.02 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.92 $\pm$ 0.00 0.70 $\pm$ 0.01 0.52 $\pm$ 0.00 0.86 $\pm$ 0.02 0.85 $\pm$ 0.00 0.81 $\pm$ 0.16 0.99/0.99 64 F 1.02 $\pm$ 0.00 0.96 $\pm$ 0.01 0.94 $\pm$ 0.01 1.00 $\pm$ 0.01 0.96 $\pm$ 0.00 0.97 $\pm$ 0.01 0.74 $\pm$ 0.01 0.55 $\pm$ 0.01 0.91 $\pm$ 0.01 0.87 $\pm$ 0.00 0.89 $\pm$ 0.14 0.97/0.96 128 F 1.03 $\pm$ 0.01 0.97 $\pm$ 0.02 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.88 $\pm$ 0.01 1.07 $\pm$ 0.02 0.98 $\pm$ 0.00 0.90 $\pm$ 0.00 0.98 $\pm$ 0.05 0.88/0.86 256 F 1.03 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.97 $\pm$ 0.00 1.00 $\pm$ 0.02 1.04 $\pm$ 0.02 1.00 $\pm$ 0.01 0.93 $\pm$ 0.00 1.00 $\pm$ 0.03 0.50/0.47
Table 5: Relative In-Distribution ROUGE-1 scores for various tasks and methods
F.4 Absolute Rouge-1 Performance and Compression Rate
Table 6 provides full results for absolute performance of Rouge-1, which shows the same trends as the results for absolute performance of Rouge-L (Table 4).
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 24.44 $\pm$ 0.00 1.60 $\pm$ 0.00 19.13 $\pm$ 0.00 39.22 $\pm$ 0.00 10.42 $\pm$ 0.00 39.88 $\pm$ 0.00 8.05 $\pm$ 0.00 6.96 $\pm$ 0.00 17.82 $\pm$ 0.00 55.03 $\pm$ 0.00 22.43 $\pm$ 16.49 1.00 / 1.00 lora 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 78.43 $\pm$ 0.00 74.90 $\pm$ 0.00 26.87 $\pm$ 0.00 95.00 $\pm$ 0.00 68.66 $\pm$ 0.00 81.14 $\pm$ 20.67 0.00/0.00 TIES 10 76.50 $\pm$ 0.00 49.00 $\pm$ 1.73 44.33 $\pm$ 4.04 9.80 $\pm$ 0.58 78.56 $\pm$ 0.96 40.44 $\pm$ 0.00 53.10 $\pm$ 0.67 15.48 $\pm$ 0.12 77.67 $\pm$ 1.15 54.89 $\pm$ 0.06 49.98 $\pm$ 23.33 1.00 / 1.00 50 55.80 $\pm$ 0.00 35.00 $\pm$ 0.00 18.00 $\pm$ 5.20 2.42 $\pm$ 0.50 85.78 $\pm$ 0.96 26.75 $\pm$ 0.00 49.96 $\pm$ 0.00 16.73 $\pm$ 0.00 30.00 $\pm$ 3.46 53.87 $\pm$ 0.02 37.43 $\pm$ 23.49 1.00 / 1.00 100 52.43 $\pm$ 0.00 34.00 $\pm$ 0.00 19.67 $\pm$ 4.62 1.09 $\pm$ 1.66 83.33 $\pm$ 0.00 28.57 $\pm$ 0.00 48.89 $\pm$ 0.00 15.18 $\pm$ 0.42 1.00 $\pm$ 0.00 53.44 $\pm$ 0.02 33.76 $\pm$ 25.22 1.00 / 1.00 500 35.18 $\pm$ 0.00 22.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 78.00 $\pm$ 0.00 24.32 $\pm$ 0.00 43.80 $\pm$ 0.04 9.96 $\pm$ 0.13 1.00 $\pm$ 0.00 27.90 $\pm$ 0.03 24.40 $\pm$ 23.79 1.00 / 1.00 SVD SVD 2 93.15 $\pm$ 2.77 92.24 $\pm$ 1.85 99.09 $\pm$ 0.18 93.44 $\pm$ 0.14 93.89 $\pm$ 0.35 77.33 $\pm$ 0.29 75.40 $\pm$ 1.01 26.90 $\pm$ 2.68 95.06 $\pm$ 1.35 67.71 $\pm$ 0.49 81.33 $\pm$ 20.85 0.88 / 0.88 SVD 4 94.01 $\pm$ 3.60 89.21 $\pm$ 0.71 99.05 $\pm$ 0.09 93.65 $\pm$ 0.03 94.66 $\pm$ 0.63 78.42 $\pm$ 0.23 74.09 $\pm$ 1.12 26.47 $\pm$ 2.06 93.98 $\pm$ 0.77 69.37 $\pm$ 0.21 81.22 $\pm$ 20.80 0.75 / 0.75 SVD 8 95.00 $\pm$ 0.00 87.40 $\pm$ 0.59 99.05 $\pm$ 0.09 93.65 $\pm$ 0.03 94.36 $\pm$ 0.38 78.21 $\pm$ 0.03 75.57 $\pm$ 0.00 26.88 $\pm$ 0.27 95.51 $\pm$ 1.09 69.33 $\pm$ 0.08 83.02 $\pm$ 19.87 0.50 / 0.50 SVD 16 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 78.44 $\pm$ 0.03 74.73 $\pm$ 0.18 26.87 $\pm$ 0.00 95.00 $\pm$ 0.00 68.62 $\pm$ 0.04 80.76 $\pm$ 21.05 0.00 / 0.00 10 diagonal (D) 16 D 96.67 $\pm$ 0.58 87.00 $\pm$ 1.00 99.00 $\pm$ 0.00 94.00 $\pm$ 0.67 93.11 $\pm$ 0.38 76.08 $\pm$ 0.17 77.26 $\pm$ 1.47 30.15 $\pm$ 0.72 94.00 $\pm$ 1.73 68.25 $\pm$ 0.18 81.55 $\pm$ 20.03 1.00 / 0.90 32 D 95.67 $\pm$ 0.58 90.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.00 $\pm$ 0.33 94.89 $\pm$ 0.51 77.46 $\pm$ 0.24 72.53 $\pm$ 1.00 27.98 $\pm$ 0.71 94.67 $\pm$ 0.58 69.16 $\pm$ 0.41 81.44 $\pm$ 20.80 1.00 / 0.80 64 D 95.00 $\pm$ 0.00 88.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.78 $\pm$ 0.38 78.28 $\pm$ 0.07 75.47 $\pm$ 0.58 26.53 $\pm$ 0.25 96.00 $\pm$ 0.00 69.36 $\pm$ 0.05 81.64 $\pm$ 21.06 1.00 / 0.60 128 D 95.00 $\pm$ 0.00 86.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 78.45 $\pm$ 0.16 75.46 $\pm$ 0.51 26.64 $\pm$ 0.23 95.00 $\pm$ 0.00 68.70 $\pm$ 0.14 81.29 $\pm$ 20.92 1.00 / 0.20 256 D 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 78.43 $\pm$ 0.00 74.90 $\pm$ 0.00 26.87 $\pm$ 0.00 95.00 $\pm$ 0.00 68.59 $\pm$ 0.03 81.18 $\pm$ 20.86 1.00 / -0.60 10 full (F) 16 F 97.00 $\pm$ 0.00 91.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.56 $\pm$ 0.69 77.64 $\pm$ 0.25 75.78 $\pm$ 1.25 28.71 $\pm$ 0.09 96.00 $\pm$ 1.00 68.69 $\pm$ 0.08 82.09 $\pm$ 20.68 1.00/0.90 32 F 96.67 $\pm$ 0.58 89.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.22 $\pm$ 0.19 94.44 $\pm$ 0.19 77.84 $\pm$ 0.21 72.24 $\pm$ 0.59 26.84 $\pm$ 0.50 94.67 $\pm$ 0.58 69.55 $\pm$ 0.08 81.38 $\pm$ 21.11 0.99/0.79 64 F 95.00 $\pm$ 0.00 88.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.56 $\pm$ 0.38 78.19 $\pm$ 0.08 75.97 $\pm$ 0.58 26.43 $\pm$ 0.34 96.00 $\pm$ 0.00 69.38 $\pm$ 0.11 81.69 $\pm$ 21.07 0.97/0.57 128 F 95.00 $\pm$ 0.00 86.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 78.46 $\pm$ 0.03 74.90 $\pm$ 0.00 26.72 $\pm$ 0.13 95.00 $\pm$ 0.00 68.65 $\pm$ 0.03 81.24 $\pm$ 20.91 0.88/0.07 256 F 95.00 $\pm$ 0.00 86.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.33 $\pm$ 0.00 78.44 $\pm$ 0.03 74.79 $\pm$ 0.19 26.87 $\pm$ 0.00 95.00 $\pm$ 0.00 68.64 $\pm$ 0.03 81.17 $\pm$ 20.86 0.50/-1.10 50 diagonal (D) 16 D 92.76 $\pm$ 3.53 84.67 $\pm$ 1.15 99.00 $\pm$ 0.00 86.17 $\pm$ 5.81 79.83 $\pm$ 6.08 73.55 $\pm$ 1.39 51.72 $\pm$ 3.78 23.75 $\pm$ 2.66 83.90 $\pm$ 6.43 59.05 $\pm$ 0.94 73.44 $\pm$ 22.08 1.00 / 0.98 32 D 95.33 $\pm$ 2.08 87.33 $\pm$ 2.08 99.00 $\pm$ 0.00 92.60 $\pm$ 0.29 90.35 $\pm$ 1.00 75.43 $\pm$ 1.33 63.84 $\pm$ 1.64 26.97 $\pm$ 3.21 93.33 $\pm$ 1.15 61.94 $\pm$ 0.32 78.61 $\pm$ 21.60 1.00 / 0.96 64 D 97.00 $\pm$ 0.00 90.33 $\pm$ 1.53 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 93.00 $\pm$ 0.58 76.27 $\pm$ 0.49 74.39 $\pm$ 0.90 29.28 $\pm$ 0.81 95.67 $\pm$ 0.58 64.84 $\pm$ 0.27 81.36 $\pm$ 20.83 1.00 / 0.92 128 D 96.33 $\pm$ 0.58 92.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.00 $\pm$ 0.58 77.24 $\pm$ 0.19 73.76 $\pm$ 1.25 29.58 $\pm$ 0.93 95.00 $\pm$ 0.00 69.04 $\pm$ 0.54 81.92 $\pm$ 20.44 1.00 / 0.84 256 D 95.67 $\pm$ 0.58 88.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 94.67 $\pm$ 0.67 78.45 $\pm$ 0.14 72.86 $\pm$ 2.07 27.00 $\pm$ 0.77 95.33 $\pm$ 0.58 69.61 $\pm$ 0.18 81.45 $\pm$ 21.00 1.00 / 0.68 50 full (F) 16 F 94.06 $\pm$ 3.54 85.67 $\pm$ 1.15 98.67 $\pm$ 0.58 90.35 $\pm$ 1.37 89.97 $\pm$ 1.78 74.46 $\pm$ 0.58 49.03 $\pm$ 7.07 27.14 $\pm$ 3.94 92.33 $\pm$ 1.53 60.26 $\pm$ 1.03 76.19 $\pm$ 22.80 1.00/0.98 32 F 97.00 $\pm$ 0.00 85.67 $\pm$ 1.53 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 92.22 $\pm$ 0.69 75.86 $\pm$ 0.22 71.68 $\pm$ 0.65 29.26 $\pm$ 0.70 95.67 $\pm$ 1.53 63.88 $\pm$ 0.10 80.39 $\pm$ 20.81 0.99/0.95 64 F 96.67 $\pm$ 0.58 91.00 $\pm$ 2.00 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.22 $\pm$ 0.51 77.17 $\pm$ 0.38 77.11 $\pm$ 0.51 29.75 $\pm$ 0.03 95.33 $\pm$ 0.58 68.13 $\pm$ 0.75 82.09 $\pm$ 20.33 0.97/0.89 128 F 97.00 $\pm$ 0.00 91.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.33 $\pm$ 0.00 94.11 $\pm$ 0.51 77.23 $\pm$ 0.17 73.67 $\pm$ 0.58 27.62 $\pm$ 1.12 95.00 $\pm$ 1.00 69.40 $\pm$ 0.16 81.74 $\pm$ 20.97 0.88/0.72 256 F 95.00 $\pm$ 0.00 88.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.44 $\pm$ 0.19 77.97 $\pm$ 0.24 75.47 $\pm$ 0.58 26.96 $\pm$ 0.09 96.00 $\pm$ 0.00 69.28 $\pm$ 0.05 81.58 $\pm$ 20.92 0.50/0.18 100 diagonal (D) 16 D 76.43 $\pm$ 7.07 76.67 $\pm$ 4.93 91.61 $\pm$ 2.75 89.99 $\pm$ 1.07 47.89 $\pm$ 8.62 63.17 $\pm$ 1.31 22.23 $\pm$ 5.27 14.46 $\pm$ 2.89 73.93 $\pm$ 3.13 57.17 $\pm$ 1.05 61.35 $\pm$ 25.78 1.00 / 0.99 32 D 90.10 $\pm$ 5.85 84.00 $\pm$ 1.00 99.00 $\pm$ 0.00 85.52 $\pm$ 5.34 75.88 $\pm$ 12.57 71.15 $\pm$ 3.61 46.10 $\pm$ 7.39 21.04 $\pm$ 6.76 86.67 $\pm$ 1.86 58.64 $\pm$ 1.02 71.81 $\pm$ 23.39 1.00 / 0.98 64 D 95.56 $\pm$ 2.49 86.67 $\pm$ 0.58 99.00 $\pm$ 0.00 92.24 $\pm$ 1.68 90.89 $\pm$ 1.17 74.57 $\pm$ 0.50 67.07 $\pm$ 3.81 29.78 $\pm$ 1.92 91.67 $\pm$ 2.31 60.28 $\pm$ 1.51 78.77 $\pm$ 20.77 1.00 / 0.96 128 D 96.00 $\pm$ 0.00 87.33 $\pm$ 1.15 99.00 $\pm$ 0.00 93.89 $\pm$ 0.19 93.00 $\pm$ 0.58 76.68 $\pm$ 0.18 74.84 $\pm$ 2.23 29.79 $\pm$ 0.50 93.67 $\pm$ 0.58 63.49 $\pm$ 0.34 80.77 $\pm$ 20.47 1.00 / 0.92 256 D 95.00 $\pm$ 0.00 91.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.56 $\pm$ 0.19 93.11 $\pm$ 0.19 76.93 $\pm$ 0.23 75.13 $\pm$ 0.84 29.75 $\pm$ 0.73 95.33 $\pm$ 0.58 67.89 $\pm$ 1.34 81.67 $\pm$ 20.28 1.00 / 0.84 100 full (F) 16 F 90.70 $\pm$ 1.07 83.00 $\pm$ 2.65 96.00 $\pm$ 3.00 91.22 $\pm$ 2.94 87.94 $\pm$ 0.54 73.07 $\pm$ 0.93 49.41 $\pm$ 2.04 24.17 $\pm$ 4.22 82.33 $\pm$ 2.08 58.18 $\pm$ 0.44 73.60 $\pm$ 22.23 1.00/0.99 32 F 95.33 $\pm$ 1.53 85.00 $\pm$ 1.00 99.00 $\pm$ 0.00 93.50 $\pm$ 0.22 91.44 $\pm$ 0.84 75.00 $\pm$ 0.19 65.09 $\pm$ 2.23 30.20 $\pm$ 0.81 91.67 $\pm$ 0.58 60.92 $\pm$ 0.26 78.72 $\pm$ 20.72 0.99/0.97 64 F 97.00 $\pm$ 0.00 85.67 $\pm$ 1.53 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 92.56 $\pm$ 0.19 76.01 $\pm$ 0.13 73.96 $\pm$ 0.89 29.46 $\pm$ 0.21 94.33 $\pm$ 1.53 64.07 $\pm$ 0.37 80.58 $\pm$ 20.59 0.97/0.93 128 F 96.33 $\pm$ 0.58 90.33 $\pm$ 0.58 99.00 $\pm$ 0.00 93.00 $\pm$ 0.00 93.89 $\pm$ 0.19 77.04 $\pm$ 0.30 77.33 $\pm$ 1.01 29.49 $\pm$ 0.35 96.00 $\pm$ 0.00 68.76 $\pm$ 0.25 82.12 $\pm$ 20.35 0.88/0.80 256 F 96.33 $\pm$ 0.58 88.67 $\pm$ 0.58 99.00 $\pm$ 0.00 93.67 $\pm$ 0.00 94.89 $\pm$ 0.19 78.16 $\pm$ 0.18 73.40 $\pm$ 0.12 26.86 $\pm$ 0.68 96.00 $\pm$ 0.00 69.47 $\pm$ 0.23 81.64 $\pm$ 21.15 0.50/0.34 500 diagonal (D) 16 D 54.44 $\pm$ 6.87 47.00 $\pm$ 2.83 82.21 $\pm$ 3.59 73.38 $\pm$ 14.97 80.13 $\pm$ 3.68 57.42 $\pm$ 5.29 18.33 $\pm$ 1.33 12.19 $\pm$ 0.30 72.67 $\pm$ 6.03 55.79 $\pm$ 0.20 55.64 $\pm$ 24.25 1.00 / 1.00 32 D 58.08 $\pm$ 11.52 47.00 $\pm$ 7.07 82.06 $\pm$ 1.69 78.62 $\pm$ 11.23 85.57 $\pm$ 1.48 59.19 $\pm$ 3.70 22.76 $\pm$ 3.95 13.15 $\pm$ 1.94 75.33 $\pm$ 4.04 56.07 $\pm$ 0.52 58.16 $\pm$ 24.56 1.00 / 1.00 64 D 69.21 $\pm$ 2.03 54.50 $\pm$ 9.19 88.33 $\pm$ 4.04 91.11 $\pm$ 0.38 88.78 $\pm$ 0.38 67.71 $\pm$ 2.59 34.79 $\pm$ 6.86 13.80 $\pm$ 1.95 77.67 $\pm$ 2.31 56.78 $\pm$ 0.73 64.61 $\pm$ 24.79 1.00 / 0.99 128 D 79.77 $\pm$ 0.37 79.50 $\pm$ 2.12 95.89 $\pm$ 2.83 91.89 $\pm$ 1.39 88.67 $\pm$ 0.00 70.27 $\pm$ 1.73 46.64 $\pm$ 10.58 14.63 $\pm$ 0.25 81.00 $\pm$ 5.00 56.88 $\pm$ 0.55 70.20 $\pm$ 24.63 1.00 / 0.98 256 D 93.83 $\pm$ 2.52 85.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 90.56 $\pm$ 0.38 73.25 $\pm$ 1.86 51.14 $\pm$ 3.86 22.93 $\pm$ 3.86 87.33 $\pm$ 2.31 58.48 $\pm$ 0.20 75.20 $\pm$ 23.90 1.00 / 0.97 500 full (F) 16 F 54.30 $\pm$ 1.13 37.00 $\pm$ 5.66 77.67 $\pm$ 0.58 91.00 $\pm$ 0.00 90.56 $\pm$ 0.19 67.63 $\pm$ 0.45 48.81 $\pm$ 0.35 14.70 $\pm$ 0.65 79.00 $\pm$ 1.00 57.66 $\pm$ 0.19 62.69 $\pm$ 23.46 1.00/1.00 32 F 75.10 $\pm$ 4.92 46.50 $\pm$ 3.54 91.67 $\pm$ 1.53 91.56 $\pm$ 0.19 91.56 $\pm$ 0.38 72.03 $\pm$ 0.15 52.63 $\pm$ 0.86 13.93 $\pm$ 0.02 81.67 $\pm$ 1.53 58.50 $\pm$ 0.20 68.24 $\pm$ 24.29 0.99/0.99 64 F 96.94 $\pm$ 0.42 82.50 $\pm$ 0.71 93.33 $\pm$ 0.58 93.89 $\pm$ 0.69 90.67 $\pm$ 0.00 75.99 $\pm$ 0.64 55.63 $\pm$ 1.07 14.74 $\pm$ 0.27 86.33 $\pm$ 0.58 59.43 $\pm$ 0.05 74.69 $\pm$ 25.01 0.97/0.96 128 F 97.67 $\pm$ 0.58 83.50 $\pm$ 2.12 98.00 $\pm$ 0.00 93.56 $\pm$ 0.19 92.00 $\pm$ 0.00 75.80 $\pm$ 0.16 66.19 $\pm$ 0.81 28.67 $\pm$ 0.49 93.00 $\pm$ 0.00 61.53 $\pm$ 0.13 78.84 $\pm$ 21.50 0.88/0.86 256 F 98.00 $\pm$ 0.00 88.50 $\pm$ 0.71 99.00 $\pm$ 0.00 93.78 $\pm$ 0.19 93.00 $\pm$ 0.88 76.33 $\pm$ 0.29 74.60 $\pm$ 1.21 27.82 $\pm$ 0.42 95.33 $\pm$ 0.58 63.70 $\pm$ 0.14 80.75 $\pm$ 21.60 0.50/0.47
Table 6: Absolute In-Distribution ROUGE-1 scores for various tasks and methods
F.5 Relative Exact-Match Performance and Compression Rate
Table 7 provides full results for relative performance of exact-match, which shows the same trends as the results for relative performance of Rouge-L (Table 3).
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.02 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.01 1.00 / 1.00 lora 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 / 0.00 TIES 10 0.69 $\pm$ 0.00 0.57 $\pm$ 0.02 0.45 $\pm$ 0.04 0.10 $\pm$ 0.01 0.57 $\pm$ 0.03 0.00 $\pm$ 0.00 0.39 $\pm$ 0.01 0.21 $\pm$ 0.00 0.82 $\pm$ 0.01 0.00 $\pm$ 0.00 0.38 $\pm$ 0.28 1.00 / 1.00 50 0.45 $\pm$ 0.00 0.41 $\pm$ 0.00 0.18 $\pm$ 0.05 0.03 $\pm$ 0.01 0.70 $\pm$ 0.02 0.00 $\pm$ 0.00 0.36 $\pm$ 0.00 0.21 $\pm$ 0.00 0.32 $\pm$ 0.04 0.00 $\pm$ 0.00 0.27 $\pm$ 0.22 1.00 / 1.00 100 0.41 $\pm$ 0.00 0.40 $\pm$ 0.00 0.20 $\pm$ 0.05 0.01 $\pm$ 0.02 0.65 $\pm$ 0.00 0.00 $\pm$ 0.00 0.36 $\pm$ 0.00 0.21 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.23 $\pm$ 0.22 1.00 / 1.00 500 0.22 $\pm$ 0.00 0.26 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.60 $\pm$ 0.00 0.00 $\pm$ 0.00 0.32 $\pm$ 0.00 0.07 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.15 $\pm$ 0.20 1.00 / 1.00 SVD SVD 2 0.98 $\pm$ 0.03 1.07 $\pm$ 0.02 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.01 0.98 $\pm$ 0.03 0.94 $\pm$ 0.01 1.03 $\pm$ 0.17 1.00 $\pm$ 0.01 0.15 $\pm$ 0.29 0.91 $\pm$ 0.28 0.88 / 0.88 SVD 4 0.99 $\pm$ 0.04 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.02 1.11 $\pm$ 0.00 0.97 $\pm$ 0.02 0.99 $\pm$ 0.13 0.99 $\pm$ 0.01 0.90 $\pm$ 0.17 1.00 $\pm$ 0.08 0.75 / 0.75 SVD 8 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.02 $\pm$ 0.05 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.02 0.50 / 0.50 SVD 16 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 / 0.00 10 diagonal (D) 16 D 1.02 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.02 0.96 $\pm$ 0.01 1.11 $\pm$ 0.11 0.89 $\pm$ 0.03 1.19 $\pm$ 0.04 0.99 $\pm$ 0.02 0.33 $\pm$ 0.58 0.95 $\pm$ 0.27 1.00 / 0.90 32 D 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01 1.02 $\pm$ 0.02 1.11 $\pm$ 0.00 0.93 $\pm$ 0.01 1.10 $\pm$ 0.04 1.00 $\pm$ 0.01 0.67 $\pm$ 0.58 0.98 $\pm$ 0.19 1.00 / 0.80 64 D 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.11 $\pm$ 0.00 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 0.67 $\pm$ 0.58 0.98 $\pm$ 0.19 1.00 / 0.60 128 D 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 / 0.20 256 D 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 / -0.60 10 full (F) 16 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.97 $\pm$ 0.03 1.15 $\pm$ 0.06 0.92 $\pm$ 0.02 1.17 $\pm$ 0.04 1.01 $\pm$ 0.01 0.67 $\pm$ 0.58 1.00 $\pm$ 0.20 1.00/0.90 32 F 1.02 $\pm$ 0.01 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01 1.00 $\pm$ 0.01 1.11 $\pm$ 0.00 0.92 $\pm$ 0.01 1.02 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.05 0.99/0.79 64 F 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.07 $\pm$ 0.06 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.03 0.97/0.57 128 F 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.88/0.07 256 F 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.50/-1.10 50 diagonal (D) 16 D 0.91 $\pm$ 0.06 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.91 $\pm$ 0.09 0.78 $\pm$ 0.05 0.89 $\pm$ 0.29 0.34 $\pm$ 0.06 0.50 $\pm$ 0.45 0.86 $\pm$ 0.07 0.00 $\pm$ 0.00 0.72 $\pm$ 0.35 1.00 / 0.98 32 D 1.00 $\pm$ 0.02 1.02 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.90 $\pm$ 0.03 0.85 $\pm$ 0.42 0.56 $\pm$ 0.04 0.98 $\pm$ 0.23 0.98 $\pm$ 0.01 0.00 $\pm$ 0.00 0.83 $\pm$ 0.34 1.00 / 0.96 64 D 1.02 $\pm$ 0.00 1.05 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.95 $\pm$ 0.02 1.15 $\pm$ 0.17 0.81 $\pm$ 0.03 1.14 $\pm$ 0.00 1.01 $\pm$ 0.01 0.00 $\pm$ 0.00 0.91 $\pm$ 0.33 1.00 / 0.92 128 D 1.01 $\pm$ 0.01 1.08 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.95 $\pm$ 0.02 1.04 $\pm$ 0.06 0.92 $\pm$ 0.03 1.21 $\pm$ 0.07 1.00 $\pm$ 0.00 0.67 $\pm$ 0.58 0.99 $\pm$ 0.20 1.00 / 0.84 256 D 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.01 $\pm$ 0.02 1.11 $\pm$ 0.00 0.95 $\pm$ 0.04 1.02 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.04 1.00 / 0.68 50 full (F) 16 F 0.96 $\pm$ 0.05 1.00 $\pm$ 0.01 1.00 $\pm$ 0.01 0.95 $\pm$ 0.04 0.87 $\pm$ 0.01 1.04 $\pm$ 0.06 0.31 $\pm$ 0.08 0.98 $\pm$ 0.23 0.97 $\pm$ 0.02 0.00 $\pm$ 0.00 0.81 $\pm$ 0.35 1.00/0.98 32 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.92 $\pm$ 0.03 1.15 $\pm$ 0.06 0.73 $\pm$ 0.04 1.17 $\pm$ 0.04 1.01 $\pm$ 0.02 0.00 $\pm$ 0.00 0.90 $\pm$ 0.33 0.99/0.95 64 F 1.02 $\pm$ 0.01 1.06 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.96 $\pm$ 0.02 1.22 $\pm$ 0.00 0.94 $\pm$ 0.01 1.17 $\pm$ 0.04 1.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.94 $\pm$ 0.33 0.97/0.89 128 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.99 $\pm$ 0.02 1.15 $\pm$ 0.06 0.92 $\pm$ 0.01 1.10 $\pm$ 0.08 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.02 $\pm$ 0.07 0.88/0.72 256 F 1.00 $\pm$ 0.00 1.02 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.04 $\pm$ 0.06 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.02 0.50/0.18 100 diagonal (D) 16 D 0.54 $\pm$ 0.16 0.89 $\pm$ 0.06 0.90 $\pm$ 0.04 0.89 $\pm$ 0.05 0.42 $\pm$ 0.08 0.44 $\pm$ 0.00 0.08 $\pm$ 0.02 0.00 $\pm$ 0.00 0.76 $\pm$ 0.05 0.00 $\pm$ 0.00 0.49 $\pm$ 0.36 1.00 / 0.99 32 D 0.85 $\pm$ 0.15 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.86 $\pm$ 0.13 0.70 $\pm$ 0.14 0.74 $\pm$ 0.28 0.28 $\pm$ 0.07 0.48 $\pm$ 0.55 0.91 $\pm$ 0.02 0.00 $\pm$ 0.00 0.68 $\pm$ 0.36 1.00 / 0.98 64 D 1.00 $\pm$ 0.04 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.02 0.88 $\pm$ 0.04 1.07 $\pm$ 0.06 0.58 $\pm$ 0.09 1.10 $\pm$ 0.04 0.96 $\pm$ 0.02 0.00 $\pm$ 0.00 0.86 $\pm$ 0.32 1.00 / 0.96 128 D 1.01 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.95 $\pm$ 0.02 1.11 $\pm$ 0.00 0.81 $\pm$ 0.06 1.21 $\pm$ 0.00 0.99 $\pm$ 0.01 0.00 $\pm$ 0.00 0.91 $\pm$ 0.33 1.00 / 0.92 256 D 1.00 $\pm$ 0.00 1.06 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.96 $\pm$ 0.01 1.11 $\pm$ 0.11 0.92 $\pm$ 0.02 1.21 $\pm$ 0.07 1.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.93 $\pm$ 0.33 1.00 / 0.84 100 full (F) 16 F 0.85 $\pm$ 0.03 0.97 $\pm$ 0.03 0.97 $\pm$ 0.03 0.95 $\pm$ 0.06 0.80 $\pm$ 0.02 0.81 $\pm$ 0.17 0.29 $\pm$ 0.04 0.60 $\pm$ 0.34 0.87 $\pm$ 0.02 0.00 $\pm$ 0.00 0.71 $\pm$ 0.33 1.00/0.99 32 F 0.99 $\pm$ 0.02 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.90 $\pm$ 0.03 1.04 $\pm$ 0.06 0.55 $\pm$ 0.04 1.07 $\pm$ 0.07 0.96 $\pm$ 0.01 0.00 $\pm$ 0.00 0.85 $\pm$ 0.32 0.99/0.97 64 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.94 $\pm$ 0.01 1.04 $\pm$ 0.06 0.78 $\pm$ 0.01 1.14 $\pm$ 0.00 0.99 $\pm$ 0.02 0.00 $\pm$ 0.00 0.89 $\pm$ 0.31 0.97/0.93 128 F 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.01 1.15 $\pm$ 0.06 0.94 $\pm$ 0.01 1.21 $\pm$ 0.00 1.01 $\pm$ 0.00 0.33 $\pm$ 0.58 0.97 $\pm$ 0.28 0.88/0.80 256 F 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.19 $\pm$ 0.06 0.93 $\pm$ 0.01 1.02 $\pm$ 0.04 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 1.02 $\pm$ 0.06 0.50/0.34 500 diagonal (D) 16 D 0.22 $\pm$ 0.10 0.55 $\pm$ 0.03 0.81 $\pm$ 0.05 0.33 $\pm$ 0.49 0.70 $\pm$ 0.03 0.15 $\pm$ 0.17 0.03 $\pm$ 0.01 0.00 $\pm$ 0.00 0.76 $\pm$ 0.06 0.00 $\pm$ 0.00 0.35 $\pm$ 0.35 1.00 / 1.00 32 D 0.27 $\pm$ 0.18 0.55 $\pm$ 0.08 0.82 $\pm$ 0.02 0.49 $\pm$ 0.37 0.75 $\pm$ 0.01 0.22 $\pm$ 0.11 0.05 $\pm$ 0.05 0.02 $\pm$ 0.04 0.79 $\pm$ 0.04 0.00 $\pm$ 0.00 0.39 $\pm$ 0.34 1.00 / 1.00 64 D 0.40 $\pm$ 0.04 0.63 $\pm$ 0.11 0.89 $\pm$ 0.04 0.91 $\pm$ 0.01 0.80 $\pm$ 0.01 0.48 $\pm$ 0.06 0.13 $\pm$ 0.04 0.05 $\pm$ 0.08 0.82 $\pm$ 0.02 0.00 $\pm$ 0.00 0.51 $\pm$ 0.35 1.00 / 0.99 128 D 0.61 $\pm$ 0.04 0.92 $\pm$ 0.02 0.97 $\pm$ 0.03 0.93 $\pm$ 0.05 0.80 $\pm$ 0.00 0.74 $\pm$ 0.17 0.22 $\pm$ 0.11 0.12 $\pm$ 0.08 0.85 $\pm$ 0.05 0.00 $\pm$ 0.00 0.61 $\pm$ 0.36 1.00 / 0.98 256 D 0.95 $\pm$ 0.02 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.86 $\pm$ 0.01 0.85 $\pm$ 0.28 0.28 $\pm$ 0.06 0.55 $\pm$ 0.39 0.92 $\pm$ 0.02 0.00 $\pm$ 0.00 0.73 $\pm$ 0.36 1.00 / 0.97 500 full (F) 16 F 0.21 $\pm$ 0.02 0.43 $\pm$ 0.07 0.78 $\pm$ 0.01 0.90 $\pm$ 0.00 0.86 $\pm$ 0.01 0.59 $\pm$ 0.06 0.21 $\pm$ 0.01 0.12 $\pm$ 0.04 0.83 $\pm$ 0.01 0.00 $\pm$ 0.00 0.50 $\pm$ 0.34 1.00/1.00 32 F 0.54 $\pm$ 0.08 0.54 $\pm$ 0.04 0.93 $\pm$ 0.02 0.92 $\pm$ 0.01 0.90 $\pm$ 0.01 0.63 $\pm$ 0.13 0.26 $\pm$ 0.02 0.14 $\pm$ 0.00 0.86 $\pm$ 0.02 0.00 $\pm$ 0.00 0.57 $\pm$ 0.34 0.99/0.99 64 F 0.99 $\pm$ 0.03 0.96 $\pm$ 0.01 0.94 $\pm$ 0.01 1.01 $\pm$ 0.03 0.87 $\pm$ 0.00 1.04 $\pm$ 0.17 0.36 $\pm$ 0.00 0.14 $\pm$ 0.00 0.91 $\pm$ 0.01 0.00 $\pm$ 0.00 0.71 $\pm$ 0.39 0.97/0.96 128 F 1.02 $\pm$ 0.01 0.97 $\pm$ 0.02 0.99 $\pm$ 0.00 1.00 $\pm$ 0.01 0.92 $\pm$ 0.00 1.15 $\pm$ 0.06 0.61 $\pm$ 0.01 1.07 $\pm$ 0.00 0.98 $\pm$ 0.00 0.00 $\pm$ 0.00 0.87 $\pm$ 0.33 0.88/0.86 256 F 1.03 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.95 $\pm$ 0.03 1.00 $\pm$ 0.00 0.78 $\pm$ 0.01 1.07 $\pm$ 0.00 1.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.88 $\pm$ 0.31 0.50/0.47
Table 7: Relative In-Distribution exact match scores for various tasks and methods
F.6 Absolute Exact-Match Performance and Compression Rate
Table 8 provides full results for absolute performance of exact-match, which shows the same trends as the results for absolute performance of Rouge-L (Table 4).
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.02 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.01 1.00 / 1.00 lora 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 / 0.00 TIES 10 0.69 $\pm$ 0.00 0.57 $\pm$ 0.02 0.45 $\pm$ 0.04 0.10 $\pm$ 0.01 0.57 $\pm$ 0.03 0.00 $\pm$ 0.00 0.39 $\pm$ 0.01 0.21 $\pm$ 0.00 0.82 $\pm$ 0.01 0.00 $\pm$ 0.00 0.38 $\pm$ 0.28 1.00 / 1.00 50 0.45 $\pm$ 0.00 0.41 $\pm$ 0.00 0.18 $\pm$ 0.05 0.03 $\pm$ 0.01 0.70 $\pm$ 0.02 0.00 $\pm$ 0.00 0.36 $\pm$ 0.00 0.21 $\pm$ 0.00 0.32 $\pm$ 0.04 0.00 $\pm$ 0.00 0.27 $\pm$ 0.22 1.00 / 1.00 100 0.41 $\pm$ 0.00 0.40 $\pm$ 0.00 0.20 $\pm$ 0.05 0.01 $\pm$ 0.02 0.65 $\pm$ 0.00 0.00 $\pm$ 0.00 0.36 $\pm$ 0.00 0.21 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.23 $\pm$ 0.22 1.00 / 1.00 500 0.22 $\pm$ 0.00 0.26 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.60 $\pm$ 0.00 0.00 $\pm$ 0.00 0.32 $\pm$ 0.00 0.07 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.15 $\pm$ 0.20 1.00 / 1.00 SVD SVD 2 0.98 $\pm$ 0.03 1.07 $\pm$ 0.02 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.98 $\pm$ 0.01 0.98 $\pm$ 0.03 0.94 $\pm$ 0.01 1.03 $\pm$ 0.17 1.00 $\pm$ 0.01 0.15 $\pm$ 0.29 0.91 $\pm$ 0.28 0.88 / 0.88 SVD 4 0.99 $\pm$ 0.04 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.02 1.11 $\pm$ 0.00 0.97 $\pm$ 0.02 0.99 $\pm$ 0.13 0.99 $\pm$ 0.01 0.90 $\pm$ 0.17 1.00 $\pm$ 0.08 0.75 / 0.75 SVD 8 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.02 $\pm$ 0.05 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.02 0.50 / 0.50 SVD 16 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 / 0.00 10 diagonal (D) 16 D 1.02 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.02 0.96 $\pm$ 0.01 1.11 $\pm$ 0.11 0.89 $\pm$ 0.03 1.19 $\pm$ 0.04 0.99 $\pm$ 0.02 0.33 $\pm$ 0.58 0.95 $\pm$ 0.27 1.00 / 0.90 32 D 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01 1.02 $\pm$ 0.02 1.11 $\pm$ 0.00 0.93 $\pm$ 0.01 1.10 $\pm$ 0.04 1.00 $\pm$ 0.01 0.67 $\pm$ 0.58 0.98 $\pm$ 0.19 1.00 / 0.80 64 D 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.11 $\pm$ 0.00 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 0.67 $\pm$ 0.58 0.98 $\pm$ 0.19 1.00 / 0.60 128 D 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 / 0.20 256 D 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 / -0.60 10 full (F) 16 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.97 $\pm$ 0.03 1.15 $\pm$ 0.06 0.92 $\pm$ 0.02 1.17 $\pm$ 0.04 1.01 $\pm$ 0.01 0.67 $\pm$ 0.58 1.00 $\pm$ 0.20 1.00/0.90 32 F 1.02 $\pm$ 0.01 1.04 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.01 1.00 $\pm$ 0.01 1.11 $\pm$ 0.00 0.92 $\pm$ 0.01 1.02 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.05 0.99/0.79 64 F 1.00 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.07 $\pm$ 0.06 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.03 0.97/0.57 128 F 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.88/0.07 256 F 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.50/-1.10 50 diagonal (D) 16 D 0.91 $\pm$ 0.06 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.91 $\pm$ 0.09 0.78 $\pm$ 0.05 0.89 $\pm$ 0.29 0.34 $\pm$ 0.06 0.50 $\pm$ 0.45 0.86 $\pm$ 0.07 0.00 $\pm$ 0.00 0.72 $\pm$ 0.35 1.00 / 0.98 32 D 1.00 $\pm$ 0.02 1.02 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.90 $\pm$ 0.03 0.85 $\pm$ 0.42 0.56 $\pm$ 0.04 0.98 $\pm$ 0.23 0.98 $\pm$ 0.01 0.00 $\pm$ 0.00 0.83 $\pm$ 0.34 1.00 / 0.96 64 D 1.02 $\pm$ 0.00 1.05 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.95 $\pm$ 0.02 1.15 $\pm$ 0.17 0.81 $\pm$ 0.03 1.14 $\pm$ 0.00 1.01 $\pm$ 0.01 0.00 $\pm$ 0.00 0.91 $\pm$ 0.33 1.00 / 0.92 128 D 1.01 $\pm$ 0.01 1.08 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.95 $\pm$ 0.02 1.04 $\pm$ 0.06 0.92 $\pm$ 0.03 1.21 $\pm$ 0.07 1.00 $\pm$ 0.00 0.67 $\pm$ 0.58 0.99 $\pm$ 0.20 1.00 / 0.84 256 D 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.01 $\pm$ 0.02 1.11 $\pm$ 0.00 0.95 $\pm$ 0.04 1.02 $\pm$ 0.04 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.04 1.00 / 0.68 50 full (F) 16 F 0.96 $\pm$ 0.05 1.00 $\pm$ 0.01 1.00 $\pm$ 0.01 0.95 $\pm$ 0.04 0.87 $\pm$ 0.01 1.04 $\pm$ 0.06 0.31 $\pm$ 0.08 0.98 $\pm$ 0.23 0.97 $\pm$ 0.02 0.00 $\pm$ 0.00 0.81 $\pm$ 0.35 1.00/0.98 32 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.92 $\pm$ 0.03 1.15 $\pm$ 0.06 0.73 $\pm$ 0.04 1.17 $\pm$ 0.04 1.01 $\pm$ 0.02 0.00 $\pm$ 0.00 0.90 $\pm$ 0.33 0.99/0.95 64 F 1.02 $\pm$ 0.01 1.06 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.96 $\pm$ 0.02 1.22 $\pm$ 0.00 0.94 $\pm$ 0.01 1.17 $\pm$ 0.04 1.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.94 $\pm$ 0.33 0.97/0.89 128 F 1.02 $\pm$ 0.00 1.06 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.99 $\pm$ 0.02 1.15 $\pm$ 0.06 0.92 $\pm$ 0.01 1.10 $\pm$ 0.08 1.00 $\pm$ 0.01 1.00 $\pm$ 0.00 1.02 $\pm$ 0.07 0.88/0.72 256 F 1.00 $\pm$ 0.00 1.02 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 1.04 $\pm$ 0.06 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 1.01 $\pm$ 0.02 0.50/0.18 100 diagonal (D) 16 D 0.54 $\pm$ 0.16 0.89 $\pm$ 0.06 0.90 $\pm$ 0.04 0.89 $\pm$ 0.05 0.42 $\pm$ 0.08 0.44 $\pm$ 0.00 0.08 $\pm$ 0.02 0.00 $\pm$ 0.00 0.76 $\pm$ 0.05 0.00 $\pm$ 0.00 0.49 $\pm$ 0.36 1.00 / 0.99 32 D 0.85 $\pm$ 0.15 0.98 $\pm$ 0.01 1.00 $\pm$ 0.00 0.86 $\pm$ 0.13 0.70 $\pm$ 0.14 0.74 $\pm$ 0.28 0.28 $\pm$ 0.07 0.48 $\pm$ 0.55 0.91 $\pm$ 0.02 0.00 $\pm$ 0.00 0.68 $\pm$ 0.36 1.00 / 0.98 64 D 1.00 $\pm$ 0.04 1.01 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.02 0.88 $\pm$ 0.04 1.07 $\pm$ 0.06 0.58 $\pm$ 0.09 1.10 $\pm$ 0.04 0.96 $\pm$ 0.02 0.00 $\pm$ 0.00 0.86 $\pm$ 0.32 1.00 / 0.96 128 D 1.01 $\pm$ 0.00 1.02 $\pm$ 0.01 1.00 $\pm$ 0.00 1.01 $\pm$ 0.01 0.95 $\pm$ 0.02 1.11 $\pm$ 0.00 0.81 $\pm$ 0.06 1.21 $\pm$ 0.00 0.99 $\pm$ 0.01 0.00 $\pm$ 0.00 0.91 $\pm$ 0.33 1.00 / 0.92 256 D 1.00 $\pm$ 0.00 1.06 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.96 $\pm$ 0.01 1.11 $\pm$ 0.11 0.92 $\pm$ 0.02 1.21 $\pm$ 0.07 1.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.93 $\pm$ 0.33 1.00 / 0.84 100 full (F) 16 F 0.85 $\pm$ 0.03 0.97 $\pm$ 0.03 0.97 $\pm$ 0.03 0.95 $\pm$ 0.06 0.80 $\pm$ 0.02 0.81 $\pm$ 0.17 0.29 $\pm$ 0.04 0.60 $\pm$ 0.34 0.87 $\pm$ 0.02 0.00 $\pm$ 0.00 0.71 $\pm$ 0.33 1.00/0.99 32 F 0.99 $\pm$ 0.02 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.90 $\pm$ 0.03 1.04 $\pm$ 0.06 0.55 $\pm$ 0.04 1.07 $\pm$ 0.07 0.96 $\pm$ 0.01 0.00 $\pm$ 0.00 0.85 $\pm$ 0.32 0.99/0.97 64 F 1.02 $\pm$ 0.00 1.00 $\pm$ 0.02 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.94 $\pm$ 0.01 1.04 $\pm$ 0.06 0.78 $\pm$ 0.01 1.14 $\pm$ 0.00 0.99 $\pm$ 0.02 0.00 $\pm$ 0.00 0.89 $\pm$ 0.31 0.97/0.93 128 F 1.01 $\pm$ 0.01 1.05 $\pm$ 0.01 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.01 1.15 $\pm$ 0.06 0.94 $\pm$ 0.01 1.21 $\pm$ 0.00 1.01 $\pm$ 0.00 0.33 $\pm$ 0.58 0.97 $\pm$ 0.28 0.88/0.80 256 F 1.01 $\pm$ 0.01 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.02 $\pm$ 0.01 1.19 $\pm$ 0.06 0.93 $\pm$ 0.01 1.02 $\pm$ 0.04 1.01 $\pm$ 0.00 1.00 $\pm$ 0.00 1.02 $\pm$ 0.06 0.50/0.34 500 diagonal (D) 16 D 0.22 $\pm$ 0.10 0.55 $\pm$ 0.03 0.81 $\pm$ 0.05 0.33 $\pm$ 0.49 0.70 $\pm$ 0.03 0.15 $\pm$ 0.17 0.03 $\pm$ 0.01 0.00 $\pm$ 0.00 0.76 $\pm$ 0.06 0.00 $\pm$ 0.00 0.35 $\pm$ 0.35 1.00 / 1.00 32 D 0.27 $\pm$ 0.18 0.55 $\pm$ 0.08 0.82 $\pm$ 0.02 0.49 $\pm$ 0.37 0.75 $\pm$ 0.01 0.22 $\pm$ 0.11 0.05 $\pm$ 0.05 0.02 $\pm$ 0.04 0.79 $\pm$ 0.04 0.00 $\pm$ 0.00 0.39 $\pm$ 0.34 1.00 / 1.00 64 D 0.40 $\pm$ 0.04 0.63 $\pm$ 0.11 0.89 $\pm$ 0.04 0.91 $\pm$ 0.01 0.80 $\pm$ 0.01 0.48 $\pm$ 0.06 0.13 $\pm$ 0.04 0.05 $\pm$ 0.08 0.82 $\pm$ 0.02 0.00 $\pm$ 0.00 0.51 $\pm$ 0.35 1.00 / 0.99 128 D 0.61 $\pm$ 0.04 0.92 $\pm$ 0.02 0.97 $\pm$ 0.03 0.93 $\pm$ 0.05 0.80 $\pm$ 0.00 0.74 $\pm$ 0.17 0.22 $\pm$ 0.11 0.12 $\pm$ 0.08 0.85 $\pm$ 0.05 0.00 $\pm$ 0.00 0.61 $\pm$ 0.36 1.00 / 0.98 256 D 0.95 $\pm$ 0.02 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.86 $\pm$ 0.01 0.85 $\pm$ 0.28 0.28 $\pm$ 0.06 0.55 $\pm$ 0.39 0.92 $\pm$ 0.02 0.00 $\pm$ 0.00 0.73 $\pm$ 0.36 1.00 / 0.97 500 full (F) 16 F 0.21 $\pm$ 0.02 0.43 $\pm$ 0.07 0.78 $\pm$ 0.01 0.90 $\pm$ 0.00 0.86 $\pm$ 0.01 0.59 $\pm$ 0.06 0.21 $\pm$ 0.01 0.12 $\pm$ 0.04 0.83 $\pm$ 0.01 0.00 $\pm$ 0.00 0.50 $\pm$ 0.34 1.00/1.00 32 F 0.54 $\pm$ 0.08 0.54 $\pm$ 0.04 0.93 $\pm$ 0.02 0.92 $\pm$ 0.01 0.90 $\pm$ 0.01 0.63 $\pm$ 0.13 0.26 $\pm$ 0.02 0.14 $\pm$ 0.00 0.86 $\pm$ 0.02 0.00 $\pm$ 0.00 0.57 $\pm$ 0.34 0.99/0.99 64 F 0.99 $\pm$ 0.03 0.96 $\pm$ 0.01 0.94 $\pm$ 0.01 1.01 $\pm$ 0.03 0.87 $\pm$ 0.00 1.04 $\pm$ 0.17 0.36 $\pm$ 0.00 0.14 $\pm$ 0.00 0.91 $\pm$ 0.01 0.00 $\pm$ 0.00 0.71 $\pm$ 0.39 0.97/0.96 128 F 1.02 $\pm$ 0.01 0.97 $\pm$ 0.02 0.99 $\pm$ 0.00 1.00 $\pm$ 0.01 0.92 $\pm$ 0.00 1.15 $\pm$ 0.06 0.61 $\pm$ 0.01 1.07 $\pm$ 0.00 0.98 $\pm$ 0.00 0.00 $\pm$ 0.00 0.87 $\pm$ 0.33 0.88/0.86 256 F 1.03 $\pm$ 0.00 1.03 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.01 0.95 $\pm$ 0.03 1.00 $\pm$ 0.00 0.78 $\pm$ 0.01 1.07 $\pm$ 0.00 1.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.88 $\pm$ 0.31 0.50/0.47
Table 8: Absolute In-Distribution exact match scores for various tasks and methods
F.7 Loss and Compression Rate
Table 9 provides full results for test loss (cross-entropy), which shows the same trends as the results for relative performance of Rouge-L (Table 3).
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 8.59 $\pm$ 0.08 9.15 $\pm$ 0.00 2.55 $\pm$ 0.00 2.88 $\pm$ 0.00 2.34 $\pm$ 0.00 3.46 $\pm$ 0.04 6.40 $\pm$ 0.18 5.55 $\pm$ 0.00 8.60 $\pm$ 0.00 2.67 $\pm$ 0.00 5.19 $\pm$ 2.65 1.00 / 1.00 lora 0.36 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.76 $\pm$ 0.02 1.17 $\pm$ 0.07 1.94 $\pm$ 0.00 0.16 $\pm$ 0.00 0.85 $\pm$ 0.00 0.57 $\pm$ 0.59 0.00 / 0.00 SVD SVD 2 0.32 $\pm$ 0.01 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.76 $\pm$ 0.02 1.13 $\pm$ 0.08 1.94 $\pm$ 0.00 0.13 $\pm$ 0.00 0.97 $\pm$ 0.00 0.57 $\pm$ 0.60 0.88 / 0.88 SVD 4 0.33 $\pm$ 0.01 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.76 $\pm$ 0.02 1.14 $\pm$ 0.08 1.94 $\pm$ 0.00 0.14 $\pm$ 0.00 0.86 $\pm$ 0.00 0.56 $\pm$ 0.59 0.75 / 0.75 SVD 8 0.35 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.77 $\pm$ 0.02 1.16 $\pm$ 0.07 1.94 $\pm$ 0.00 0.15 $\pm$ 0.00 0.84 $\pm$ 0.00 0.51 $\pm$ 0.58 0.50 / 0.50 SVD 16 0.36 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.76 $\pm$ 0.02 1.14 $\pm$ 0.06 1.94 $\pm$ 0.00 0.16 $\pm$ 0.00 0.85 $\pm$ 0.00 0.56 $\pm$ 0.59 0.00 / 0.00 10 diagonal (D) 16 D 0.33 $\pm$ 0.01 0.15 $\pm$ 0.01 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.76 $\pm$ 0.03 1.13 $\pm$ 0.08 1.95 $\pm$ 0.01 0.14 $\pm$ 0.00 1.00 $\pm$ 0.02 0.57 $\pm$ 0.61 1.00 / 0.90 32 D 0.33 $\pm$ 0.01 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.75 $\pm$ 0.02 1.11 $\pm$ 0.07 1.93 $\pm$ 0.00 0.14 $\pm$ 0.01 0.88 $\pm$ 0.00 0.55 $\pm$ 0.60 1.00 / 0.80 64 D 0.35 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.75 $\pm$ 0.02 1.11 $\pm$ 0.07 1.94 $\pm$ 0.00 0.15 $\pm$ 0.00 0.84 $\pm$ 0.00 0.55 $\pm$ 0.59 1.00 / 0.60 128 D 0.35 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.75 $\pm$ 0.02 1.11 $\pm$ 0.07 1.94 $\pm$ 0.00 0.16 $\pm$ 0.00 0.84 $\pm$ 0.00 0.56 $\pm$ 0.59 1.00 / 0.20 256 D 0.36 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.75 $\pm$ 0.02 1.12 $\pm$ 0.07 1.94 $\pm$ 0.00 0.16 $\pm$ 0.00 0.85 $\pm$ 0.00 0.56 $\pm$ 0.59 1.00 / -0.60 10 full (F) 16 F 0.33 $\pm$ 0.00 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.76 $\pm$ 0.02 1.20 $\pm$ 0.02 1.95 $\pm$ 0.00 0.13 $\pm$ 0.00 0.97 $\pm$ 0.00 0.57 $\pm$ 0.61 1.00/0.90 32 F 0.33 $\pm$ 0.01 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.75 $\pm$ 0.02 1.11 $\pm$ 0.07 1.94 $\pm$ 0.00 0.14 $\pm$ 0.00 0.86 $\pm$ 0.00 0.55 $\pm$ 0.60 0.99/0.79 64 F 0.34 $\pm$ 0.01 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.75 $\pm$ 0.02 1.11 $\pm$ 0.07 1.94 $\pm$ 0.00 0.15 $\pm$ 0.00 0.84 $\pm$ 0.00 0.55 $\pm$ 0.59 0.97/0.57 128 F 0.35 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.75 $\pm$ 0.02 1.12 $\pm$ 0.07 1.94 $\pm$ 0.00 0.16 $\pm$ 0.00 0.84 $\pm$ 0.00 0.56 $\pm$ 0.59 0.88/0.07 256 F 0.36 $\pm$ 0.01 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.75 $\pm$ 0.02 1.12 $\pm$ 0.07 1.94 $\pm$ 0.00 0.16 $\pm$ 0.00 0.85 $\pm$ 0.00 0.56 $\pm$ 0.59 0.50/-1.10 50 diagonal (D) 16 D 0.61 $\pm$ 0.06 0.19 $\pm$ 0.02 0.03 $\pm$ 0.01 0.29 $\pm$ 0.04 0.36 $\pm$ 0.04 0.95 $\pm$ 0.05 1.73 $\pm$ 0.21 2.66 $\pm$ 0.22 0.32 $\pm$ 0.11 1.98 $\pm$ 0.01 0.91 $\pm$ 0.88 1.00 / 0.98 32 D 0.37 $\pm$ 0.02 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.19 $\pm$ 0.03 0.18 $\pm$ 0.01 0.85 $\pm$ 0.05 1.37 $\pm$ 0.14 2.12 $\pm$ 0.05 0.16 $\pm$ 0.00 1.65 $\pm$ 0.03 0.71 $\pm$ 0.73 1.00 / 0.96 64 D 0.33 $\pm$ 0.02 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.79 $\pm$ 0.02 1.12 $\pm$ 0.08 1.97 $\pm$ 0.01 0.13 $\pm$ 0.01 1.13 $\pm$ 0.03 0.59 $\pm$ 0.63 1.00 / 0.92 128 D 0.33 $\pm$ 0.01 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.76 $\pm$ 0.03 1.10 $\pm$ 0.05 1.93 $\pm$ 0.01 0.14 $\pm$ 0.00 0.93 $\pm$ 0.01 0.56 $\pm$ 0.60 1.00 / 0.84 256 D 0.34 $\pm$ 0.01 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.76 $\pm$ 0.03 1.11 $\pm$ 0.05 1.93 $\pm$ 0.00 0.15 $\pm$ 0.00 0.85 $\pm$ 0.00 0.55 $\pm$ 0.59 1.00 / 0.68 50 full (F) 16 F 0.47 $\pm$ 0.06 0.17 $\pm$ 0.00 0.02 $\pm$ 0.00 0.20 $\pm$ 0.02 0.19 $\pm$ 0.04 0.86 $\pm$ 0.03 1.71 $\pm$ 0.10 2.20 $\pm$ 0.04 0.17 $\pm$ 0.01 1.84 $\pm$ 0.07 0.78 $\pm$ 0.80 1.00/0.98 32 F 0.36 $\pm$ 0.02 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.14 $\pm$ 0.00 0.11 $\pm$ 0.00 0.80 $\pm$ 0.03 1.14 $\pm$ 0.08 2.00 $\pm$ 0.01 0.14 $\pm$ 0.00 1.32 $\pm$ 0.02 0.62 $\pm$ 0.65 0.99/0.95 64 F 0.33 $\pm$ 0.01 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.77 $\pm$ 0.03 1.10 $\pm$ 0.06 1.94 $\pm$ 0.00 0.13 $\pm$ 0.00 1.02 $\pm$ 0.00 0.57 $\pm$ 0.61 0.97/0.89 128 F 0.33 $\pm$ 0.01 0.16 $\pm$ 0.00 0.00 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.76 $\pm$ 0.03 1.11 $\pm$ 0.05 1.93 $\pm$ 0.00 0.14 $\pm$ 0.00 0.87 $\pm$ 0.00 0.55 $\pm$ 0.60 0.88/0.72 256 F 0.35 $\pm$ 0.01 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.76 $\pm$ 0.03 1.11 $\pm$ 0.05 1.94 $\pm$ 0.00 0.15 $\pm$ 0.00 0.84 $\pm$ 0.00 0.55 $\pm$ 0.59 0.50/0.18 100 diagonal (D) 16 D 1.69 $\pm$ 0.49 0.26 $\pm$ 0.04 0.18 $\pm$ 0.07 0.34 $\pm$ 0.02 1.01 $\pm$ 0.20 1.45 $\pm$ 0.10 3.59 $\pm$ 0.25 3.72 $\pm$ 0.72 0.44 $\pm$ 0.20 2.37 $\pm$ 0.09 1.51 $\pm$ 1.32 1.00 / 0.99 32 D 0.67 $\pm$ 0.24 0.18 $\pm$ 0.01 0.06 $\pm$ 0.05 0.31 $\pm$ 0.06 0.35 $\pm$ 0.08 1.04 $\pm$ 0.15 1.97 $\pm$ 0.13 2.88 $\pm$ 0.70 0.22 $\pm$ 0.01 2.12 $\pm$ 0.07 0.98 $\pm$ 0.98 1.00 / 0.98 64 D 0.39 $\pm$ 0.06 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.18 $\pm$ 0.02 0.14 $\pm$ 0.01 0.86 $\pm$ 0.02 1.39 $\pm$ 0.07 2.18 $\pm$ 0.04 0.17 $\pm$ 0.00 1.79 $\pm$ 0.02 0.73 $\pm$ 0.76 1.00 / 0.96 128 D 0.32 $\pm$ 0.00 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.79 $\pm$ 0.02 1.19 $\pm$ 0.02 2.00 $\pm$ 0.01 0.14 $\pm$ 0.01 1.24 $\pm$ 0.04 0.61 $\pm$ 0.65 1.00 / 0.92 256 D 0.32 $\pm$ 0.00 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.77 $\pm$ 0.02 1.16 $\pm$ 0.00 1.94 $\pm$ 0.00 0.13 $\pm$ 0.00 0.96 $\pm$ 0.01 0.56 $\pm$ 0.61 1.00 / 0.84 100 full (F) 16 F 0.66 $\pm$ 0.07 0.19 $\pm$ 0.01 0.03 $\pm$ 0.01 0.25 $\pm$ 0.02 0.29 $\pm$ 0.02 0.99 $\pm$ 0.07 2.50 $\pm$ 0.51 2.63 $\pm$ 0.03 0.24 $\pm$ 0.02 2.21 $\pm$ 0.08 1.00 $\pm$ 1.01 1.00/0.99 32 F 0.40 $\pm$ 0.00 0.17 $\pm$ 0.00 0.01 $\pm$ 0.00 0.15 $\pm$ 0.01 0.13 $\pm$ 0.01 0.85 $\pm$ 0.02 1.53 $\pm$ 0.12 2.17 $\pm$ 0.06 0.15 $\pm$ 0.01 1.93 $\pm$ 0.04 0.75 $\pm$ 0.80 0.99/0.97 64 F 0.34 $\pm$ 0.01 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.11 $\pm$ 0.00 0.79 $\pm$ 0.01 1.23 $\pm$ 0.07 1.98 $\pm$ 0.01 0.15 $\pm$ 0.00 1.26 $\pm$ 0.01 0.61 $\pm$ 0.65 0.97/0.93 128 F 0.32 $\pm$ 0.00 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.77 $\pm$ 0.02 1.16 $\pm$ 0.01 1.94 $\pm$ 0.00 0.13 $\pm$ 0.00 0.99 $\pm$ 0.01 0.57 $\pm$ 0.61 0.88/0.80 256 F 0.33 $\pm$ 0.00 0.16 $\pm$ 0.00 0.00 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.76 $\pm$ 0.02 1.15 $\pm$ 0.01 1.93 $\pm$ 0.00 0.14 $\pm$ 0.00 0.86 $\pm$ 0.00 0.56 $\pm$ 0.60 0.50/0.34 500 diagonal (D) 16 D 2.95 $\pm$ 0.28 0.73 $\pm$ 0.29 0.27 $\pm$ 0.09 0.67 $\pm$ 0.28 0.52 $\pm$ 0.07 2.06 $\pm$ 0.30 4.85 $\pm$ 0.31 3.94 $\pm$ 0.42 0.50 $\pm$ 0.05 2.50 $\pm$ 0.03 1.94 $\pm$ 1.59 1.00 / 1.00 32 D 2.33 $\pm$ 0.30 0.62 $\pm$ 0.17 0.24 $\pm$ 0.05 0.50 $\pm$ 0.16 0.37 $\pm$ 0.07 1.86 $\pm$ 0.25 4.73 $\pm$ 0.35 3.81 $\pm$ 0.59 0.39 $\pm$ 0.04 2.46 $\pm$ 0.05 1.77 $\pm$ 1.57 1.00 / 1.00 64 D 1.67 $\pm$ 0.18 0.43 $\pm$ 0.16 0.13 $\pm$ 0.04 0.29 $\pm$ 0.02 0.23 $\pm$ 0.02 1.32 $\pm$ 0.28 3.99 $\pm$ 0.36 3.41 $\pm$ 0.28 0.32 $\pm$ 0.05 2.35 $\pm$ 0.11 1.45 $\pm$ 1.39 1.00 / 0.99 128 D 1.12 $\pm$ 0.02 0.23 $\pm$ 0.00 0.04 $\pm$ 0.03 0.21 $\pm$ 0.04 0.22 $\pm$ 0.03 1.08 $\pm$ 0.06 3.05 $\pm$ 0.87 3.09 $\pm$ 0.37 0.26 $\pm$ 0.03 2.31 $\pm$ 0.04 1.19 $\pm$ 1.21 1.00 / 0.98 256 D 0.54 $\pm$ 0.03 0.18 $\pm$ 0.01 0.01 $\pm$ 0.00 0.16 $\pm$ 0.01 0.15 $\pm$ 0.01 0.92 $\pm$ 0.08 2.42 $\pm$ 0.14 2.51 $\pm$ 0.13 0.19 $\pm$ 0.01 2.09 $\pm$ 0.02 0.94 $\pm$ 0.99 1.00 / 0.97 500 full (F) 16 F 2.14 $\pm$ 0.06 0.70 $\pm$ 0.04 0.28 $\pm$ 0.00 0.27 $\pm$ 0.01 0.21 $\pm$ 0.00 1.14 $\pm$ 0.04 3.06 $\pm$ 0.27 2.71 $\pm$ 0.01 0.34 $\pm$ 0.01 2.21 $\pm$ 0.01 1.33 $\pm$ 1.09 1.00/1.00 32 F 1.17 $\pm$ 0.07 0.48 $\pm$ 0.03 0.08 $\pm$ 0.04 0.21 $\pm$ 0.01 0.17 $\pm$ 0.00 0.99 $\pm$ 0.04 2.69 $\pm$ 0.10 2.47 $\pm$ 0.02 0.25 $\pm$ 0.02 2.11 $\pm$ 0.04 1.08 $\pm$ 0.99 0.99/0.99 64 F 0.51 $\pm$ 0.03 0.21 $\pm$ 0.00 0.02 $\pm$ 0.00 0.17 $\pm$ 0.01 0.14 $\pm$ 0.00 0.88 $\pm$ 0.04 2.19 $\pm$ 0.14 2.34 $\pm$ 0.03 0.20 $\pm$ 0.00 1.97 $\pm$ 0.02 0.89 $\pm$ 0.91 0.97/0.96 128 F 0.39 $\pm$ 0.01 0.16 $\pm$ 0.00 0.01 $\pm$ 0.00 0.13 $\pm$ 0.00 0.11 $\pm$ 0.00 0.81 $\pm$ 0.03 1.42 $\pm$ 0.07 2.03 $\pm$ 0.01 0.16 $\pm$ 0.00 1.71 $\pm$ 0.01 0.71 $\pm$ 0.74 0.88/0.86 256 F 0.32 $\pm$ 0.01 0.15 $\pm$ 0.00 0.01 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.77 $\pm$ 0.01 1.18 $\pm$ 0.04 1.96 $\pm$ 0.00 0.14 $\pm$ 0.01 1.25 $\pm$ 0.00 0.61 $\pm$ 0.65 0.50/0.47
Table 9: Absolute In-Distribution test loss for various tasks and methods
F.8 Agreement and Compression Rate
Table 10 provides full results for agreement, which shows the same trends as the results for relative performance of Rouge-L (Table 3). Note that agreement measures the exact match in task generations between the uncompressed LoRA model and the compressed LoRA model, rather than comparing to the task’s ground truth data. The comparison is very strict and requires an exact match between the generations of the two models (LoRA and the compressed LoRA), comparing each sample one at a time.
Model Type Method Type Tasks Average Para. Saved task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.10 $\pm$ 0.30 1.00 / 1.00 lora 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 0.00 / 0.00 TIES 10 41.00 $\pm$ 0.00 53.67 $\pm$ 0.58 44.33 $\pm$ 4.04 10.33 $\pm$ 0.58 46.33 $\pm$ 4.04 1.00 $\pm$ 0.00 8.00 $\pm$ 0.00 8.00 $\pm$ 0.00 76.67 $\pm$ 1.15 1.00 $\pm$ 0.00 29.03 $\pm$ 25.69 1.00 / 1.00 50 24.00 $\pm$ 0.00 38.67 $\pm$ 0.58 17.67 $\pm$ 4.62 2.00 $\pm$ 0.00 56.33 $\pm$ 0.58 1.00 $\pm$ 0.00 8.00 $\pm$ 0.00 8.00 $\pm$ 0.00 29.67 $\pm$ 2.89 0.00 $\pm$ 0.00 18.53 $\pm$ 18.07 1.00 / 1.00 100 22.00 $\pm$ 0.00 38.00 $\pm$ 0.00 18.67 $\pm$ 4.62 1.00 $\pm$ 1.73 51.67 $\pm$ 4.62 1.00 $\pm$ 0.00 8.00 $\pm$ 0.00 7.33 $\pm$ 0.58 2.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.97 $\pm$ 17.20 1.00 / 1.00 500 8.00 $\pm$ 0.00 25.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 59.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.00 $\pm$ 0.00 6.00 $\pm$ 0.00 2.00 $\pm$ 0.00 0.00 $\pm$ 0.00 9.90 $\pm$ 18.12 1.00 / 1.00 SVD SVD 2 88.33 $\pm$ 0.65 91.91 $\pm$ 0.94 100.00 $\pm$ 0.00 97.25 $\pm$ 0.45 92.83 $\pm$ 0.39 76.50 $\pm$ 1.51 66.00 $\pm$ 1.41 58.08 $\pm$ 1.16 98.67 $\pm$ 0.49 5.83 $\pm$ 0.94 77.42 $\pm$ 27.69 0.88 / 0.88 SVD 4 93.00 $\pm$ 0.00 96.64 $\pm$ 0.50 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 96.75 $\pm$ 0.87 88.83 $\pm$ 1.53 90.67 $\pm$ 1.23 72.17 $\pm$ 0.58 98.67 $\pm$ 0.49 16.67 $\pm$ 1.78 85.24 $\pm$ 24.39 0.75 / 0.75 SVD 8 98.89 $\pm$ 0.60 98.55 $\pm$ 0.52 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 99.42 $\pm$ 0.51 93.44 $\pm$ 0.73 97.22 $\pm$ 0.44 82.78 $\pm$ 1.64 99.00 $\pm$ 0.00 60.00 $\pm$ 0.87 93.70 $\pm$ 11.59 0.50 / 0.50 SVD 16 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 99.67 $\pm$ 0.50 99.50 $\pm$ 0.55 99.67 $\pm$ 0.50 100.00 $\pm$ 0.00 98.11 $\pm$ 0.78 99.69 $\pm$ 0.68 0.00 / 0.00 10 diagonal (D) 16 D 83.33 $\pm$ 1.53 88.33 $\pm$ 0.58 100.00 $\pm$ 0.00 97.00 $\pm$ 2.00 88.33 $\pm$ 1.15 57.00 $\pm$ 1.00 48.67 $\pm$ 3.21 50.67 $\pm$ 4.93 97.67 $\pm$ 1.53 5.33 $\pm$ 1.15 71.63 $\pm$ 29.53 1.00 / 0.90 32 D 93.00 $\pm$ 1.00 95.33 $\pm$ 0.58 100.00 $\pm$ 0.00 98.00 $\pm$ 1.00 93.67 $\pm$ 1.53 80.67 $\pm$ 2.31 78.67 $\pm$ 1.15 68.00 $\pm$ 1.73 98.33 $\pm$ 0.58 14.67 $\pm$ 2.52 82.03 $\pm$ 24.99 1.00 / 0.80 64 D 99.00 $\pm$ 0.00 97.00 $\pm$ 1.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 98.00 $\pm$ 0.00 90.67 $\pm$ 1.53 95.33 $\pm$ 1.15 79.67 $\pm$ 1.53 99.00 $\pm$ 0.00 55.00 $\pm$ 4.36 91.37 $\pm$ 13.78 1.00 / 0.60 128 D 100.00 $\pm$ 0.00 99.33 $\pm$ 0.58 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 96.67 $\pm$ 1.53 98.33 $\pm$ 1.15 95.67 $\pm$ 2.31 100.00 $\pm$ 0.00 91.67 $\pm$ 3.51 98.17 $\pm$ 2.94 1.00 / 0.20 256 D 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 99.33 $\pm$ 1.15 100.00 $\pm$ 0.00 95.00 $\pm$ 1.00 99.43 $\pm$ 1.57 1.00 / -0.60 10 full (F) 16 F 83.00 $\pm$ 2.00 93.00 $\pm$ 1.00 100.00 $\pm$ 0.00 98.33 $\pm$ 0.58 91.67 $\pm$ 0.58 64.33 $\pm$ 3.21 59.33 $\pm$ 1.15 52.67 $\pm$ 1.53 98.33 $\pm$ 0.58 6.33 $\pm$ 1.15 74.70 $\pm$ 28.71 1.00/0.90 32 F 91.33 $\pm$ 0.58 96.00 $\pm$ 1.00 100.00 $\pm$ 0.00 98.33 $\pm$ 0.58 94.33 $\pm$ 0.58 84.00 $\pm$ 2.00 83.00 $\pm$ 1.73 70.33 $\pm$ 1.53 99.00 $\pm$ 1.00 22.00 $\pm$ 2.65 83.83 $\pm$ 22.82 0.99/0.79 64 F 99.00 $\pm$ 0.00 97.33 $\pm$ 0.58 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 99.33 $\pm$ 1.15 91.33 $\pm$ 1.53 96.33 $\pm$ 0.58 81.67 $\pm$ 2.31 99.00 $\pm$ 0.00 58.33 $\pm$ 1.53 92.23 $\pm$ 12.76 0.97/0.57 128 F 99.67 $\pm$ 0.58 99.33 $\pm$ 0.58 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 97.67 $\pm$ 1.15 100.00 $\pm$ 0.00 95.67 $\pm$ 1.15 100.00 $\pm$ 0.00 91.00 $\pm$ 1.00 98.33 $\pm$ 2.89 0.88/0.07 256 F 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 99.67 $\pm$ 0.58 99.67 $\pm$ 0.58 99.67 $\pm$ 0.58 100.00 $\pm$ 0.00 98.00 $\pm$ 1.00 99.70 $\pm$ 0.70 0.50/-1.10 50 diagonal (D) 16 D 52.67 $\pm$ 4.51 86.67 $\pm$ 3.06 100.00 $\pm$ 0.00 85.00 $\pm$ 3.46 65.33 $\pm$ 3.79 25.33 $\pm$ 5.03 10.00 $\pm$ 1.00 10.67 $\pm$ 10.26 81.00 $\pm$ 6.56 0.00 $\pm$ 0.00 51.67 $\pm$ 36.18 1.00 / 0.98 32 D 69.67 $\pm$ 3.21 88.67 $\pm$ 1.53 100.00 $\pm$ 0.00 95.00 $\pm$ 2.00 80.00 $\pm$ 3.00 36.67 $\pm$ 5.13 17.00 $\pm$ 2.65 26.33 $\pm$ 5.03 95.00 $\pm$ 2.00 0.00 $\pm$ 0.00 60.83 $\pm$ 36.02 1.00 / 0.96 64 D 79.67 $\pm$ 2.52 91.00 $\pm$ 1.00 100.00 $\pm$ 0.00 97.67 $\pm$ 0.58 88.00 $\pm$ 1.00 52.00 $\pm$ 1.00 36.67 $\pm$ 5.69 41.33 $\pm$ 1.15 96.00 $\pm$ 1.00 0.33 $\pm$ 0.58 68.27 $\pm$ 32.66 1.00 / 0.92 128 D 90.00 $\pm$ 1.00 91.33 $\pm$ 0.58 100.00 $\pm$ 0.00 98.33 $\pm$ 0.58 90.67 $\pm$ 2.08 73.67 $\pm$ 2.08 63.67 $\pm$ 1.53 56.33 $\pm$ 0.58 98.00 $\pm$ 0.00 7.33 $\pm$ 1.15 76.93 $\pm$ 27.79 1.00 / 0.84 256 D 94.67 $\pm$ 0.58 96.33 $\pm$ 0.58 100.00 $\pm$ 0.00 99.67 $\pm$ 0.58 96.33 $\pm$ 1.15 87.33 $\pm$ 0.58 87.00 $\pm$ 2.65 71.67 $\pm$ 1.53 99.67 $\pm$ 0.58 31.67 $\pm$ 1.15 86.43 $\pm$ 20.41 1.00 / 0.68 50 full (F) 16 F 61.67 $\pm$ 3.06 89.67 $\pm$ 1.15 99.67 $\pm$ 0.58 90.67 $\pm$ 2.52 78.33 $\pm$ 3.51 34.00 $\pm$ 1.00 7.00 $\pm$ 3.46 25.00 $\pm$ 6.24 90.00 $\pm$ 1.00 0.00 $\pm$ 0.00 57.60 $\pm$ 36.59 1.00/0.98 32 F 71.00 $\pm$ 1.00 89.00 $\pm$ 1.73 100.00 $\pm$ 0.00 98.00 $\pm$ 0.00 85.00 $\pm$ 1.00 47.00 $\pm$ 1.73 29.00 $\pm$ 3.00 35.00 $\pm$ 2.00 98.00 $\pm$ 1.00 0.00 $\pm$ 0.00 65.20 $\pm$ 34.00 0.99/0.95 64 F 81.67 $\pm$ 0.58 93.67 $\pm$ 1.15 100.00 $\pm$ 0.00 98.33 $\pm$ 0.58 90.67 $\pm$ 2.08 61.67 $\pm$ 1.53 54.33 $\pm$ 1.53 51.33 $\pm$ 1.15 98.33 $\pm$ 0.58 3.33 $\pm$ 0.58 73.33 $\pm$ 29.89 0.97/0.89 128 F 91.00 $\pm$ 1.00 94.33 $\pm$ 0.58 100.00 $\pm$ 0.00 99.00 $\pm$ 0.00 93.33 $\pm$ 1.53 81.67 $\pm$ 0.58 75.00 $\pm$ 1.73 67.67 $\pm$ 2.08 98.67 $\pm$ 0.58 16.67 $\pm$ 0.58 81.73 $\pm$ 24.46 0.88/0.72 256 F 97.00 $\pm$ 0.00 98.00 $\pm$ 0.00 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 99.67 $\pm$ 0.58 92.00 $\pm$ 0.00 94.33 $\pm$ 1.15 79.67 $\pm$ 1.53 99.00 $\pm$ 0.00 57.33 $\pm$ 2.52 91.70 $\pm$ 13.11 0.50/0.18 100 diagonal (D) 16 D 33.00 $\pm$ 8.19 79.33 $\pm$ 5.69 89.33 $\pm$ 5.51 80.00 $\pm$ 3.61 35.33 $\pm$ 6.03 4.00 $\pm$ 1.73 3.00 $\pm$ 1.00 0.00 $\pm$ 0.00 71.33 $\pm$ 4.51 0.00 $\pm$ 0.00 39.53 $\pm$ 36.15 1.00 / 0.99 32 D 51.00 $\pm$ 7.81 90.00 $\pm$ 1.00 100.00 $\pm$ 0.00 88.00 $\pm$ 7.00 58.67 $\pm$ 11.68 17.67 $\pm$ 10.26 7.67 $\pm$ 2.89 9.33 $\pm$ 12.86 86.33 $\pm$ 2.52 0.00 $\pm$ 0.00 50.87 $\pm$ 38.43 1.00 / 0.98 64 D 68.00 $\pm$ 2.65 87.33 $\pm$ 1.53 100.00 $\pm$ 0.00 94.33 $\pm$ 4.04 80.33 $\pm$ 2.08 38.00 $\pm$ 3.00 19.67 $\pm$ 4.51 28.33 $\pm$ 1.53 92.67 $\pm$ 1.15 0.33 $\pm$ 0.58 60.90 $\pm$ 34.91 1.00 / 0.96 128 D 82.00 $\pm$ 2.00 90.00 $\pm$ 2.00 100.00 $\pm$ 0.00 97.33 $\pm$ 0.58 85.33 $\pm$ 0.58 55.33 $\pm$ 2.08 34.33 $\pm$ 3.79 36.67 $\pm$ 2.52 94.67 $\pm$ 0.58 0.00 $\pm$ 0.00 67.57 $\pm$ 32.95 1.00 / 0.92 256 D 90.00 $\pm$ 1.00 93.00 $\pm$ 2.00 100.00 $\pm$ 0.00 97.67 $\pm$ 0.58 91.67 $\pm$ 0.58 71.67 $\pm$ 5.13 59.67 $\pm$ 1.53 58.00 $\pm$ 0.00 97.67 $\pm$ 0.58 4.00 $\pm$ 1.00 76.33 $\pm$ 28.88 1.00 / 0.84 100 full (F) 16 F 49.00 $\pm$ 2.00 89.67 $\pm$ 3.21 97.00 $\pm$ 3.00 84.33 $\pm$ 3.06 65.33 $\pm$ 2.52 20.67 $\pm$ 8.33 6.33 $\pm$ 2.08 8.33 $\pm$ 4.73 81.33 $\pm$ 2.08 0.00 $\pm$ 0.00 50.20 $\pm$ 37.06 1.00/0.99 32 F 65.00 $\pm$ 3.46 90.33 $\pm$ 1.53 100.00 $\pm$ 0.00 96.33 $\pm$ 1.53 80.00 $\pm$ 2.65 41.33 $\pm$ 3.21 16.00 $\pm$ 0.00 29.33 $\pm$ 2.08 92.00 $\pm$ 2.65 0.00 $\pm$ 0.00 61.03 $\pm$ 35.43 0.99/0.97 64 F 72.33 $\pm$ 0.58 89.67 $\pm$ 1.53 100.00 $\pm$ 0.00 97.67 $\pm$ 0.58 86.00 $\pm$ 1.00 53.00 $\pm$ 1.00 35.33 $\pm$ 1.53 38.00 $\pm$ 1.73 94.67 $\pm$ 0.58 0.00 $\pm$ 0.00 66.67 $\pm$ 32.54 0.97/0.93 128 F 84.33 $\pm$ 1.53 92.33 $\pm$ 1.53 100.00 $\pm$ 0.00 98.00 $\pm$ 0.00 91.33 $\pm$ 0.58 68.67 $\pm$ 0.58 56.00 $\pm$ 1.00 57.67 $\pm$ 1.15 99.00 $\pm$ 0.00 5.33 $\pm$ 0.58 75.27 $\pm$ 28.67 0.88/0.80 256 F 91.67 $\pm$ 1.15 96.67 $\pm$ 0.58 100.00 $\pm$ 0.00 100.00 $\pm$ 0.00 94.33 $\pm$ 0.58 84.67 $\pm$ 0.58 78.00 $\pm$ 0.00 69.67 $\pm$ 0.58 99.00 $\pm$ 0.00 22.00 $\pm$ 1.00 83.60 $\pm$ 23.08 0.50/0.34 500 diagonal (D) 16 D 8.00 $\pm$ 3.61 51.50 $\pm$ 3.54 79.67 $\pm$ 4.93 28.00 $\pm$ 42.44 56.67 $\pm$ 2.89 0.67 $\pm$ 1.15 0.33 $\pm$ 0.58 0.00 $\pm$ 0.00 71.67 $\pm$ 6.03 0.00 $\pm$ 0.00 28.90 $\pm$ 33.54 1.00 / 1.00 32 D 14.33 $\pm$ 11.02 52.50 $\pm$ 9.19 80.67 $\pm$ 2.08 43.00 $\pm$ 31.48 60.67 $\pm$ 0.58 0.67 $\pm$ 1.15 1.67 $\pm$ 1.15 0.00 $\pm$ 0.00 74.33 $\pm$ 4.04 0.00 $\pm$ 0.00 32.10 $\pm$ 33.43 1.00 / 1.00 64 D 25.67 $\pm$ 1.15 62.50 $\pm$ 12.02 87.33 $\pm$ 4.04 78.33 $\pm$ 1.15 65.33 $\pm$ 3.06 5.33 $\pm$ 3.51 3.67 $\pm$ 1.15 1.00 $\pm$ 1.73 76.67 $\pm$ 2.31 0.00 $\pm$ 0.00 39.83 $\pm$ 35.80 1.00 / 0.99 128 D 38.33 $\pm$ 3.21 85.50 $\pm$ 2.12 96.00 $\pm$ 3.00 81.33 $\pm$ 2.31 65.67 $\pm$ 1.15 11.67 $\pm$ 4.73 5.33 $\pm$ 2.08 2.00 $\pm$ 1.00 80.00 $\pm$ 5.00 0.00 $\pm$ 0.00 45.24 $\pm$ 37.78 1.00 / 0.98 256 D 53.33 $\pm$ 0.58 91.00 $\pm$ 2.83 100.00 $\pm$ 0.00 89.00 $\pm$ 2.65 76.00 $\pm$ 2.00 20.67 $\pm$ 6.81 6.00 $\pm$ 1.73 12.00 $\pm$ 8.00 86.33 $\pm$ 2.31 0.00 $\pm$ 0.00 52.14 $\pm$ 38.64 1.00 / 0.97 500 full (F) 16 F 8.33 $\pm$ 2.08 41.00 $\pm$ 5.66 76.67 $\pm$ 0.58 78.00 $\pm$ 0.00 72.67 $\pm$ 0.58 6.00 $\pm$ 0.00 5.67 $\pm$ 0.58 0.00 $\pm$ 0.00 78.00 $\pm$ 1.00 0.00 $\pm$ 0.00 36.48 $\pm$ 35.46 1.00/1.00 32 F 33.67 $\pm$ 4.16 51.00 $\pm$ 1.41 92.67 $\pm$ 1.53 77.00 $\pm$ 1.73 75.00 $\pm$ 2.00 14.33 $\pm$ 1.53 8.00 $\pm$ 0.00 0.00 $\pm$ 0.00 80.67 $\pm$ 1.53 0.00 $\pm$ 0.00 42.97 $\pm$ 35.80 0.99/0.99 64 F 56.00 $\pm$ 2.65 85.50 $\pm$ 0.71 94.33 $\pm$ 0.58 89.33 $\pm$ 2.89 74.33 $\pm$ 1.15 36.33 $\pm$ 1.15 9.00 $\pm$ 1.00 2.67 $\pm$ 1.15 84.00 $\pm$ 1.00 0.00 $\pm$ 0.00 52.03 $\pm$ 36.92 0.97/0.96 128 F 69.33 $\pm$ 0.58 88.50 $\pm$ 0.71 99.00 $\pm$ 0.00 96.33 $\pm$ 1.53 80.33 $\pm$ 1.15 45.00 $\pm$ 2.00 16.33 $\pm$ 0.58 31.00 $\pm$ 1.73 92.00 $\pm$ 0.00 0.00 $\pm$ 0.00 60.86 $\pm$ 35.07 0.88/0.86 256 F 79.67 $\pm$ 0.58 89.50 $\pm$ 0.71 100.00 $\pm$ 0.00 97.67 $\pm$ 0.58 87.33 $\pm$ 0.58 57.00 $\pm$ 1.00 35.00 $\pm$ 1.00 42.00 $\pm$ 1.00 95.00 $\pm$ 1.00 0.00 $\pm$ 0.00 67.59 $\pm$ 32.67 0.50/0.47
Table 10: Absolute In-Distribution agreement for various tasks and methods
F.9 Reconstruction Error and Compression Rate
Table 11 provides the full results of the experiments behind Figure 4 for every evaluation task.
Model Type Method Type Tasks Average task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 SVD SVD 2 0.29 $\pm$ 0.00 0.43 $\pm$ 0.00 0.31 $\pm$ 0.00 0.40 $\pm$ 0.00 0.38 $\pm$ 0.00 0.31 $\pm$ 0.00 0.37 $\pm$ 0.00 0.31 $\pm$ 0.00 0.42 $\pm$ 0.00 0.30 $\pm$ 0.00 0.35 $\pm$ 0.05 SVD 4 0.16 $\pm$ 0.00 0.24 $\pm$ 0.00 0.16 $\pm$ 0.00 0.25 $\pm$ 0.00 0.23 $\pm$ 0.00 0.17 $\pm$ 0.00 0.22 $\pm$ 0.00 0.16 $\pm$ 0.00 0.25 $\pm$ 0.00 0.16 $\pm$ 0.00 0.20 $\pm$ 0.04 SVD 8 0.06 $\pm$ 0.00 0.09 $\pm$ 0.00 0.06 $\pm$ 0.00 0.11 $\pm$ 0.00 0.10 $\pm$ 0.00 0.07 $\pm$ 0.00 0.09 $\pm$ 0.00 0.06 $\pm$ 0.00 0.11 $\pm$ 0.00 0.06 $\pm$ 0.00 0.08 $\pm$ 0.02 10 diagonal (D) 16 D 0.37 $\pm$ 0.02 0.51 $\pm$ 0.02 0.36 $\pm$ 0.01 0.57 $\pm$ 0.02 0.55 $\pm$ 0.00 0.39 $\pm$ 0.02 0.49 $\pm$ 0.01 0.36 $\pm$ 0.02 0.53 $\pm$ 0.03 0.39 $\pm$ 0.01 0.45 $\pm$ 0.08 32 D 0.21 $\pm$ 0.01 0.28 $\pm$ 0.00 0.20 $\pm$ 0.01 0.35 $\pm$ 0.00 0.33 $\pm$ 0.01 0.22 $\pm$ 0.01 0.31 $\pm$ 0.01 0.20 $\pm$ 0.01 0.32 $\pm$ 0.01 0.22 $\pm$ 0.00 0.26 $\pm$ 0.06 64 D 0.10 $\pm$ 0.00 0.11 $\pm$ 0.01 0.09 $\pm$ 0.00 0.18 $\pm$ 0.01 0.18 $\pm$ 0.00 0.10 $\pm$ 0.00 0.15 $\pm$ 0.01 0.09 $\pm$ 0.00 0.14 $\pm$ 0.00 0.09 $\pm$ 0.00 0.12 $\pm$ 0.04 128 D 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.02 $\pm$ 0.00 0.03 $\pm$ 0.00 0.04 $\pm$ 0.00 0.02 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.03 $\pm$ 0.01 256 D 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 10 full (F) 16 F 0.35 $\pm$ 0.00 0.46 $\pm$ 0.00 0.34 $\pm$ 0.00 0.51 $\pm$ 0.00 0.47 $\pm$ 0.01 0.36 $\pm$ 0.01 0.45 $\pm$ 0.01 0.35 $\pm$ 0.01 0.49 $\pm$ 0.00 0.35 $\pm$ 0.01 0.41 $\pm$ 0.06 32 F 0.20 $\pm$ 0.00 0.24 $\pm$ 0.00 0.20 $\pm$ 0.00 0.30 $\pm$ 0.00 0.29 $\pm$ 0.00 0.22 $\pm$ 0.00 0.27 $\pm$ 0.00 0.20 $\pm$ 0.00 0.27 $\pm$ 0.00 0.21 $\pm$ 0.00 0.24 $\pm$ 0.04 64 F 0.10 $\pm$ 0.00 0.10 $\pm$ 0.00 0.09 $\pm$ 0.00 0.13 $\pm$ 0.00 0.13 $\pm$ 0.00 0.10 $\pm$ 0.00 0.12 $\pm$ 0.00 0.09 $\pm$ 0.00 0.12 $\pm$ 0.00 0.10 $\pm$ 0.00 0.11 $\pm$ 0.02 128 F 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 256 F 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 50 diagonal (D) 16 D 0.66 $\pm$ 0.01 0.69 $\pm$ 0.01 0.88 $\pm$ 0.01 0.76 $\pm$ 0.03 0.95 $\pm$ 0.02 0.91 $\pm$ 0.01 0.83 $\pm$ 0.02 0.88 $\pm$ 0.03 0.72 $\pm$ 0.02 0.88 $\pm$ 0.02 0.82 $\pm$ 0.10 32 D 0.50 $\pm$ 0.01 0.52 $\pm$ 0.02 0.73 $\pm$ 0.01 0.58 $\pm$ 0.03 0.88 $\pm$ 0.03 0.79 $\pm$ 0.03 0.72 $\pm$ 0.01 0.75 $\pm$ 0.01 0.57 $\pm$ 0.02 0.75 $\pm$ 0.01 0.68 $\pm$ 0.12 64 D 0.34 $\pm$ 0.01 0.37 $\pm$ 0.01 0.52 $\pm$ 0.00 0.38 $\pm$ 0.01 0.71 $\pm$ 0.02 0.58 $\pm$ 0.01 0.54 $\pm$ 0.00 0.56 $\pm$ 0.00 0.44 $\pm$ 0.01 0.58 $\pm$ 0.01 0.50 $\pm$ 0.11 128 D 0.21 $\pm$ 0.01 0.22 $\pm$ 0.01 0.31 $\pm$ 0.00 0.22 $\pm$ 0.00 0.51 $\pm$ 0.01 0.42 $\pm$ 0.01 0.38 $\pm$ 0.00 0.39 $\pm$ 0.00 0.27 $\pm$ 0.00 0.40 $\pm$ 0.00 0.33 $\pm$ 0.10 256 D 0.10 $\pm$ 0.00 0.12 $\pm$ 0.00 0.16 $\pm$ 0.00 0.10 $\pm$ 0.00 0.29 $\pm$ 0.01 0.21 $\pm$ 0.00 0.19 $\pm$ 0.00 0.23 $\pm$ 0.01 0.15 $\pm$ 0.00 0.20 $\pm$ 0.00 0.18 $\pm$ 0.06 50 full (F) 16 F 0.57 $\pm$ 0.01 0.60 $\pm$ 0.01 0.86 $\pm$ 0.01 0.71 $\pm$ 0.02 0.95 $\pm$ 0.01 0.88 $\pm$ 0.01 0.81 $\pm$ 0.00 0.83 $\pm$ 0.01 0.67 $\pm$ 0.01 0.86 $\pm$ 0.01 0.78 $\pm$ 0.12 32 F 0.47 $\pm$ 0.01 0.48 $\pm$ 0.01 0.71 $\pm$ 0.00 0.55 $\pm$ 0.01 0.78 $\pm$ 0.01 0.69 $\pm$ 0.01 0.69 $\pm$ 0.00 0.65 $\pm$ 0.01 0.53 $\pm$ 0.01 0.71 $\pm$ 0.00 0.63 $\pm$ 0.11 64 F 0.33 $\pm$ 0.00 0.35 $\pm$ 0.00 0.45 $\pm$ 0.00 0.36 $\pm$ 0.00 0.56 $\pm$ 0.00 0.50 $\pm$ 0.01 0.47 $\pm$ 0.00 0.49 $\pm$ 0.00 0.39 $\pm$ 0.01 0.49 $\pm$ 0.00 0.44 $\pm$ 0.08 128 F 0.19 $\pm$ 0.00 0.21 $\pm$ 0.00 0.25 $\pm$ 0.00 0.19 $\pm$ 0.00 0.35 $\pm$ 0.00 0.30 $\pm$ 0.00 0.28 $\pm$ 0.00 0.31 $\pm$ 0.00 0.24 $\pm$ 0.00 0.30 $\pm$ 0.00 0.26 $\pm$ 0.05 256 F 0.09 $\pm$ 0.00 0.10 $\pm$ 0.00 0.10 $\pm$ 0.00 0.08 $\pm$ 0.00 0.16 $\pm$ 0.00 0.13 $\pm$ 0.00 0.12 $\pm$ 0.00 0.15 $\pm$ 0.00 0.11 $\pm$ 0.00 0.13 $\pm$ 0.00 0.12 $\pm$ 0.02 100 diagonal (D) 16 D 0.90 $\pm$ 0.01 0.85 $\pm$ 0.01 0.87 $\pm$ 0.03 0.88 $\pm$ 0.02 0.68 $\pm$ 0.02 0.91 $\pm$ 0.01 0.97 $\pm$ 0.01 0.98 $\pm$ 0.01 0.96 $\pm$ 0.01 1.00 $\pm$ 0.00 0.90 $\pm$ 0.09 32 D 0.83 $\pm$ 0.02 0.77 $\pm$ 0.00 0.77 $\pm$ 0.01 0.78 $\pm$ 0.00 0.55 $\pm$ 0.02 0.79 $\pm$ 0.01 0.94 $\pm$ 0.02 0.94 $\pm$ 0.03 0.87 $\pm$ 0.00 0.98 $\pm$ 0.01 0.82 $\pm$ 0.12 64 D 0.67 $\pm$ 0.01 0.63 $\pm$ 0.00 0.59 $\pm$ 0.02 0.63 $\pm$ 0.01 0.40 $\pm$ 0.00 0.62 $\pm$ 0.01 0.86 $\pm$ 0.02 0.82 $\pm$ 0.02 0.71 $\pm$ 0.03 0.93 $\pm$ 0.00 0.68 $\pm$ 0.15 128 D 0.49 $\pm$ 0.01 0.47 $\pm$ 0.00 0.42 $\pm$ 0.01 0.45 $\pm$ 0.00 0.27 $\pm$ 0.02 0.44 $\pm$ 0.01 0.73 $\pm$ 0.01 0.69 $\pm$ 0.02 0.59 $\pm$ 0.02 0.80 $\pm$ 0.02 0.53 $\pm$ 0.16 256 D 0.32 $\pm$ 0.00 0.31 $\pm$ 0.00 0.26 $\pm$ 0.01 0.30 $\pm$ 0.00 0.15 $\pm$ 0.01 0.28 $\pm$ 0.00 0.51 $\pm$ 0.02 0.51 $\pm$ 0.02 0.40 $\pm$ 0.01 0.61 $\pm$ 0.01 0.36 $\pm$ 0.14 100 full (F) 16 F 0.88 $\pm$ 0.00 0.82 $\pm$ 0.00 0.84 $\pm$ 0.01 0.86 $\pm$ 0.00 0.67 $\pm$ 0.01 0.88 $\pm$ 0.01 0.99 $\pm$ 0.00 0.96 $\pm$ 0.01 0.91 $\pm$ 0.01 1.00 $\pm$ 0.00 0.88 $\pm$ 0.09 32 F 0.78 $\pm$ 0.00 0.72 $\pm$ 0.00 0.73 $\pm$ 0.00 0.74 $\pm$ 0.00 0.52 $\pm$ 0.00 0.74 $\pm$ 0.01 0.94 $\pm$ 0.01 0.89 $\pm$ 0.00 0.77 $\pm$ 0.02 0.99 $\pm$ 0.00 0.78 $\pm$ 0.13 64 F 0.60 $\pm$ 0.00 0.57 $\pm$ 0.00 0.57 $\pm$ 0.00 0.57 $\pm$ 0.00 0.39 $\pm$ 0.00 0.56 $\pm$ 0.00 0.76 $\pm$ 0.00 0.73 $\pm$ 0.00 0.60 $\pm$ 0.00 0.83 $\pm$ 0.01 0.62 $\pm$ 0.12 128 F 0.40 $\pm$ 0.00 0.38 $\pm$ 0.00 0.35 $\pm$ 0.00 0.37 $\pm$ 0.00 0.25 $\pm$ 0.00 0.37 $\pm$ 0.00 0.52 $\pm$ 0.00 0.54 $\pm$ 0.00 0.45 $\pm$ 0.00 0.60 $\pm$ 0.00 0.42 $\pm$ 0.10 256 F 0.21 $\pm$ 0.00 0.20 $\pm$ 0.00 0.18 $\pm$ 0.00 0.19 $\pm$ 0.00 0.13 $\pm$ 0.00 0.19 $\pm$ 0.00 0.30 $\pm$ 0.00 0.34 $\pm$ 0.00 0.26 $\pm$ 0.00 0.38 $\pm$ 0.00 0.24 $\pm$ 0.08 500 diagonal (D) 16 D 0.97 $\pm$ 0.00 0.73 $\pm$ 0.00 0.96 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.01 0.96 $\pm$ 0.01 0.90 $\pm$ 0.00 0.92 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.94 $\pm$ 0.08 32 D 0.96 $\pm$ 0.00 0.70 $\pm$ 0.00 0.92 $\pm$ 0.01 0.98 $\pm$ 0.01 0.96 $\pm$ 0.01 0.93 $\pm$ 0.01 0.86 $\pm$ 0.00 0.89 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.92 $\pm$ 0.09 64 D 0.90 $\pm$ 0.01 0.65 $\pm$ 0.00 0.86 $\pm$ 0.01 0.96 $\pm$ 0.02 0.90 $\pm$ 0.01 0.87 $\pm$ 0.01 0.81 $\pm$ 0.00 0.83 $\pm$ 0.01 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 0.88 $\pm$ 0.10 128 D 0.82 $\pm$ 0.01 0.60 $\pm$ 0.00 0.76 $\pm$ 0.00 0.90 $\pm$ 0.02 0.83 $\pm$ 0.01 0.78 $\pm$ 0.02 0.74 $\pm$ 0.00 0.74 $\pm$ 0.01 0.97 $\pm$ 0.01 1.00 $\pm$ 0.00 0.81 $\pm$ 0.12 256 D 0.59 $\pm$ 0.02 0.51 $\pm$ 0.00 0.56 $\pm$ 0.01 0.81 $\pm$ 0.02 0.70 $\pm$ 0.02 0.55 $\pm$ 0.01 0.57 $\pm$ 0.01 0.54 $\pm$ 0.01 0.91 $\pm$ 0.01 1.00 $\pm$ 0.01 0.67 $\pm$ 0.17 500 full (F) 16 F 0.94 $\pm$ 0.00 0.67 $\pm$ 0.00 0.88 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.90 $\pm$ 0.00 0.82 $\pm$ 0.00 0.83 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 0.90 $\pm$ 0.10 32 F 0.88 $\pm$ 0.00 0.61 $\pm$ 0.00 0.81 $\pm$ 0.00 0.97 $\pm$ 0.01 0.94 $\pm$ 0.00 0.84 $\pm$ 0.00 0.75 $\pm$ 0.00 0.77 $\pm$ 0.00 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00 0.86 $\pm$ 0.12 64 F 0.80 $\pm$ 0.00 0.55 $\pm$ 0.00 0.72 $\pm$ 0.00 0.86 $\pm$ 0.00 0.82 $\pm$ 0.01 0.76 $\pm$ 0.00 0.67 $\pm$ 0.00 0.70 $\pm$ 0.00 0.94 $\pm$ 0.00 0.99 $\pm$ 0.00 0.78 $\pm$ 0.13 128 F 0.64 $\pm$ 0.00 0.46 $\pm$ 0.00 0.60 $\pm$ 0.00 0.74 $\pm$ 0.00 0.65 $\pm$ 0.00 0.63 $\pm$ 0.00 0.56 $\pm$ 0.00 0.58 $\pm$ 0.00 0.85 $\pm$ 0.00 0.96 $\pm$ 0.00 0.67 $\pm$ 0.14 256 F 0.43 $\pm$ 0.00 0.35 $\pm$ 0.00 0.44 $\pm$ 0.00 0.55 $\pm$ 0.00 0.49 $\pm$ 0.00 0.45 $\pm$ 0.00 0.40 $\pm$ 0.00 0.42 $\pm$ 0.00 0.67 $\pm$ 0.00 0.84 $\pm$ 0.00 0.50 $\pm$ 0.14
Table 11: Reconstruction error In-Distribution for various tasks and methods
F.10 Reconstruction Error: Trained vs. Random
Table 12 provides the reconstruction error on random (untrained) LoRA matrices. Comparing with Table 11, we find that reconstruction error is consistently higher on random (untrained LoRA) matrices than on trained LoRA matrices. This demonstrates that after training, LoRAs have a shared structure that JD exploits.
Model Type Method Type Tasks Average task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 10 full (F) 16 F 0.46 $\pm$ 0.01 0.63 $\pm$ 0.00 0.50 $\pm$ 0.00 0.55 $\pm$ 0.01 0.50 $\pm$ 0.00 0.49 $\pm$ 0.01 0.50 $\pm$ 0.01 0.50 $\pm$ 0.01 0.61 $\pm$ 0.01 0.47 $\pm$ 0.01 0.52 $\pm$ 0.06 32 F 0.30 $\pm$ 0.01 0.37 $\pm$ 0.00 0.31 $\pm$ 0.00 0.35 $\pm$ 0.00 0.34 $\pm$ 0.00 0.31 $\pm$ 0.00 0.33 $\pm$ 0.00 0.31 $\pm$ 0.00 0.38 $\pm$ 0.00 0.30 $\pm$ 0.00 0.33 $\pm$ 0.03 64 F 0.15 $\pm$ 0.00 0.15 $\pm$ 0.00 0.16 $\pm$ 0.00 0.17 $\pm$ 0.00 0.17 $\pm$ 0.00 0.16 $\pm$ 0.00 0.16 $\pm$ 0.00 0.16 $\pm$ 0.00 0.17 $\pm$ 0.00 0.15 $\pm$ 0.00 0.16 $\pm$ 0.01 50 full (F) 16 F 0.80 $\pm$ 0.02 0.82 $\pm$ 0.01 0.85 $\pm$ 0.01 0.90 $\pm$ 0.02 0.78 $\pm$ 0.01 0.95 $\pm$ 0.01 0.76 $\pm$ 0.01 0.75 $\pm$ 0.00 0.79 $\pm$ 0.01 0.82 $\pm$ 0.01 0.82 $\pm$ 0.06 32 F 0.65 $\pm$ 0.01 0.67 $\pm$ 0.01 0.72 $\pm$ 0.01 0.76 $\pm$ 0.02 0.65 $\pm$ 0.01 0.82 $\pm$ 0.02 0.66 $\pm$ 0.01 0.65 $\pm$ 0.01 0.67 $\pm$ 0.02 0.69 $\pm$ 0.00 0.69 $\pm$ 0.06 64 F 0.50 $\pm$ 0.01 0.52 $\pm$ 0.00 0.52 $\pm$ 0.00 0.55 $\pm$ 0.01 0.52 $\pm$ 0.01 0.62 $\pm$ 0.00 0.54 $\pm$ 0.01 0.51 $\pm$ 0.00 0.54 $\pm$ 0.01 0.57 $\pm$ 0.00 0.54 $\pm$ 0.03 100 full (F) 16 F 0.93 $\pm$ 0.02 0.90 $\pm$ 0.02 0.93 $\pm$ 0.01 0.91 $\pm$ 0.02 0.88 $\pm$ 0.03 0.98 $\pm$ 0.01 0.96 $\pm$ 0.01 0.78 $\pm$ 0.00 0.82 $\pm$ 0.00 0.93 $\pm$ 0.02 0.90 $\pm$ 0.06 32 F 0.87 $\pm$ 0.01 0.81 $\pm$ 0.01 0.85 $\pm$ 0.02 0.80 $\pm$ 0.01 0.79 $\pm$ 0.02 0.91 $\pm$ 0.00 0.90 $\pm$ 0.01 0.74 $\pm$ 0.01 0.70 $\pm$ 0.02 0.85 $\pm$ 0.02 0.82 $\pm$ 0.07 64 F 0.65 $\pm$ 0.04 0.69 $\pm$ 0.01 0.71 $\pm$ 0.01 0.67 $\pm$ 0.01 0.64 $\pm$ 0.01 0.76 $\pm$ 0.01 0.77 $\pm$ 0.01 0.67 $\pm$ 0.00 0.61 $\pm$ 0.00 0.75 $\pm$ 0.06 0.69 $\pm$ 0.06 500 full (F) 16 F 0.98 $\pm$ 0.04 0.98 $\pm$ 0.01 0.99 $\pm$ 0.01 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.96 $\pm$ 0.05 0.93 $\pm$ 0.10 0.94 $\pm$ 0.09 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.05 32 F 0.92 $\pm$ 0.07 0.84 $\pm$ 0.20 0.92 $\pm$ 0.10 0.98 $\pm$ 0.02 0.97 $\pm$ 0.02 0.89 $\pm$ 0.08 0.82 $\pm$ 0.13 0.84 $\pm$ 0.11 0.99 $\pm$ 0.00 0.99 $\pm$ 0.02 0.92 $\pm$ 0.10 64 F 0.80 $\pm$ 0.00 0.67 $\pm$ 0.21 0.78 $\pm$ 0.11 0.90 $\pm$ 0.07 0.86 $\pm$ 0.08 0.76 $\pm$ 0.00 0.67 $\pm$ 0.00 0.70 $\pm$ 0.00 0.96 $\pm$ 0.03 0.99 $\pm$ 0.00 0.81 $\pm$ 0.13
Table 12: Reconstruction error on random LoRAs The error is larger in comparison to reconstructing trained (i.e., non-random) LoRAs in Table 11 for the corresponding compression methods.
F.11 Performance of Compressing Unseen/New LoRAs
Table 13 provides results for JD-full, where we project new unseen LoRAs into the shared basis. Our findings indicate that while compressing a new LoRA within the original space improves upon the baseline, it can degrades performance compared to uncompressed LoRAs
Model Type Method Type Tasks task036 task200 task220 task279 task281 task392 task520 task957 task1195 task1664 Base 11.11 21.02 26.81 8.59 48.08 12.78 15.06 46.42 61.35 43.01 Lora 87.37 94.00 99.00 94.33 80.25 94.00 99.00 47.94 97.86 100.00 10 16 14.49 27.85 82.15 62.23 53.49 63.31 43.66 48.73 69.81 54.85 32 14.64 46.41 95.83 69.89 53.63 72.66 74.86 48.12 71.25 57.27 64 19.63 67.00 98.16 71.66 55.32 80.51 79.60 48.13 76.02 63.92 50 16 16.09 82.00 97.16 70.14 55.26 76.01 83.74 47.71 77.92 61.46 32 44.28 83.00 99.00 72.66 56.68 86.55 86.29 47.74 79.44 68.21 64 57.76 88.00 99.00 74.00 56.86 92.33 88.05 47.75 84.29 71.29 100 16 22.18 84.00 98.00 71.66 55.49 82.38 88.12 48.27 78.71 66.81 32 48.45 86.00 99.00 73.66 56.85 88.32 88.05 48.02 81.81 67.19 64 60.60 90.00 99.00 74.00 59.63 92.33 90.00 48.18 84.92 71.52 500 16 60.99 86.00 99.00 73.00 56.58 90.66 91.00 47.20 85.16 71.21 32 64.09 90.00 99.00 73.66 59.58 92.00 91.00 47.37 86.84 75.23 64 66.29 89.00 100.00 71.33 59.79 93.00 92.00 47.56 88.29 74.98
Table 13: Compression new unseen LoRAs. ROUGE-L scores for various tasks and methods
F.12 Convergence
Table 14 presents outcomes where the JD-Full algorithm is executed until convergence. Our convergence criterion is defined as follows:
$\max\left(\|U_{t+1}-U_{t}U_{t}^{\top}U_{t+1}\|_{\mathrm{Fro}}/\|U_{t+1}\|_{\mathrm{Fro}},\|V_{t+1}-V_{t}V_{t}^{\top}V_{t+1}\|_{\mathrm{Fro}}/\|V_{t+1}\|_{\mathrm{Fro}}\right)<\tau$ (19)
where the tolerance threshold $\tau$ is set to 0.001. Due to the slow per-iteration computation times of the primary JD-Full algorithm, which quickly reaches an approximate optimum but then has a long tail of convergence for final digits of precision, we devised an alternative eigenvalue iteration algorithm (Appendix B.2) optimized for GPU acceleration. Our analysis indicates that adherence to this convergence criterion does not significantly alter the results.
Model Type Method Type Tasks Average task039 task190 task280 task290 task391 task442 task620 task1342 task1391 task1598 base 24.44 1.60 19.13 39.22 10.27 35.46 7.85 6.22 17.82 38.87 20.09 lora 95.00 86.00 99.00 93.67 94.33 74.88 74.40 26.68 95.00 50.32 78.93 10 full (F) 32 F 97.00 90.00 99.00 93.33 94.67 74.09 72.13 27.83 94.00 50.71 79.28 64 F 95.00 89.00 99.00 93.67 94.67 74.29 74.80 26.63 96.00 51.04 79.41 50 full (F) 32 F 96.00 88.00 99.00 93.67 92.33 72.30 75.97 29.89 94.00 45.68 78.68 64 F 98.00 89.00 99.00 93.67 93.33 72.74 76.50 29.33 96.00 45.71 79.33 100 full (F) 32 F 92.10 83.00 99.00 93.67 92.00 71.09 63.29 27.87 88.00 42.36 75.24 64 F 97.00 87.00 99.00 93.67 92.33 72.23 74.69 29.98 95.00 44.71 78.56 500 full (F) 32 F 68.92 43.00 87.00 91.67 90.67 70.08 51.16 14.40 83.00 41.97 64.19 64 F 93.50 78.00 91.00 92.33 90.33 72.55 57.49 15.44 85.00 42.31 71.80
Table 14: Performance with convergence In-Distribution Rouge-L
F.13 Out-of-distribution Performance (LoRA-hub)
For completeness, we incorporate results using the protocol of LoRA-hub [9]. That is, 100 LoRA-adapters are sampled, independent of the evaluation task, representing a measure of out-of-distribution performance. This also means that each result on a task is averaged across all 100 LoRA-adapters (as there is no a priori LoRA-to-task mapping). These results were obtained without normalizing the LoRA-adapters before applying the JD algorithms, a step we later identified as beneficial. We present performance comparison in Table 16. Table 15 presents the average agreement between uncompressed and compressed LoRA across 10 evaluation tasks. Results per task for JD-diagonal and JD-full are shown in Table 17 and Table 18, respectively.
From these tables, we find that the JD algorithms successfully maintain performance in this out-of-distribution context.
Table 15: Agreement Comparison. 100 LoRAs
Configuration Agreement (%) Base Model 83.015 Uncompressed LoRAs 100.000 Joint Compression Diagonal Rank 8 87.032 Rank 16 88.908 Rank 32 91.545 Rank 64 94.659 Full Rank 8 87.686 Rank 16 90.163 Rank 32 94.018 Rank 64 96.918
Table 16: Performance Comparison. 100 LoRAs
Configuration Average Performance Base Model 32.28 Uncompressed LoRAs 48.32 Join Compression Diagonal Rank 8 41.90 Rank 16 45.44 Rank 32 46.89 Rank 64 47.43 Full Rank 8 43.88 Rank 16 45.79 Rank 32 46.83 Rank 64 47.66
Table 17: Task-Based Performance Evaluation Across Different Models and Ranks
Task Base Model LoRA Diagonal R8 Diagonal R16 Diagonal R32 Diagonal R64 Causal Judgement 57.47 64.37 55.17 58.62 58.62 58.62 Date Understanding 15.33 23.33 20.67 22.00 21.33 22.67 Formal Fallacies 51.33 56.00 52.67 52.67 53.33 54.67 Hyperbaton 6.67 68.00 57.33 63.33 67.33 68.00 Logical Deduction (5 Objects) 21.33 37.33 32.00 36.67 37.33 37.33 Logical Deduction (7 Objects) 12.67 44.00 31.33 42.67 44.67 45.33 Movie Recommendation 62.67 67.33 62.00 64.67 66.67 67.33 Object Counting 34.67 38.00 35.33 36.67 36.67 38.00 Snarks 50.00 61.54 53.85 56.41 58.97 57.69 Temporal Sequences 16.67 23.33 18.67 20.67 24.00 24.67 Average 32.88 48.32 41.90 45.44 46.89 47.43
Table 18: Task-Based Performance Evaluation Across Different Models and Ranks
Task Base Model LoRA Full R8 Full R16 Full R32 Full R64 Causal Judgement 57.47 64.37 56.32 57.47 58.62 60.92 Date Understanding 15.33 23.33 19.33 22.00 22.67 22.67 Formal Fallacies 51.33 56.00 51.33 52.67 53.33 56.00 Hyperbaton 6.67 68.00 63.33 66.00 69.33 68.00 Logical Deduction (5 Objects) 21.33 37.33 35.33 36.00 35.33 37.33 Logical Deduction (7 Objects) 12.67 44.00 40.00 44.67 44.67 44.67 Movie Recommendation 62.67 67.33 63.33 65.33 67.33 67.33 Object Counting 34.67 38.00 35.33 36.67 37.33 37.33 Snarks 50.00 61.54 53.85 55.13 57.69 58.97 Temporal Sequences 16.67 23.33 20.67 22.00 22.00 23.33 Average 32.88 48.32 43.88 45.79 46.83 47.66